### v1

# String graphs are quasi-isometric to planar graphs

Abstract
We prove that for every countable string graph $S$, there is a planar graph $G$ with $V(G)=V(S)$ such that
$\frac{1}{23660800}d_{S}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{S}(u,v)$
for all $u,v\in V(S)$, where $d_{S}(u,v)$, $d_{G}(u,v)$ denotes the distance between $u$ and $v$ in $S$ and $G$ respectively. In other words, string graphs are quasi-isometric to planar graphs.
This theorem lifts a number of theorems from planar graphs to string graphs, we give some examples. String graphs have Assouad-Nagata (and asymptotic dimension) at most 2. Connected, locally finite, quasi-transitive string graphs are accessible. A finitely generated group $\Gamma$ is virtually a free product of free and surface groups if and only if $\Gamma$ is quasi-isometric to a string graph.
Two further corollaries are that countable planar metric graphs and complete Riemannian planes are also quasi-isometric to planar graphs, which answers a question of Georgakopoulos and Papasoglu. For finite string graphs and planar metric graphs, our proofs yield polynomial time (for string graphs, this is in terms of the size of a representation given in the input) algorithms for generating such quasi-isometric planar graphs.
1 Introduction
A string graph is an intersection graph of a collection $\mathcal{S}$ of path-connected subsets of the plane, where the vertices are $\mathcal{S}$, and $S_{1},S_{2}\in\mathcal{S}$ are adjacent if they intersect. String graphs were first formally defined by Ehrlich, Even, and Tarjan [28] but were present in earlier works of Benzer [9] on the topology of genetic structures and of Sinden [74] on electrical networks realizable by printed circuit. String graphs are the most general class of intersection graphs in the plane and significantly generalize planar graphs, which by the Koebe-Andreev-Thurston circle packing theorem [50] can be viewed as intersection graphs of discs in the plane with disjoint interiors. String graphs can be thought of as a dense or induced minor analogue of planar graphs.
There are several important structural results for string graphs. For instance, Pach and Tóth [67] proved that recognizing string graphs is decidable. Fox and Pach [31, 32, 34] (see also [62]) proved a separator theorem for string graphs. Every incomparability graph is a string graph and Fox and Pach [33] proved a partial converse of this for dense string graphs. This was used by Tomon [75] to prove that string graphs have the Erdős-Hajanl property. Very recently, Abrishami, Briański, Davies, Du, Masaříková, Rzążewski, and Walczak [1] classified string graphs with large chromatic number.
Every planar graph is a string graph, although the structure of string graphs is much more complicated. For instance, while planar graphs can be recognized in linear time [47], recognition of string graphs is NP-hard [52, 73]. There are at most $c^{n}n!$ planar graphs on $n$ vertices [22, 65], while the number of string graphs on $n$ vertices is $2^{(\frac{3}{4}+o(1)){n\choose 2}}$ [68]. Every planar graph is 4-colourable [5, 71], however there are triangle-free string graphs with arbitrarily large chromatic number [69].
We prove a partial converse to the fact that planar graphs are string graphs. In a coarse sense, every string graph is equivalent to some planar graph. For vertices $u,v$ of a graph $G$, we let $d_{G}(u,v)$ denote the distance between $u$ and $v$ in $G$.
Theorem 1.1.
Let $S$ be a countable string graph.  [^1] We remark that Theorem 1.1 does not hold for uncountable string graphs as can be seen by Wagner’s [78] characterization of uncountable planar graphs and the fact that the graph obtained from $|\mathbb{R}|$ 1-ended infinite paths by identifying their ends and then for each vertex adding $|\mathbb{R}|$ new adjacent vertices of degree 1 is a string graph. Then there is a planar graph $G$ with $V(G)=V(S)$ such that
$\frac{1}{23660800}d_{S}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{S}(u,v)$
for all $u,v\in V(S)$.
In other words, string graphs are quasi-isometric (or even bi-Lipschitz equivalent) to planar graphs. For finite string graphs the proof of Theorem˜1.1 yields a polynomial time (with respect to the size of a given representation of $S$ in the input) algorithm for generating such a planar graph $G$. Unfortunately, are string graphs that require exponentially sized representations [53]. However, natural subclasses such as segment intersection graphs or more generally intersection graphs of convex compact path-connected subsets of $\mathbb{R}^{2}$, do have polynomially sized representations.
Theorem˜1.1 has numerous applications. We first discuss some direct corollaries for string graphs. Afterwards, we shall discuss two further quasi-isometry results on metric planar graphs and complete Riemannian planes.
Asymptotic dimension was introduced by Gromov [43] for metric spaces and groups (via their Cayley graphs). This notion has had significant impact in geometric group theory, for instance Yu [79] proved that finitely generated groups with finite asymptotic dimension satisfy the Novikov conjecture. For a survey of asymptotic dimension from this point of view, see [8]. In coarse graph theory, asymptotic dimension can be thought of as a coarse analogue of graph colouring. There has recently been significant progress on the study of asymptotic dimension in the intersection of structural graph theory and geometric group theory [1, 10, 11, 19, 23, 27, 55]. There are also applications of asymptotic dimension to graph colouring and $\chi$-boundedness, for instance hereditary classes of graphs with bounded asymptotic dimension are either $\chi$-bounded or contain all Burling graphs [1].
One of the most natural classes of graphs to consider is of course planar graphs. Ostrovskii and Rosenthal [66] proved that planar graphs have bounded asymptotic dimension. Fujiwara and Papasoglu [35] improved this bound to 3 and Jørgensen and Lang [49] proved that planar graphs have asymptotic dimension and furthermore Assouad–Nagata dimension at most 2, which is tight due to the infinite planar grid. Asymptotic dimension and Assouad–Nagata dimension are persevered by quasi-isometries. Thus, by Theorem˜1.1, we obtain the following.
Theorem 1.2.
String graphs have Assouad–Nagata dimension at most 2.
Theorem˜1.2 settles the 2-dimensional case of a conjecture of Davies, Georgakopoulos, Hatzel, and McCarty [19] that sphere intersection graphs in $\mathbb{R}^{d}$ have asymptotic dimension at most $d$. For $d\geqslant 3$, such graphs have unbounded Assouad–Nagata dimension [19].
Bonamy, Bousquet, Esperet, Groenland, Liu, Pirot, and Scott [11] gave a proof that planar graphs have Assouad–Nagata dimension at most 2, which yields a polynomial time algorithm for computing the corresponding covers. Thus, the corresponding covers in Theorem˜1.2 can also be computed in polynomial time (in terms of the size of the provided representation of the string graph given in the input). For some algorithmic applications of such covers, see [7, 12, 13, 30, 48, 54, 56].
A well-known conjecture (see [44]) states that planar graphs admit embeddings into $\ell_{1}$ with bounded distortion. Theorem˜1.1 implies that this is equivalent to string graphs having embeddings into $\ell_{1}$ with bounded distortion. Rao [70] showed that $n$-vertex planar graphs have embeddings into $\ell_{1}$ with distortion at most $O(\sqrt{\log n})$, and Theorem˜1.1 extends this to string graphs.
Theorem 1.3.
Every $n$-vertex string graph has a $O(\sqrt{\log n})$-distortion embedding into $\ell_{1}$.
For a more extensive survey and discussion of applications of such embeddings, see [44].
Dunwoody [26] showed that every connected, locally finite, quasi-transitive, planar graph is accessible (see also [29, 45, 46]). MacManus [57] recently gave a remarkable extension of this theorem by showing that every connected, locally finite, quasi-transitive graph that is quasi-isometric to a planar graph is accessible. Thus, by Theorem˜1.1, Dunwoody’s [26] theorem extends from planar graphs to string graphs.
Theorem 1.4.
Every connected, locally finite, quasi-transitive string graphs is accessible.
There has been significant interest in planar Cayley graphs [6, 25, 24, 26, 38, 40, 41, 39, 61, 76] and Cayley graphs quasi-isometric to planar graphs [14, 15, 37, 57, 58, 59, 63, 77]. Closely related to this, there are several results characterizing certain finitely presented groups $\Gamma$ in terms of being quasi-isometric to certain Riemannian surfaces. For instance, $\Gamma$ is quasi-isometric to a complete Riemannian plane if and only if $\Gamma$ is a virtual surface group [14, 15, 37, 59, 63, 77]. Maillot also showed that $\Gamma$ is quasi-isometric to a complete, simply connected, planar Riemannian surface with non-empty geodesic boundary if and only if $\Gamma$ is virtually free. In both these cases, $\Gamma$ is quasi-isometric to a planar graph since both surface groups and free groups have planar Cayley graphs.
MacManus [57] used his aforementioned theorem on accessibility to extend and unify these two theorems on finitely generated groups. By Theorem˜1.1, we obtain the following further extension of MacManus’s [57] theorem (our contribution is the equivalence between the first two bullets, the rest is given in [57]).
Theorem 1.5.
Let $\Gamma$ be a finitely generated group. Then the following are equivalent.
  • $\Gamma$ is quasi-isometric to a string graph.
  • $\Gamma$ is quasi-isometric to a planar graph.
  • $\Gamma$ is quasi-isometric to a planar Cayley graph.
  • Some finite index subgroup of $\Gamma$ admits a planar Cayley graph.
  • $\Gamma$ is virtually a free product of free and surface groups.
We now turn our attention to some further quasi-isometry results on metric planar graphs and complete Riemannian planes.
A metric graph $H$ is a graph along with an assignment of non-negative edge lengths (in this paper, the length of edges in (non-metric) graphs is otherwise always uniformly equal to 1). We can view metric graphs as a length space in which the edges are replaces by intervals of the corresponding lengths. When the edge lengths are within $D\subseteq(0,\infty)$, we say that $H$ is an $D$-metric graph.
Georgakopoulos and Papasoglu [42] asked whether every planar metric graph is quasi-isometric to a planar graph. As a straightforward application of Theorem˜1.1 (see Lemma˜2.4), we answer this in positive (note that every planar metric graph can be turned into a $(0,1]$-metric planar graph by subdividing edges appropriately). Such a planar graph $G$ can be generated in polynomial time.
Theorem 1.6.
Let $H$ be a countable $(0,1]$-metric planar graph. Then there is a planar graph $G$ with $V(G)=V(H)$ such that
$\frac{1}{47321600}d_{H}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{H}(u,v)+1$
for all $u,v\in V(H)$.
Beyond the general interest in metric planar graphs and also in groups acting on planar surfaces [39], one other motivation of Georgakopoulos and Papasoglu [42] was their coarse Kuratowski conjecture that graphs forbidding $K_{5}$ and $K_{3,3}$ as $K$-fat minors are quasi-isometric to planar graphs. Although Theorem˜1.6 and Theorem˜1.1 provide substantial progress towards this conjecture, we remark that the closely related graph minor analogue that graphs forbidding $H$ as a $K$-fat minor are quasi-isometric to $H$-minor-free graphs is false [2, 21], (although it is known to be true for some small graphs $H$ [3, 4, 16, 36, 42, 60, 64]). Such $K$-fat $H$-minor-free graphs need not even be quasi-isometric to $H^{\prime}$-minor-free graphs for any graph $H^{\prime}$ [2]. However, with Distel and Hickingbotham [18], we will use the main technical result (see Theorem˜5.1) of this paper to prove an approximate graph minor version of Theorem˜1.6 that metric $H$-minor-free graphs are quasi-isometric to $H^{\prime}$-minor-free graphs. We remark that the constants of the quasi-isometries necessarily depend on $H$ [20].
It is well known (see [11, 59, 72]) that complete Riemannian planes are quasi-isometric to locally finite [^2] A metric graph is locally finite if every bounded radius ball contains at most a finite number of vertices. metric planar graphs. This was used for instance to show that complete Riemannian planes have Assouad–Nagata dimension at most 2 [11] and in Maillot’s [59] aforementioned classification finitely generated groups virtually free groups as being those quasi-isometric to certain planar Riemannian surfaces. By Theorem˜1.6, we obtain the following refinement, which significantly generalizes the previously discussed fact that finitely generated groups quasi-isometric to complete Riemannian planes are quasi-isometric to planar graphs.
Theorem 1.7.
Complete Riemannian planes are quasi-isometric to locally finite planar graphs.
In [17] we will use the main technical result (see Theorem˜5.1) of the present paper to extend Theorem˜1.7 to higher genus Riemannian surfaces and also refine the theorem so that the quasi-isometry from the graph to the Riemannian surface is a continuous embedding. Such a result does not hold for complete Riemannian surfaces of unbounded genus [17], (although, it does if we use metric graphs [11, 59, 72] instead of graphs).
In Section˜2 we introduce basic definitions, deduce Theorem˜1.6 from Theorem˜1.1, and also reduce proving Theorem˜1.1 to its finite case (Theorem˜2.5). We also introduce the necessary framework for our approach to proving Theorem˜1.1, namely through viewing string graphs as region intersection graphs, and the notion of “impressions” which provides a convenient way of finding quasi-isometries from region intersection graphs. The rest of the proof in Sections˜3, 4 and 5 is essentially a delicate technical induction. In Section˜3 we prove a weakening of Theorem˜1.1 for the subclass outerstring graphs (Theorem˜3.1). This is used in Section˜4 to find a technical structure tailored for use in the inductive argument that we call an “encasing” (see Lemma˜4.4). Finally, in Section˜5, we prove Theorem˜1.1 using Lemma˜4.4.
2 Preliminaries
[…]
Given a graph $G$ and a collection of connected vertex sets $\mathcal{H}$ of $G$, we let $\textbf{RIG}(G,\mathcal{H})$ be the graph with vertex set $\mathcal{H}$ where, $H_{1},H_{2}\in\mathcal{H}$ are adjacent if they intersect. We call $\textbf{RIG}(G,\mathcal{H})$ a region intersection graph. Region intersection graphs were introduced by Lee [54] as a generalization of string graphs (usually we consider region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ in which $G$ forbids some graph as a minor, since otherwise the class contains all graphs). Finite string graphs are exactly finite region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ for which $G$ is planar [54] and this is how we will view string graphs to prove Theorem˜1.1 (or its finite version, Theorem˜2.5).
If $\mathcal{M}$ is a collection of disjoint connected vertex sets in $G$, then we let $\textbf{IM}(G,\mathcal{M})$ be the graph with vertex set $\mathcal{M}$ where $M_{1},M_{2}\in\mathcal{M}$ are adjacent if there is an edge in $G$ between $M_{1}$ and $M_{2}$ (which case, we say that they touch). This is one way to define induced minors of a graph $G$.
Let $G$ be a graph and $\mathcal{H}$ a collection of connected vertex sets of $G$. If every edge of $G$ is contained in $G[H]$ for some $H\in\mathcal{H}$, then we say that $\mathcal{H}$ spans $G$. More generally, we say that $\mathcal{H}$ $z$-spans $G$ if for every edge $uv$ of $G$, there exists $H_{u},H_{v}\in\mathcal{H}$ with $u\in H_{u}$ and $v\in H_{v}$, such that $d_{\textbf{RIG}(G,\mathcal{H})}(H_{u},H_{v})\leqslant z$. Note that $\mathcal{H}$ spans $G$ if and only if it 0-spans $G$.
For a graph $G$ and two collections of connected vertex sets $\mathcal{H},\mathcal{M}$ of $G$ with the sets in $\mathcal{M}$ being disjoint, we say that $(G,\mathcal{M})$ is a $(x,y)$-impression of $(G,\mathcal{H})$ if
  • for every $M\in\mathcal{M}$, $I_{\mathcal{H}}(M)$ has weak diameter at most $x$ in $\textbf{RIG}(G,\mathcal{H})$,
  • for every $H\in\mathcal{H}$, $I_{\mathcal{M}}(H)$ has weak diameter at most $y$ in $\textbf{IM}(G,\mathcal{M})$, and
  • $\bigcup_{M\in\mathcal{M}}M=V(G)$.
Impressions provide a convenient way of obtaining a quasi-isometry from a region intersection graph $\textbf{RIG}(G,\mathcal{H})$ to a minor of $G$.
Proposition 2.1.
Let $(G,\mathcal{M})$ be a $(x,y)$-impression of $(G,\mathcal{H})$ where $\mathcal{H}$ $z$-spans $G$. Then there is a function $f:\mathcal{H}\to\mathcal{M}$ such that for all $H_{1},H_{2}\in\mathcal{H}$, we have
$\frac{1}{x+z}d_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2})-\frac{x}{x+z}\leqslant d_{\textbf{IM}(G,\mathcal{M})}(f(H_{1}),f(H_{2}))\leqslant 2yd_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2}),$
and for every $M\in\mathcal{M}$ there exists some $H\in\mathcal{H}$ such that $d_{\textbf{IM}(G,\mathcal{M})}(f(H),M)\leqslant y$.
Proof.
Let $f:\mathcal{H}\to\mathcal{M}$ be chosen so that $f(H)\in I_{\mathcal{M}}(H)$ for all $H\in\mathcal{H}$.
If $H_{1}$ and $H_{2}$ are adjacent in $\textbf{RIG}(G,\mathcal{H})$, then they intersect. So then $I_{\mathcal{M}}(H_{1}\cup H_{2})$ is a connected set in $\textbf{IM}(G,\mathcal{M})$ and furthermore has weak diameter at most $2y$. It follows for (possibly non-adjacent $H_{1},H_{2}\in\mathcal{H}$) that $d_{\textbf{IM}(G,\mathcal{M})}(f(H_{1}),f(H_{2}))\leqslant 2yd_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2})$.
Suppose that $d_{\textbf{IM}(G,\mathcal{M})}(f(H_{0}),f(H_{t}))=t$ and let $M_{0},\ldots,M_{t}$ be a path in $\textbf{IM}(G,\mathcal{M})$ between $M_{0}=f(H_{0})$ and $M_{t}=f(H_{t})$. For each $1\leqslant i<t$, there exists some $u_{i}v_{i}\in E(G)$ with $u_{i}\in M_{i-1}$ and $v_{i}\in M_{i}$. For each $1\leqslant i<t$, let $H_{i}^{\prime},H_{i}\in\mathcal{H}$ be such that $u_{i}\in V(H_{i}^{\prime})$, $v_{i}\in V(H_{i})$, and the distance between $H_{i}^{\prime}$ and $H_{i}$ in $\textbf{RIG}(G,\mathcal{H})$ is at most $z$. For each $1\leqslant i\leqslant t$, we have that $H_{i-1},H_{i}^{\prime}\in I_{\mathcal{H}}(M_{i-1})$. Thus, there is a path in $\textbf{RIG}(G,\mathcal{H})$ between $H_{0}$ and $H_{t}$ of length at most $(t+1)x+tz$. It now follows that for all $H_{1},H_{2}\in\mathcal{H}$, we have
$\frac{1}{x+z}d_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2})-\frac{x}{x+z}\leqslant d_{\textbf{IM}(G,\mathcal{M})}(f(H_{1}),f(H_{2}))\leqslant 2yd_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2}).$
Lastly, consider some $M\in\mathcal{M}$. Choose some $H\in I_{\mathcal{H}}(M)$. Then $M,f(H)\in I_{\mathcal{M}}(H)$, and thus $M$ is at distance at most $y$ from $f(M)$ in $\textbf{IM}(G,\mathcal{M})$. ∎
In most but not all cases we consider, we will actually have that $\mathcal{H}$ spans $G$, which means that we simply have $z=0$ in Proposition˜2.1. Given $(G,\mathcal{H})$, we could simply reduce to the case the $\mathcal{H}$ spans $G$ by removing edges of $G$ not contained in $G[H]$ for any $H\in\mathcal{H}$ (which is something we often do). However, adding edges to $G$ can also make finding impressions easier. In fact, for $G$ planar, it is convenient for inductive arguments that we can also add new connected sets to $\mathcal{H}$ in a controlled way so that our final impression can be obtained from a related impression.
Lemma 2.2.
[…]
Proof.
[…]
It is straight forward to make the quasi-isometry $f$ in Proposition˜2.1 into a bijective map by taking a minor and possibly adding some pendent edges. Note in particular that for countable graphs this preserves planarity.
Lemma 2.3.
Let $G,H$ be graphs and let $f:V(G)\to V(H)$ be a function such that for all $u,v\in V(G)$, we have
$\frac{1}{x_{1}}d_{G}(u,v)-\frac{x_{2}}{x_{1}}\leqslant d_{H}(f(u),f(v))\leqslant x_{3}d_{G}(u,v),$
and for every $v\in V(H)$ there exists some $u\in V(G)$ such that $d_{H}(f(u),v)\leqslant x_{4}$.
Then there exists some graph $H^{\prime}$ obtained from $H$ by adding vertices adjacent to individual vertices of $V(H)$ and then taking a minor, and some bijection $f^{\prime}:V(G)\to V(H^{\prime})$ such that for all $u,v\in V(G)$, we have
$\frac{1}{2x_{4}(x_{1}+x_{2})}d_{G}(u,v)\leqslant d_{H^{\prime}}(f^{\prime}(u),f^{\prime}(v))\leqslant(x_{3}+2)d_{G}(u,v).$
Proof.
Let $A=f(V(G))$. Since for every $v\in V(H)$ there exists some $u\in V(G)$ such that $d_{H}(f(u),v)\leqslant x_{4}$, we can choose a forest $F$ contained in $H$ as a subgraph with $V(F)=V(H)$ such that every connected component of $F$ contains exactly one vertex of $A$ and every vertex of $F$ is at distance at most $x_{4}$ from $A$ in $F$. Let $H^{*}$ be obtained from $H$ by for each $a\in A$, adding $|f^{-1}(a)|-1$ vertices to $H$ that are adjacent to $a$ only. Choose $f^{\prime}:V(G)\to V(H^{*})$ so that for every $a\in A$, $f^{-1}(a)$ is mapped bijectivly to $a$ and the newly added $|f^{-1}(a)|-1$ vertices adjacent to $a$. Let $H^{\prime}$ be obtained from $H^{*}$ by contracting the edges of $F$ onto $A$. Then $f^{\prime}$ is a bijection between $G$ and $H^{\prime}$.
For every $u,v\in V(G)$, we have
$\frac{1}{2x_{4}}d_{H}(f(u),f(v))\leqslant d_{H^{\prime}}(f^{\prime}(u),f^{\prime}(v))\leqslant d_{H}(f(u),f(v))+2.$
For distinct $u,v\in V(G)$, we clearly have $d_{H}(f(u),f(v))+2\leqslant 3d_{H}(f(u),f(v))\leqslant x_{3}d_{G}(u,v)+2\leqslant(x_{3}+2)d_{G}(u,v)$. For $t\geqslant 0$, we have that $\frac{1}{x_{1}+x_{2}}t\leqslant\max\{\frac{1}{x_{1}}t-\frac{x_{2}}{x_{1}},1\}$. Since $f^{\prime}$ is bijective, it therefore follows that $\frac{1}{2x_{4}(x_{1}+x_{2})}d_{G}(u,v)\leqslant d_{H^{\prime}}(f^{\prime}(u),f^{\prime}(v))$. ∎
Theorem˜1.1 for finite string graphs will be quickly deduced in Section˜5 from our main technical result (Theorem˜5.1) by applying Lemma˜2.2, Proposition˜2.1, and Lemma˜2.3 in order.
Next we show that Theorem˜1.1 implies Theorem˜1.6. This is immediate from the following lemma.
Lemma 2.4.
Let $H$ be a $(0,1]$-metric planar graph. Then there is a string graph $S$ with $V(S)=V(H)$ such that
$\frac{1}{2}d_{H}(u,v)\leqslant d_{S}(u,v)\leqslant d_{H}(u,v)+1$
for all $u,v\in V(H)$.
Proof.
Consider $H$ embedded in the plane. For each $u\in V(H)$, let $S_{u}$ be the subset of the plane consisting of all points of $H$ at distance at most 1 in $H$ from $u$. Let $S$ be the string graph with vertex set $V(H)$ in which each $u\in V(S)=V(H)$ is represented by $S_{u}$. Consider $u,v\in V(H)$ and let $P$ be a geodesic path from $u$ to $v$ in $H$. Choose vertices $x_{0},\ldots,x_{t}$ of $H$ along $P$ so that $x_{0}=u$, $x_{t}=v$ and for each $0\leqslant i<t$, the distance along $P$ between $x_{i}$ and $x_{i+1}$ is in $(1,2]$ (or if $d_{H}(u,v)\leqslant 1$, we just take $x_{0}=u$, $x_{1}=v$). So, $\lceil\frac{1}{2}d_{H}(u,v)\rceil\leqslant t\leqslant\lfloor d_{H}(u,v)\rfloor$. The lemma now follows since $x_{i}$ is adjacent to $x_{i+1}$ for each $0\leqslant i<t$. ∎
We remark that with a very similar proof to Lemma˜2.4, it is easy to extend Theorem˜1.1 to $(0,1]$-metric string graphs (again with slightly worse bounds).
For our inductive argument, it is more convenient to prove Theorem˜1.1 for finite string graphs. Section˜3, Section˜4, and Section˜5 are dedicated to proving Theorem˜1.1 for finite string graphs (Theorem˜2.5 below), which in particular, will be quickly deduced in Section˜5 from Theorem˜5.1 (which is a more technical result tailored for an inductive argument) by applying Lemma˜2.2, Proposition˜2.1, and Lemma˜2.3 in order.
Theorem 2.5.
Let $S$ be a finite string graph. Then there is a planar graph $G$ with $V(G)=V(S)$ such that
$\frac{1}{23660800}d_{S}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{S}(u,v)$
for all $u,v\in V(S)$.
We finish this section by showing that Theorem˜2.5 implies Theorem˜1.1.
Proof of Theorem˜1.1 assuming Theorem˜2.5..
Clearly it is enough to consider the case that $S$ is connected. Let $V(S)=\{x_{k}:k\geqslant 1\}$ and for each $i\leqslant 1$, let $X_{i}=\{x_{k}:1\leqslant k\leqslant i\}$. For each $i\geqslant 1$, let $\mathcal{G}_{i}$ be the collection of embedded planar graphs $G^{\prime}$ that are topologically inequivelant after possibly relabeling vertices of $V(G^{\prime})\backslash X_{i}$, such that $X_{i}\subseteq V(G^{\prime})\subseteq V(S)$, $|V(G^{\prime})|\leqslant i+\sum_{x,y\in X_{i}}162d_{S}(x,y)$, and for every $x,y\in X_{i}$, we have that
$\frac{1}{23660800}d_{S}(x,y)\leqslant d_{G^{\prime}}(x,y)\leqslant 162d_{S}(x,y).$
Note that each $\mathcal{G}_{i}$ is non-empty by Theorem˜2.5 (applied to some $S[Y]$ for some $Y$ with $X_{i}\subseteq Y$, $|Y|\leqslant|X_{i}|+\sum_{x,y\in X_{i}}d_{S}(x,y)$ such that $d_{S[Y]}(x,y)=d_{S}(x,y)$ for all $x,y\in X_{i}$). Also, each $|\mathcal{G}_{i}|$ is finite. Let $T$ be the graph with vertex set $\bigcup_{i=1}^{\infty}\mathcal{G}_{i}$ where each $G^{\prime\prime}\in\mathcal{G}_{i+1}$ is adjacent to some $G^{\prime}\in\mathcal{G}_{i}$ such that $G^{\prime}$ is topologically equivalent to a restriction of $G^{\prime\prime}$ after possibly relabeling vertices of $V(G^{\prime\prime})\backslash X_{i}$. Observe that by considering shortest paths between vertices of $X_{i}$, such a $G^{\prime}$ always exists.
By Kőnig’s infinity lemma [51], $T$ has an infinite path $G_{1},G_{2},\ldots$ in which $G_{i}\in\mathcal{G}_{i}$ for each $i\geqslant 1$. Note that for each $i\geqslant 1$, there exists some $R$ so that $V(G_{i})\subseteq V(G_{r})$ for all $r\geqslant R$. We obtain the desired countable planar graph $G$ by iteratively extending an embedded planar graph topically equivalent to $G_{i}$ to one topologically equivalent to $G_{i+1}$ for each $i\geqslant 1$ in order. ∎
3 Outerstring graphs
Outerstring graphs are equivalent to region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ for which $G$ is planar and every $H\in\mathcal{H}$ contains a vertex of the outer face of $G$. They can be thought of as an “outerplanar” analogue of string graphs, and are thus much simpler. The main aim of this section is to prove the following theorem on finding certain restricted impressions of such region intersection graphs. This will be crucial to our arguments on string graphs (or strictly speaking, Lemma˜3.4 and Lemma˜3.2 which easilly imply the following theorem will be).
Theorem 3.1.
Let $G$ be a finite planar graph and $\mathcal{H}$ be a spanning collection of connected vertex sets in $G$, each of which contains a vertex incident to the outer face of $G$. Then there is a $(770,9)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ such that each set in $\mathcal{M}$ contains a vertex incident to the outer face of $G$.
Note that if $(G,\mathcal{M})$ is an impression of $(G,\mathcal{H})$ where each graph in $\mathcal{M}$ contains a vertex incident to the outer face of $G$, then $\textbf{IM}(G,\mathcal{M})$ is an outerplanar graph.
Figure 1: The graph $F_{6}$.
By Proposition˜2.1, Theorem˜3.1 implies that outerstring graphs are quasi-isometric to outerplanar graphs. This result is already easily implied by a theorem that graphs forbidding $K_{2,3}$ as a fat minor are quasi-isometric to cacti [16, 36], where a cacti is a planar graph in which every edge is on the outer face, or equivalently, a graph in which each maximal 2-connected subgraph is either an edge of a cycle. One might hope that Theorem˜3.1 could be improved so that $\textbf{IM}(G,\mathcal{M})$ is also a cacti. However this is not possible. For every $\ell\geqslant 3$, let $F_{\ell}$ be the outerplanar graph consisting of a cycle of length $\ell$ where for every edge of this cycle, we identify the edge with an edge of another cycle of length $\ell$ (see Figure˜1 for an illustration of the graph $F_{6}$). Then it can easily be seen that there is no $x,y$ such that every $(F_{\ell},\{\{u,v\}:uv\in E(F_{\ell})\})$ has a $(x,y)$-impression $(F_{\ell},\mathcal{M}_{\ell})$ in which $\textbf{IM}(F_{\ell},\mathcal{M}_{\ell})$ is a cacti. It is however a nice exercise to show that outerplanar graphs are quasi-isometric to cacti. With this, Theorem˜3.1 and Proposition˜2.1 then yeild another proof that outerstring graphs are quasi-isometric to cacti.
[…]
Lemma 3.2.
Let $G$ be a finite outerplanar graph and $\mathcal{H}$ a spanning collection of connected vertex sets in $G$. Then there is a $(11,9)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$.
Proof.
Clearly it is enough to handle the case that $G$ is connected. Choose some $H_{0}\in\mathcal{H}$ and for every non-negative integer $i$, let $L_{i}$ be the vertices of $\textbf{RIG}(G,\mathcal{H})$ at distance $i$ from $H_{0}$ in $\textbf{RIG}(G,\mathcal{H})$. For each $i$ we define the following two vertex sets in $G$. Let $D_{i}=\bigcup_{H\in L_{i}}H$, and let $A_{i}=D_{i}\backslash D_{i-1}$ (where $D_{-1}=\emptyset$). For each $A_{i}$, let $\mathcal{M}_{i}$ be the maximal connected vertex sets of $G[A_{i}]$. Let $\mathcal{M}=\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$. See Figure˜2 for an illustration. We must show that $(G,\mathcal{M})$ is a $(11,9)$-impression of $(G,\mathcal{H})$.
Figure 2: An illustration of the resulting collection of disjoint connected sets $\mathcal{M}$. The connected sets in $\mathcal{H}$ are the red sets and the sets in $\mathcal{M}$ are the blue ones. The dashed purple lines indicate the partition of $\mathcal{M}$ into $\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$.
Next, we show for every $M\in\mathcal{M}$ that $I_{\mathcal{H}}(M)$ has weak diameter at most $11$ in $\textbf{RIG}(G,\mathcal{H})$. Let $M\in\mathcal{M}$. Then, for some non-negative integer $i$, we have that $M\in\mathcal{M}_{i}$. We may assume that $i\geqslant 2$, since otherwise we clearly have that $I_{\mathcal{H}}(M)$ has weak diameter at most $4$ in $\textbf{RIG}(G,\mathcal{H})$. Since $G$ is outerplanar, observe that there exists a $2$-seperation $(X,Y)$ of $G$ with $A_{i-2}\subseteq X$, $X\cap Y\subseteq A_{i-1}$, and $M\subseteq Y$. To see this one can apply Menger’s theorem between $\bigcup_{0\leqslant j\leqslant i-2}A_{j}$ and the connected component of $\bigcup_{j\geqslant i}A_{j}$ that contains $M$. For each $x\in X\cap Y$, let $\mathcal{H}_{x}=\{H\in L_{i-1}:x\in V(H)\}$. So, $\mathcal{H}_{x}$ is a clique in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, each $H\in I_{\mathcal{H}}(M)$ is at distance at most $2$ from $\bigcup_{x\in X\cap Y}\mathcal{H}_{x}$ in $\textbf{RIG}(G,\mathcal{H})$. Since $|X\cap Y|\leqslant 2$ and $\textbf{RIG}(G,\mathcal{H})[I_{\mathcal{H}}(M)]$ is connected, it now follows that $I_{\mathcal{H}}(M)$ has weak diameter at most $(2+1+2)+1+(2+1+2)=11$ in $\textbf{RIG}(G,\mathcal{H})$.
Now, we show for every $H\in\mathcal{H}$ that $I_{\mathcal{M}}(H)$ has weak diameter at most 9 in $\textbf{IM}(G,\mathcal{M})$. This is similar to the previous argument. Let $H\in\mathcal{H}$. Then for some $i\geqslant 0$, we have that $H\in L_{i}$. We may assume that $i\geqslant 3$, since otherwise we clearly have that $I_{\mathcal{M}}(H)$ has weak diameter at most $6$ in $\textbf{IM}(G,\mathcal{M})$. Similarly to before, by applying Menger’s theorem between $\bigcup_{0\leqslant j\leqslant i-3}A_{j}$ and the connected component of $\bigcup_{j\geqslant i-1}A_{j}$ that contains $H$, there exists a $2$-separation $(X,Y)$ of $G$ with $A_{i-3}\subseteq X$, $X\cap Y\subseteq A_{i-2}$, and $H\subseteq Y$. For each $x\in X\cap Y$, choose some $M_{x}\in\mathcal{M}_{i-2}$ with $x\in M_{x}$. Then in $\textbf{IM}(G,\mathcal{M})$, each vertex of $I_{\mathcal{M}}(H)$ is at distance at most 2 from some vertex of $\bigcup_{x\in X\cap Y}M_{x}$. Since $|X\cap Y|\leqslant 2$ and $\textbf{IM}(G,\mathcal{M})[I_{\mathcal{M}}(H)]$ is connected, it now follows that $I_{\mathcal{M}}(H)$ has weak diameter at most $(2+2)+1+(2+2)=9$ in $\textbf{IM}(G,\mathcal{M})$.
Hence $(G,\mathcal{M})$ is indeed a $(11,9)$-impression of $(G,\mathcal{H})$, as desired. ∎
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets in $G$, each of which contains a vertex incident to the outer face of $G$. For a collection disjoint connected sets $\mathcal{B}$ of $G$, we say that $(G,\mathcal{B})$ is a partial $(x,\infty)$-impression of $(G,\mathcal{H})$ if for every $B\in\mathcal{B}$, $I_{\mathcal{H}}(B)$ has weak diameter at most $x$ in $\textbf{RIG}(G,\mathcal{H})$.
We aim to find a $(x,y)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ in which each $M\in\mathcal{M}$ contains a vertex incident to the outer face of $G$. In light of Lemma˜3.2, we can find such a $(x,y)$-impression by first finding such a $(x^{\prime},\infty)$-impression and then refining this impression. The following theorem comes close to giving us such a $(x^{\prime},\infty)$-impression (not every $M\in\mathcal{M}$ contains a vertex incident to the outer face of $G$, but they are at distance at most $1$ from such a $M^{\prime}\in\mathcal{M}$ in $\textbf{IM}(G,\mathcal{M})$). The exact technical statement of this theorem is tailored to allow for an inductive argument on extending partial $(x^{\prime},\infty)$-impressions.
Theorem 3.3.
Let $G$ be a finite planar graph and $\mathcal{H}$ be a spanning collection of connected vertex sets in $G$, each of which contains a vertex incident to the outer face of $G$. Let $(G,\mathcal{B})$ be a partial $(30,\infty)$-impression of $(G,\mathcal{H})$ in which for every $B\in\mathcal{B}$, either $B$ contains a vertex incident to the outer face of $G$ and $I_{\mathcal{H}}(B)$ has weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$, or $B$ touches some $B^{\prime}\in\mathcal{B}$ where $B^{\prime}$ contains a vertex incident to the outer face of $G$, and for every maximal connected vertex set $K$ of $G\backslash(\bigcup_{B\in\mathcal{B}}B)$, either
  1. 1.
    $I_{\mathcal{H}}(K)$ has weak diameter at most $10$ in $\textbf{RIG}(G,\mathcal{H})$ and contains a vertex incident to the outer face of $G$, or $I_{\mathcal{H}}(K)$ has weak diameter at most $30$ in $\textbf{RIG}(G,\mathcal{H})$ and touches some $B\in\mathcal{B}$ that does contain a vertex incident incident to the outer face of $G$, or
  2. 2.
    $K$ contains a vertex incident to the outer face of $G$ and touches no $B\in\mathcal{B}$, or
  3. 3.
    $K$ contains a vertex incident to the outer face of $G$ and touches a unique $B\in\mathcal{B}$, or
  4. 4.
    $K$ contains a vertex incident to the outer face of $G$ and there exists distinct $B_{1},B_{2}\in\mathcal{B}$ satisfying the following
    • $K$ touches $B_{1}$ and $B_{2}$ and no other $\{B$$\in\mathcal{B}\backslash\{B_{1},B_{2}\}$,
    • $B_{1}$ and $B_{2}$ both contain a vertex incident to the outer face of $G$, and
    • $B_{1}$ and $B_{2}$ both contain a vertex incident to a common face $F$ distinct from the outer-face.
Then there is a $(30,\infty)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ such that for every $M\in\mathcal{M}$, either $M$ contains a vertex incident to the outer face of $G$ and $I_{\mathcal{H}}(B)$ has weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$, or $M$ touches some $M^{\prime}\in\mathcal{M}$ where $M^{\prime}$ contains a vertex incident to the outer face of $G$.
Proof.
We shall argue inductively on $|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$. Clearly the theorem holds if we have $V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)=\emptyset$. So, we may assume that there is some connected component of $G\backslash(\bigcup_{B\in\mathcal{B}}B)$, and we let $K$ be its vertex set.
Suppose first that $K$ satisfies Item˜1. Let $B^{*}=K$ and $\mathcal{B}^{*}=\mathcal{B}\cup\{B^{*}\}$. Clearly, $(G,\mathcal{B}^{*})$ satisfies the inductive hypothesis and $|V(G)\backslash(\bigcup_{B\in\mathcal{B}^{*}}B)|<|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$, so the theorem follows.
Now, suppose that $K$ satisfies Item˜2. Then simply choose some $b\in K$ incident to the outer face of $G$, let $B^{*}=\{b\}$ and $\mathcal{B}^{*}=\mathcal{B}\cup\{B^{*}\}$. Clearly, $(G,\mathcal{B}^{*})$ satisfies the inductive hypothesis and $|V(G)\backslash(\bigcup_{B\in\mathcal{B}^{*}}B)|<|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$, so the theorem follows.
[…]
Lastly, we can now assume that $K$ satisfies Item˜4. Let $B_{1},B_{2},F$ be as in Item˜4. Note that the intersection of $K$ with the vertices of the outer face of $G$ induces a connected graph since $B_{1}$ and $B_{2}$ both contain a vertex on the outer face of $G$ and the face $F$ (which is distinct from the outer face). Let $Q$ be the vertex set of this connected induced subgraph. Let $p_{1}p_{2}\ldots p_{t}$ be the walk in $G[Q]$ along the outer face of $G$ (and containing all of) $Q$, where $p_{1}$ touches $B_{1}$ and $p_{t}$ touches $B_{2}$. Choose $1\leqslant\ell\leqslant t$ maximum such that that there exists a path with vertex set $L$ between $p_{\ell}$ and $N_{G}(B_{1})$ such that $L$ is contained in $K\cap(H_{\ell}\cup H^{\prime}_{\ell})$ for some intersecting (but not necessarily distinct) $H_{\ell},H^{\prime}_{\ell}\in\mathcal{H}$. Similarly, choose $1\leqslant r\leqslant t$ minimum such that that there exists a path with vertex set $R$ between $p_{r}$ and $N_{G}(B_{2})$ such that $R$ is contained in $K\cap(H_{r}\cup H^{\prime}_{r})$ for some intersecting (but not necessarily distinct) $H_{r},H^{\prime}_{r}\in\mathcal{H}$. Note that such $\ell,r$ exist since $K$ touches exactly $B_{1},B_{2}$ from $\mathcal{B}$, and $K$ (and every $H\in\mathcal{H}$) contains a vertex incident to the outer face of $G$. Let $L^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(L)}H\right)\right]$ containing $L$, and similarly, let $R^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(R)}H\right)\right]$ containing $R$. Clearly $I_{\mathcal{H}}(L^{*})$ and $I_{\mathcal{H}}(R^{*})$ both have weak diameter at most 5 in $\textbf{RIG}(G,\mathcal{H})$. See Figure˜3 for an illustration (in a case where $L^{*}$ and $R^{*}$ do not intersect). We split into two cases depending on whether or not $L^{*}$ and $R^{*}$ intersect.
Figure 3: An illustration of the constructed connected sets $L^{*}$ and RjRj (coloured purple) in a case where they do not intersect. Both $B_{1}$ and $B_{2}$ are coloured blue. The connected sets in $\mathcal{H}$ are coloured red or orange, with the orange ones being $H_{\ell},H_{\ell}^{\prime},H_{r},H_{r}^{\prime}$. The two orange paths indicate $L$ and $R$.
[…]
We may now assume that $L^{*}$ and $R^{*}$ do not intersect. This in particular implies that $\ell<r$. Let $\mathcal{B}^{*}=\mathcal{B}\cup\{L^{*},R^{*}\}$. Consider the vertex set $K^{\prime}$ of a connected component of $G[K]\backslash(L^{*}\cup R^{*})$. Suppose that $K^{\prime}$ does not contain a vertex incident to the outer face of $G$. Then $K$ touches at least one of $B_{1},B_{2},L^{*},R^{*}$, and for every $H\in I_{\mathcal{H}}(K^{\prime})$, we have that $H$ intersects at least one of $B_{1},B_{2},L^{*},R^{*}$. Since $I_{\mathcal{H}}(B_{1}),I_{\mathcal{H}}(B_{2})$ both have weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$ and $I_{\mathcal{H}}(L^{*}),I_{\mathcal{H}}(R^{*})$ both have weak diameter at most 5 in $\textbf{RIG}(G,\mathcal{H})$, it follows that $I_{\mathcal{H}}(K^{\prime})$ has weak diameter at most 30 in $\textbf{RIG}(G,\mathcal{H})$. Thus in this case, $K^{\prime}$ satisfies Item˜1. So, we may now assume that $K^{\prime}$ does contain a vertex incident to the outer face of $G$. If $K^{\prime}$ does not touch both $L^{*}$ and $R^{*}$, then as before, $K^{\prime}$ satisfies either Item˜3 or Item˜4. So we assume now that $K^{\prime}$ touches both $L^{*}$ and $R^{*}$. Showing that $K^{\prime}$ satisfies Item˜4 will complete the proof.
Let $Q^{\prime}$ be the vertices $p_{i}$ of the walk $Q$ with $\ell<i<r$. Note that $Q^{\prime}$ is non-empty since $L^{*}$ and $R^{*}$ do not intersect (and $L^{*}$ contains both $p_{i}$ and $p_{i+1}$). Let $\mathcal{H}_{Q^{\prime}}=I_{\mathcal{H}}(Q^{\prime})$. Observe that by our choice of $L$ and $R$, there is no intersecting pair $H_{1}\in\mathcal{H}_{Q^{\prime}}$, $H_{2}\in\mathcal{H}$ such that there exists a path in $G[K\cap(H_{1}\cup H_{2})]$ between $Q^{\prime}$ and $N_{G}(B_{1}\cup B_{2})$. By our choice of $L^{*}$ and $R^{*}$, we therefore have that $I_{\mathcal{H}}(K^{\prime})\subseteq\mathcal{H}_{Q^{\prime}}$. In particular, $K^{\prime}$ does not touch $B_{1}$ or $B_{2}$. It now follows that there is some face $F$ of $G$ distinct from the outer face containing a vertex of $L^{*}$ and a vertex of $R^{*}$. Thus $K^{\prime}$ satisfies Item˜4.
So, $(G,\mathcal{B}^{*})$ satisfies the inductive hypothesis and $|V(G)\backslash(\bigcup_{B\in\mathcal{B}^{*}}B)|<|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$, so the theorem follows. ∎
Now we quickly refine the previous theorem so that every graph in $\mathcal{M}$ contains a vertex incident to the outer face of $G$.
Lemma 3.4.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$, each of which contains a vertex incident to the outer face of $G$. Then there is a $(70,\infty)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ such that for every $\in\mathcal{M}, MM$ contains a vertex incident to the outer face of $G$.
Proof.
By Theorem˜3.3 (applied with $\mathcal{B}=\emptyset$), there is a $(30,\infty)$-impression $(G,\mathcal{M}^{\prime})$ of $(G,\mathcal{H})$ such that for every $M\in\mathcal{M}^{\prime}$, either $M$ contains a vertex incident to the outer-face of $G$ and $I_{\mathcal{H}}(B)$ has weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$, or $M$ touches some $M^{\prime}\in\mathcal{M}^{\prime}$ where $M^{\prime}$ contains a vertex incident to the outer face of $G$.
Let $\mathcal{M}_{1}$ be the elements $M$ of $\mathcal{M}^{\prime}$ that contain a vertex incident to the outer face of $G$, and let $\mathcal{M}_{2}$ be the elements $M$ of $\mathcal{M}^{\prime}$ that do not contain a vertex incident to the outer face of $G$. Then $\mathcal{M}^{\prime}=\mathcal{M}_{1}\cup\mathcal{M}_{2}$. Choose some function $f:\mathcal{M}_{2}\to\mathcal{M}_{1}$ such that $M$ touches $f(M)$ for every $M\in\mathcal{M}_{2}$. Let $\mathcal{M}=\{M\cup f^{-1}(M):M\in\mathcal{M}_{1})\}$. Then for every $M\in\mathcal{M}$, we have that $I_{\mathcal{H}}(M)$ has weak diameter at most $30+10+30=70$ in $G$. Then $(G,\mathcal{M})$ is a $(70,\infty)$-impression of $(G,\mathcal{H})$ such that for every $\in\mathcal{M}, MM$ contains a vertex incident to the outer face of $G$, as desired. ∎
We can now quickly prove Theorem˜3.1 using Lemma˜3.4 and Lemma˜3.2. We remark that in our proof of Theorem˜2.5, it actually turns out to be more convenient to use Lemma˜3.4 and Lemma˜3.2 separately as below instead of Theorem˜3.1. This is essentially because we end up needing to apply Lemma˜3.2 another time later anyway. However for the sake of completeness for outerstring graphs, we still complete the proof of Theorem˜3.1 here.
Proof of Theorem˜3.1..
By Lemma˜3.4, there is a $(70,\infty)$-impression $(G,\mathcal{M}^{\prime})$ of $(G,\mathcal{H})$ such that for every $\in\mathcal{M}^{\prime}, MM$ contains a vertex incident to the outer face of $G$. Let $G^{\prime}=\textbf{IM}(G,\mathcal{M}^{\prime})$. Then $G^{\prime}$ is outerplanar. For each $H\in\mathcal{H}$, let $H^{\prime}$ be the connected vertex set of $G^{\prime}$ that is exactly the union of all $M^{\prime}\in\mathcal{M}^{\prime}$ such that $M^{\prime}$ intersects $H$. Let $\mathcal{H}^{\prime}=\{H^{\prime}:H\in\mathcal{H}\}$.
By Lemma˜3.2, there is a $(11,9)$-impression $(G^{\prime},\mathcal{M}^{*})$ of $(G^{\prime},\mathcal{H}^{\prime})$. Let $\mathcal{M}=\{\bigcup_{M^{\prime}\in M^{*}}M^{\prime}:M^{*}\in\mathcal{M}\}$. Clearly for every $H\in\mathcal{H}$, $I_{\mathcal{M}}(H)$ has weak diameter at most $9$ in $\textbf{IM}(G,\mathcal{M})$. Also, for every $M\in\mathcal{M}$, we have that $I_{\mathcal{H}}(M)$ has weak diameter at most $11\cdot 70=770$ in $\textbf{RIG}(G,\mathcal{H})$.
Hence, $(G,\mathcal{M})$ is a $(770,9)$-impression of $(G,\mathcal{H})$, as desired. ∎
4 Encasings
The key to carrying out our inductive proof of Theorem˜2.5 in Section˜5 will be that we can find what we shall call a $(a,b,c,d)$-encasing. The aim of this section is to find these encasings. We show that we can find $(a,b,c,d)$-encasing by first finding a $(a,\infty,\infty,\infty)$-encasing, then refining it to a $(a,\infty,c,\infty)$-encasing, then further refining it to a $(a,\infty,c,d)$-encasing, and then finally refining it to a $(a,b,c,d)$-encasing. To ease into things, we delay the full definition of an $(a,b,c,d)$-encasing for now, and start simply with $(a,\infty,\infty,\infty)$-encasings.
Let $G$ be a finite planar graph and let $\mathcal{H}$ be a spanning collection of connected vertex sets in $G$. Let $\mathcal{B}$ be a collection of disjoint connected vertex sets in $G$. We say that $\mathcal{B}$ encircles $(G,\mathcal{H})$ if every $B\in\mathcal{B}$ contains a vertex incident to the outer face of $G$, and every $H\in\mathcal{H}$ that contains a vertex incident to the outer face of $G$ is contained in $\bigcup_{B\in\mathcal{B}}B$. We say that $\mathcal{B}$ is a $(a,\infty,\infty,\infty)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$.
We begin by finding a $(a,\infty,\infty,\infty)$-encasing. The lemma has an extra condition on the vertices contained in $\bigcup_{B\in\mathcal{B}}B$ that we will need later.
Lemma 4.1.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(210,\infty,\infty,\infty)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ with $\bigcup_{B\in\mathcal{B}}B=\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$, where $\mathcal{H}_{0}$ are the elements of $\mathcal{H}$ that contain a vertex incident to the outer face of $G$ and $\mathcal{H}_{1}$ are the elements of $\mathcal{H}\backslash\mathcal{H}_{0}$ that intersect a element of $\mathcal{H}_{0}$.
Proof.
Let $G^{*}$ be the subgraph of $G$ with edges $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}E(G[H])$ and no isolated vertices. Note that $G^{*}$ has the same outer face as $G$. For each $H\in\mathcal{H}_{1}$, choose some $H^{\prime}\in\mathcal{H}_{0}$ that intersects $H$ and let $H^{*}=H\cup H^{\prime}$. Let $\mathcal{H}^{*}=\mathcal{H}_{0}\cup\{H^{*}:H\in\mathcal{H}_{1}\}$. Then every vertex on the outer face of $G$ is contained in some element of $\mathcal{H}^{*}$. By Lemma˜3.4, there is a $(70,\infty)$-impression $(G^{*},\mathcal{B})$ of $(G^{*},\mathcal{H}^{*})$ such that every vertex of $G^{*}$ incident to the outer face of $G^{*}$ is contained in some $B\in\mathcal{B}$. In particular, every vertex of $G$ incident to the outer face of $G$ is contained in some $B\in\mathcal{B}$. Furthermore, observe that as $I_{\mathcal{H}^{*}}(B)$ has weak diameter at most 70 in $\textbf{RIG}(G^{*},\mathcal{H}^{*})$ for each $B\in\mathcal{B}$, we also have that $I_{\mathcal{H}^{*}}(B)$ has weak diameter at most $3\cdot 70=210$ in $\textbf{RIG}(G,\mathcal{H})$ for each $B\in\mathcal{B}$. ∎
Let $G$ be a finite planar graph, $\mathcal{H}$ a spanning collection of connected vertex sets of $G$, and let $\mathcal{B}$ be a collection of disjoint connected sets of $G$ that encircles $(G,\mathcal{H})$. Let $\mathcal{F}$ be the collection of non-empty vertex sets of $G$ that are maximal with respect to being contained in the interior of a face of $G[\bigcup_{B\in\mathcal{B}}B]$. We call these facial vertex sets. We remark that these facial vertex sets are not necessarily connected sets of $G$. Note that $\mathcal{B}\cup\mathcal{F}$ is a partition of $V(G)$. We denote by $G(\mathcal{B})$ the graph obtained from $\textbf{IM}(G,\mathcal{B})$ by for every facial vertex set $\mathcal{F}$ of $G[\bigcup_{B\in\mathcal{B}}B]$, we add $F$ as a vertex and make it adjacent to exactly the elements $B\in\mathcal{B}$ that contain a vertex incident to the face of $G[\bigcup_{B\in\mathcal{B}}B]$ that contains $F$ (we remark that $B$ and $F$ need not touch).
We say that some $H\in\mathcal{H}$ is $d$-caged by $\mathcal{B}$ if for every $u,v\in H$, there exists some sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant d$ such that
  • each $A_{i}$ is either a single element of $\mathcal{B}$ or a single element of $\mathcal{F}$ or the union of a subcollection $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$,
  • $u\in A_{0}$,
  • $v\in A_{t}$, and
  • for every $1\leqslant i<t$, we have that either $H$ intersects both $A_{i-1}$ and $A_{i}$, or neither $A_{i-1}$ or $A_{i}$ is an element of $\mathcal{F}$ and they touch.
See Figure˜4 for an illustration of some $7$-caged $H$. We say that $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H)}$ if every $H\in\mathcal{H}$ is $d$-caged by $\mathcal{B}$.
Figure 4: An illustration of some $H\in\mathcal{H}$ (coloured red) that is 7-caged by $\mathcal{B}$ (coloured blue). For $u,v\in H$, the vertex sets $A_{0},\ldots,A_{7}$ (coloured green) are as in the definition of being 7-caged. As an example illustrating the end of the proof of Lemma˜4.2, we would have that $A_{0}=A_{u}$, $A_{1}\in\mathcal{F}_{u}$, $A_{2},A_{5}\in\mathcal{A}_{u,v}$, $\{A_{3},A_{4}\}=\mathcal{V}_{(X,Y)}$, $A_{6}\in\mathcal{F}_{v}$, and $A_{7}=A_{v}$.
The motivation behind the notion of a $d$-cage is that it will allow us to “cut up” connected sets $H$ in such a way that each part is in some sense still close via the encircling $\mathcal{B}$. Being able to do this is crucial to the inductive argument of Theorem˜2.5.
We say that $\mathcal{B}$ is a $(a,\infty,\infty,d)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$, and
  • $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H})$.
In other words, $\mathcal{B}$ is a $(a,\infty,\infty,d)$-encasing of $(G,\mathcal{H})$ if it is both a $(a,\infty,\infty,\infty)$-encasing and a $d$-cage of $(G,\mathcal{H})$.
We will use our $(a,\infty,\infty,\infty)$-encasing to get a $(a,\infty,\infty,d)$-encasing. However, we again need this $(a,\infty,\infty,d)$-encasing to have some stronger structure on the vertices of $G$ that $\bigcup_{B\in\mathcal{B}}B$ contains. Let $\mathcal{H}_{0}$ be the elements of $\mathcal{H}$ that contain a vertex incident to the outer face of $G$, let $\mathcal{H}_{1}$ be the elements of $\mathcal{H}\backslash\mathcal{H}_{0}$ that intersect a element of $\mathcal{H}_{0}$, and similarly, let $\mathcal{H}_{2}$ be the elements of $\mathcal{H}\backslash(\mathcal{H}_{0}\cup\mathcal{H}_{1})$ that intersect a element of $\mathcal{H}_{1}$. We say that a $(a,\infty,\infty,d)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ is enforced if $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\subseteq\bigcup_{B\in\mathcal{B}}B\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2}}H$, and for every $v\in\left(\bigcup_{B\in\mathcal{B}}B\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, there is some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $\bigcup_{B\in\mathcal{B}}B$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$.
Lemma 4.2.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is an enforced $(840,\infty,\infty,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Proof.
We define $\mathcal{H}_{0},\mathcal{H}_{1},\mathcal{H}_{2}$ as above in the definition of an enforced $(a,\infty,\infty,d)$-encasing of $(G,\mathcal{H})$. By Lemma˜4.1, there is a $(210,\infty,\infty,\infty)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ with $\bigcup_{B\in\mathcal{B}}B=\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$. We will adjust this encasing to obtain the desired enforced $(840,\infty,\infty,7)$-encasing.
Let $G^{*}=G(\mathcal{B})$ as above, where $\mathcal{F}$ are the additional facial vertex sets. Choose some $B_{0}\in\mathcal{B}$. We say that a separation $(X,Y)$ of $G^{*}$ is 1-clean if $B_{0}\in X$, $X\cup Y=V(G^{*})$, $|X\cap Y|=1$, and subject to this, there is no other such $(X^{\prime},Y^{\prime})$ with $X^{\prime}\cap Y^{\prime}=X\cap Y$ and $|Y^{\prime}|>|Y|$. A separation $(X,Y)$ of $G^{*}$ is 2-clean if $B_{0}\in X$, $X\cup Y=V(G^{*})$, $X\cap Y$ has two elements and they are adjacent in $G^{*}$, and there is no 1-clean separation $(X^{*},Y^{*})$ with $X^{*}\cap Y^{*}\subset X\cap Y$ and $Y^{*}\subseteq Y$, and there is no other such separation $(X^{\prime},Y^{\prime})$ of $G^{*}$ with $X^{\prime}\cap Y^{\prime}=X\cap Y$ and $|Y^{\prime}|>|Y|$. By abuse of notation, we shall also call $(\{B_{0}\},V(G^{*}))$ a 1-clean separation. A separation $(X,Y)$ of $G^{*}$ is clean if it is 1-clean or 2-clean. Observe that if $(X,Y)$ is a clean separation, then $X\cap Y\subseteq\mathcal{B}$.
[…]
Figure 5: A choice of $V^{1}_{(X,Y)},V^{2}_{(X,Y)}$ (coloured purple) as in the proof of Lemma˜4.2. Elements of $\mathcal{B}$ are coloured blue and elements of $\mathcal{H}$ are coloured red.
For a clean separation $(X,Y)$ of $G^{*}$, we let $H_{(X,Y)}$ be the vertices $v$ of $G$ contained in some vertex set of $\mathcal{F}\cap Y$ such that there exists some $H\in\mathcal{H}_{2}$ containing both $v$ and intersecting $\bigcup_{A\in X\cap Y}A$. If $(X,Y)$ is 1-clean, then we let $V_{(X,Y)}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$. Note that in this case, $V_{(X,Y)}$ is a connected set of $G$ and $I_{\mathcal{H}}(V_{X,Y})$ has weak diameter at most $3\cdot 210=630$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in V_{(X,Y)}\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=V_{(X,Y)}\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$ (or more precisely, the unique vertex set of $X\cap Y$). Otherwise, if $(X,Y)$ is 2-clean with $X\cap Y=\{B^{1},B^{2}\}$, then we choose disjoint $V_{(X,Y)}^{1}$, $V_{(X,Y)}^{2}$ such that $B^{1}\subseteq V_{(X,Y)}^{1}$, $B^{2}\subseteq V_{(X,Y)}^{2}$, $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$, and both $V_{(X,Y)}^{1}$ and $V_{(X,Y)}^{2}$ are connected sets of $G$ (see Figure˜5 for an illustration). Then $I_{\mathcal{H}}(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})$, and in particular both of $I_{\mathcal{H}}(V_{(X,Y)}^{1})$ and $I_{\mathcal{H}}(V_{(X,Y)}^{2})$ all have weak diameter at most $4\cdot 210=840$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}$, and intersects $B^{1}\cup B^{2}\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$.
Let $\mathcal{B}^{\prime}$ be the elements of $\mathcal{B}$ that do not intersect $V_{(X,Y)}$ for any $(X,Y)\in\mathcal{T}_{1}$ and do not intersect $V_{(X,Y)}^{1}$ or $V_{(X,Y)}^{2}$ for any $(X,Y)\in\mathcal{T}_{2}$. At last, let $\mathcal{B}^{*}=\mathcal{B}^{\prime}\cup\{V_{(X,Y)}:(X,Y)\in\mathcal{T}_{1}\}\cup\{V_{(X,Y)}^{1},V_{(X,Y)}^{2}:(X,Y)\in\mathcal{T}_{2}\}$.
Clearly, we have that $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\subseteq\bigcup_{B\in\mathcal{B}^{*}}B\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2}}H$, and for every $B\in\mathcal{B}^{*}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $840$ in $\textbf{RIG}(G,\mathcal{H})$. We also have that for every $v\in\left(\bigcup_{B\in\mathcal{B}^{*}}B\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, there is some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $\bigcup_{B\in\mathcal{B}^{*}}B$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$. To show that $\mathcal{B}^{*}$ provides the desired $(840,\infty,\infty,7)$-encasing of $(G,\mathcal{H})$, it now just remains to show that $\mathcal{B}^{*}$ is a $7$-cage of $(G,\mathcal{H})$.
If $H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}$, then $H$ is 0-caged by $\mathcal{B}^{*}$ since $H\subseteq\bigcup_{B\in\mathcal{B}^{*}}B$. If $H\in\mathcal{H}\backslash(\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2})$, then $H$ lies in some face of $G(\mathcal{B})$. In particular, $H$ intersects at most two elements of $\mathcal{B}^{*}$ and furthermore if $H$ does intersect two elements of $\mathcal{B^{*}}$ then they are touching, so it follows that $H$ is 3-caged by $\mathcal{B}^{*}$.
It remains to consider $H\in\mathcal{H}_{2}$. To do so, we must first introduce certain subsets $\mathcal{B}^{*}_{(X,Y)}$ and $\mathcal{F}^{*}_{(X,Y)}$ of $V(G(\mathcal{B}^{*}))$ according to clean separations $(X,Y)$ of $G^{*}$. Fix some $H\in\mathcal{H}_{2}$. For each 1-clean separation $(X,Y)\in\mathcal{T}_{1}$, let $\mathcal{V}_{(X,Y)}=\{V_{(X,Y)}\}$, and for each 2-clean separation $(X,Y)\in\mathcal{T}_{2}$, let $\mathcal{V}_{(X,Y)}=\{V^{1}_{(X,Y)},V_{(X,Y)}^{2}\}$. Let $\mathcal{F}^{*}$ be the facial vertex sets of $G(\mathcal{B}^{*})$. For each clean separation $(X,Y)\in\mathcal{T}$, let $\mathcal{F}^{*}_{(X,Y)}$ be the elements $F$ of $\mathcal{F}^{*}$ adjacent in $G(\mathcal{B}^{*})$ to a element of $\mathcal{V}_{(X,Y)}$. Now, let $\mathcal{B}^{*}_{(X,Y)}=N_{G(\mathcal{B}^{*})}(\mathcal{F}^{*}_{(X,Y)})$.
Consider some $H\in\mathcal{H}_{2}$. By our choice of $\mathcal{B}^{*}$, we have that there exists some clean separation $(X,Y)\in\mathcal{T}$ such that $H$ is entirely contained in $\bigcup_{B\in\mathcal{B}^{*}_{(X,Y)}}B\cup\bigcup_{F\in\mathcal{F}^{*}_{(X,Y)}}F$. Observe that $G(\mathcal{B}^{*})[\mathcal{B}^{*}_{(X,Y)}\cup\mathcal{F}^{*}_{(X,Y)}]$ is a planar graph in which the vertices $\mathcal{B}^{*}_{(X,Y)}$ are incident to the outer face and every $F\in\mathcal{F}^{*}_{(X,Y)}$ is adjacent to some $V\in\mathcal{V}_{(X,Y)}$, where $|\mathcal{V}_{(X,Y)}|\leqslant 2$ and if $|\mathcal{V}_{(X,Y)}|=2$, then the two vertices of $\mathcal{V}_{(X,Y)}$ are adjacent.
Consider some $u,v\in H$. Let $H_{u}$ be a subset of $H$ that contains $u$ and is maximal with respect to being a connected set of $G$ and being contained in $G[\bigcup_{B\in\mathcal{B}}B]$ (or equivalently $G[\bigcup_{B\in\mathcal{B}^{*}}B]$). Let $A_{u}$ be the union of all $B\in\mathcal{B}^{*}$ that intersect $H_{u}$. Let $\mathcal{F}_{u}$ be elements $F$ of $\mathcal{F}^{*}$ such that $H$ intersects both $F$ and $A_{u}$. We define $A_{v}$ and $\mathcal{F}_{v}$ similarly. Crucially, note that if $H$ intersects distinct $F_{1},F_{2}\in\mathcal{F}^{*}$, then $H$ intersects some $V\in\mathcal{B}$ incident to the face of $G(\mathcal{B}^{*})$ corresponding to $F_{1}$ that is also either contained in $\mathcal{V}_{(X,Y)}$ or adjacent to some element of $\mathcal{V}_{(X,Y)}$ in $G(\mathcal{B}^{*})$. Let $\mathcal{A}_{u,v}$ be the collection of unions of subcollections $\mathcal{B}^{\prime}\subseteq\mathcal{B}^{*}\backslash\mathcal{V}_{(X,Y)}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$, and some $B\in\mathcal{B^{\prime}}$ touches some $V\in\mathcal{V}_{(X,Y)}$. To see that $H$ is 7-caged by $\mathcal{B}^{*}$, it is now straight forward to choose the required sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant 7$ from the collection $\{A_{u}\}\cup\mathcal{F}_{u}\cup(\mathcal{A}_{u,v}\cup\mathcal{V}_{(X,Y)})\cup\mathcal{F}_{v}\cup\{A_{v}\}$ (here we take $A_{0}=A_{u}$, $A_{t}=A_{v}$, at most one $A_{i}$ is contained in $\mathcal{F}_{u}$ (this would be $A_{1}$ if so), at most one is contained in $\mathcal{F}_{v}$ (this would be $A_{t-1}$ if so), at most two are contained in $\mathcal{A}_{u,v}$, and at most two more are contained in $\mathcal{V}_{(X,Y)}$). For an illustrative example, see Figure˜4.
So, $\mathcal{B}^{*}$ is indeed a 7-cage of $(G,\mathcal{H})$, and thus $\mathcal{B}^{*}$ is an enforced $(840,\infty,\infty,7)$-encasing of $(G,\mathcal{H})$, as desired. ∎
Let $G$ be a finite planar graph, $\mathcal{H}$ a spanning collection of connected vertex sets of $G$, and let $\mathcal{B}$ be a collection of disjoint connected sets of $G$ such that $\mathcal{B}$ encircles $(G,\mathcal{H})$. Let $\mathcal{R}$ be the interiors of the internal faces of $G[\bigcup_{B\in\mathcal{B}}B]$. For $u\in\bigcup_{B\in\mathcal{B}}B$ and $H\in\mathcal{H}$, we say that $u$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ if either $u\in H$ or $u$ is contained in one of the bounded regions of $\mathbb{R}^{2}\backslash(G[H]\cup\bigcup_{R\in\mathcal{R}}R)$. We say that $U\subseteq\bigcup_{B\in\mathcal{B}}B$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ if each $u\in U$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$. Let $F$ be the vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$. We say that a vertex $v\in F$ is $\mathcal{B}$-encroached by some $H\in\mathcal{H}$ with $v\notin H$ if there exists some connected set $U\subseteq\bigcup_{B\in\mathcal{B}}B$ of $G$ and edges $u_{1}v_{1},\ldots,u_{r}v_{r}$ between $F$ and $\bigcup_{B\in\mathcal{B}}B$ that appear consecutively along the boundary of the face of $G[\bigcup_{B\in\mathcal{B}}B]$ corresponding to $F$, such that
  • $U$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$,
  • for each $1\leqslant i<r$, there is a facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$,
  • $u_{1},v_{1},u_{r},v_{r}\in H$, and
  • $v\in\{v_{1},\ldots,v_{r}\}$.
See Figure˜6 for an illustrative example. We further say that some $H^{\prime}\in\mathcal{H}$ is $\mathcal{B}$-encroached by some $H\in\mathcal{H}$ if there is some set $F$ of vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$ and some vertex $v\in H^{\prime}\cap F$ such that $v$ is $\mathcal{B}$-encroached by $H$.
Figure 6: An encircling $\mathcal{B}$ (in blue) and some (red) $H\in\mathcal{H}$. The vertices of the facial vertex set $F$ that are encroached by $H$ are coloured orange.
Let $\mathcal{H}_{0}$ be the connected sets in $\mathcal{H}$ that contain a vertex incident to the outer face of $G$. We say that a collection of disjoint connected sets $\mathcal{B}$ of $G$ that encircles $(G,\mathcal{H})$ is a $c$-surround of $(G,\mathcal{H})$ if
  • for every pair $H_{u},H\in\mathcal{H}$ such that $H$ intersects but is not contained in $\bigcup_{B\in\mathcal{B}}B$, and $H_{u}$ contains a vertex $u\in\bigcup_{B\in\mathcal{B}}B$ that is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$, we have that $H_{u}$ is at distance at most c1c-1 from $H$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$, and
  • for every pair $H^{\prime},H\in\mathcal{H}$ such that $H$ encroaches $H^{\prime}$, we have that $H^{\prime}$ is at distance at most $c$ from $H$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$.
Note that since $\mathcal{B}$ encircles $(G,\mathcal{H})$, if $H$ encroaches $H^{\prime}$, then $H\not\in\mathcal{H}_{0}$.
The motivation behind the notion of a $c$-surround is that it will allow us to add some (not necessarily yet connected until more edges are added to $G$) sets $H^{\prime}$ to $\mathcal{H}$ such that $I_{\mathcal{H}}(H^{\prime})$ has bounded weak diameter in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ and $H^{\prime}$ is disjoint from $\bigcup_{B\in\mathcal{B}}B$. Such $H^{\prime}$ will contain for instance the vertices encroached by a given $H\in\mathcal{H}$ and will be crucial for the inductive argument of Theorem˜2.5 in which we remove the vertices of $\bigcup_{B\in\mathcal{B}}B$.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Let $\mathcal{B}$ be a collection of disjoint connected subgraphs of $G$. We say that $\mathcal{B}$ is a $(a,\infty,c,d)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$,
  • $\mathcal{B}$ is a $c$-surround of $(G,\mathcal{H})$, and
  • $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H})$.
The enforced encasing found in Lemma˜4.2 is actually a 4-surround (which was the purpose of the technical definition of enforced encasings).
Lemma 4.3.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(840,\infty,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Proof.
By Lemma˜4.2, there is an enforced $(840,\infty,\infty,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$. Let $\mathcal{H}_{0},\mathcal{H}_{1},\mathcal{H}_{2}$ be as in the definition of an enforced encasing, and furthermore, let $\mathcal{H}_{3}$ be the elements of $\mathcal{H}\backslash(\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2})$ that intersect an element of $\mathcal{H}_{2}$. We will show that $\mathcal{B}$ is in fact a $(840,\infty,4,7)$-encasing of $(G,\mathcal{H})$.
First, we show the first bullet in the definition of being a 4-surround. Consider some pair $H_{u},H^{*}\in\mathcal{H}$ such that $H^{*}$ intersects but is not contained in $\bigcup_{B\in\mathcal{B}}B$, and $H_{u}$ contains a vertex $u\in\bigcup_{B\in\mathcal{B}}B$ that is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$. Clearly $H^{*}\in\mathcal{H}_{2}\cup\mathcal{H}_{3}$, and therefore $u\not\in\bigcup_{H\in{\mathcal{H}_{0}}}H$. In particular, $H_{u},H^{*}\not\in\mathcal{H}_{0}$.
Suppose first that $u\in\left(\bigcup_{H\in\mathcal{H}_{1}}H\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}}H\right)$. Then, there is some $H_{u}^{\prime}\in\mathcal{H}_{1}$ that contains $u$ (possibly $H_{u}^{\prime}$ is not distinct from $H_{u}$). Observe that $H_{u}^{\prime}$ must intersect $H^{*}$ since $u$ is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$, $H_{u}\in\mathcal{H}_{1}$, and $H^{*}\in\mathcal{H}_{2}\cup\mathcal{H}_{3}$. Therefore, $H_{u}$ is at distance at most 2 from $H^{*}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$.
So, we may assume now that $u\in\left(\bigcup_{B\in\mathcal{B}}B\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$. Then, by the definition of an enforced encasing, there is some connected set $P_{u}\subseteq H_{u}^{\prime}\in\mathcal{H}_{2}$ of $G$ that contains $u$, is contained in $\bigcup_{B\in\mathcal{B}}B$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$. Choose $H_{u}^{\prime\prime}\in\mathcal{H}_{1}$ that intersects $P_{u}$ (and thus also $H_{u}^{\prime}$). Similarly to before, $H_{u}^{\prime\prime}\cup P_{u}$ must intersect $H^{*}$ since $u$ is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$, $u\in H_{u}^{\prime\prime}\cup P_{u}\subseteq\bigcup_{B\in\mathcal{B}}B$, $H_{u}^{\prime\prime}\in\mathcal{H}_{1}$, and $H^{*}\in\mathcal{H}_{2}\cup\mathcal{H}_{3}$. In particular, $H^{*}$ intersects at least one of $H_{u}^{\prime\prime}$ or $H_{u}^{\prime}$. Therefore, $H_{u}$ is at distance at most 3 from $H^{*}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$. Thus, $\mathcal{B}$ satisfies the first bullet of being a 4-surround.
Suppose that some $H^{\prime}\in\mathcal{H}$ is $\mathcal{B}$-encroached by some $H^{*}\in\mathcal{H}$. Clearly we have that $H^{\prime},H^{*}\not\in\mathcal{H}_{0}$. Note that $H^{*}$ intersects but is not contained in $\bigcup_{B\in\mathcal{B}}B$. By definition, there is some set $F$ of vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$ and some vertex $v\in H^{\prime}\cap F$ such that $v$ is $\mathcal{B}$-encroached by $H^{*}$. Let $u$ be a neighbour of $v$ in $\bigcup_{B\in\mathcal{B}}B$ that is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$, and let $H_{u}$ be some connected set in $\mathcal{H}$ that contains $\{u,v\}$ (possibly $H_{u}$ is not distinct from $H^{\prime}$). Then, $H^{\prime}$ is at distance at most $1+3=4$ from $H^{*}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$.
Thus, $\mathcal{B}$ is a 4-surround of $(G,\mathcal{H})$, and in particular, a $(840,\infty,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ as desired. ∎
We are finally ready to give the entire definition of a $(a,b,c,d)$-encasing. We say that $\mathcal{B}$ is a $(a,b,c,d)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$,
  • for every connected set $H^{*}\subseteq H\in\mathcal{H}$ of $G$ such that $H^{*}$ is contained in $\bigcup_{B\in\mathcal{B}}B$, we have that $I_{\mathcal{B}}(H^{*})$ has weak diameter at most $b$ in $\textbf{IM}(G,\mathcal{B})$,
  • $\mathcal{B}$ is a $c$-surround of $(G,\mathcal{H})$, and
  • $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H})$.
We can at last complete the refinement of the encasing with the use of Lemma˜3.2.
Lemma 4.4.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(9240,9,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Proof.
By Lemma˜4.3, there is a $(840,\infty,4,7)$-encasing $\mathcal{B}^{\prime}$ of $(G,\mathcal{H})$. Let $G^{\prime}=\textbf{IM}(G,\mathcal{B}^{\prime})$. Then $G^{\prime}$ is outerplanar. For every connected set $H^{*}\subseteq H\in\mathcal{H}$ of $G$ such that $H^{*}$ is contained in $\bigcup_{B\in\mathcal{B}^{\prime}}B$, let $H^{\prime}$ be the connected set of $G^{\prime}$ whose vertex set is exactly the $B\in\mathcal{B}^{\prime}$ such that $H^{*}$ intersects $B$. Let $\mathcal{H}^{\prime}$ be the collection of all of these connected vertex sets of $G^{\prime}$.
By Lemma˜3.2, there is a $(11,9)$-impression $(G^{\prime},\mathcal{M})$ of $(G^{\prime},\mathcal{B}^{\prime})$. Let $\mathcal{B}=\{\bigcup_{B\in M}B:M\in\mathcal{M}\}$. Clearly for every $H\in\mathcal{H}$, $I_{\mathcal{M}}(H)$ has weak diameter at most $9$ in $\textbf{IM}(G,\mathcal{M})$. Also, for every $M\in\mathcal{M}$, we have that $I_{\mathcal{H}}(M)$ has weak diameter at most $11\cdot 840=9240$ in $\textbf{RIG}(G,\mathcal{H})$.
Hence, $\mathcal{B}$ is a $(9240,9,4,7)$-encasing of $(G,\mathcal{H})$, as desired. ∎
5 String graphs
Lemma˜4.4 is our main technical tool for proving Theorem˜2.5. We still need to introduce the notion of fortifications in order to give the precise technical statement that will allow for the inductive proof using Lemma˜4.4.
[…]
Figure 7: A fortification $G^{\prime}$ of planar graph $G$ (which is the subgraph on the black edges. The arrows orthogonal to the additional red edges indicate their sides.
Let $G$ be a planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets $\mathcal{H}$ of $G$. Let $\mathcal{H}_{0}$ be the collection of connected sets $H\in\mathcal{H}$ of $G$ that contain a vertex incident to the outer face of $G$. we say that $(G^{\prime},\mathcal{H^{\prime}})$ is a $k$-fortification of $(G,\mathcal{H})$ if
  • $G^{\prime}$ is a fortification of $G$,
  • $\mathcal{H}^{\prime}$ spans $G^{\prime}$,
  • $\mathcal{H}^{\prime}$ can be obtained from $\mathcal{H}$ by adding connected sets $H^{\prime}$ of $G$ such that each $I_{\mathcal{H}}(H^{\prime})$ has weak diameter at most $k$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$, and
[…]
We remark that for the purposes of this paper, the last bullet could be weakened so that $H$ and $H^{\prime}$ just contain an end of $e$. However, this stronger version does not complicate the proof any further and we need this property for our applications to Riemannian surfaces [17].
We use the following technical statement to facilitate an inductive argument making use of Lemma˜4.4.
Theorem 5.1.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(9242,80)$-impression $(G^{\prime},\mathcal{M})$ of a $8$-fortification $(G^{\prime},\mathcal{H}^{\prime})$ of $(G,\mathcal{H})$ such that for every $M\in\mathcal{M}$ containing a vertex incident to the outer face of $G$, we have that $I_{\mathcal{H^{\prime}}}(M)$ has weak diameter at most $9240$ in $\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})$, and for every $H\in\mathcal{H}$ containing a vertex incident to the outer face of $G$, we have that $I_{\mathcal{M}}(H)$ has weak diameter at most $9$ in $\textbf{IM}(G^{\prime},\mathcal{M})$.
Proof.
We shall argue inductively on $|V(G)|$ (trivially the theorem holds for small $|V(G)|$). Let $\mathcal{H}_{0}$ be the collection of connected sets $H\in\mathcal{H}$ that contain a vertex incident to the outer face of $G$. By Lemma˜4.4, there is a $(9240,9,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Let $\mathcal{F}$ be the facial vertex sets of $G(\mathcal{B})$. Consider some $F\in\mathcal{F}$ and $H\in\mathcal{H}$ such that $H$ intersects both $\bigcup_{B\in\mathcal{B}}B$ and $F$. Let $R_{F,H}^{\prime}$ be the vertices of $F$ that are encroached by $H$, and let $R_{F,H}=R_{F,H}^{\prime}\cup(F\cap H)$. Consider some $H^{\prime}\in\mathcal{H}$ intersecting $R_{F,H}$ but not $H$. Then, $H^{\prime}$ intersects $R_{F,H}^{\prime}$. Since $\mathcal{B}$ is a $4$-surround of $(G,\mathcal{H})$, the distance between $H^{\prime}$ and $H$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ is at most 4. Therefore, $I_{\mathcal{H}}(R_{F,H})$ has weak diameter at most $2\cdot 4=8$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$. We wish to add edges to $G$ so that $R_{F,H}$ is a connected set.
Consider some $v\in R_{F,H}$. By definition of being encroached, there exists connected set $U\subseteq\bigcup_{B\in\mathcal{B}}B$ of $G$ and edges $u_{1}v_{1},\ldots,u_{r}v_{r}$ between $F$ and $\bigcup_{B\in\mathcal{B}}B$ that appear consecutively along the boundary of the face of $G[\bigcup_{B\in\mathcal{B}}B]$ corresponding to $F$, such that
  • $U$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$,
  • for each $1\leqslant i<r$, there is a facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$,
  • $u_{1},v_{1},u_{r},v_{r}\in H$, and
  • $v\in\{v_{1},\ldots,v_{r}\}$.
Clearly $\{v_{1},\ldots,v_{r}\}\subseteq R_{F,H}$. Let $E_{F,H}$ be the collection of such edges $v_{i}v_{i+1}$ (across each $v\in R_{F,H}$) that are not already an edge of $G$. For each such $v_{i}v_{i+1}\in E_{F,H}$, let $F_{v_{i}v_{i+1}}$ be the side of $v_{i}v_{i+1}$ drawn in the corresponding face of $G[\bigcup_{B\in\mathcal{B}}B]$ that contains the facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$ (note that the vertices of this facial walk are bounded $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ and the only two other vertices incident to this face are $v_{i},v_{i+1}$). Since $\mathcal{B}$ is a 4-surround of $(G,\mathcal{H})$, it follows that for every $H^{\prime}\in\mathcal{H}$ that contains a vertex incident to $F_{v_{i}v_{i+1}}$, the distance between $H$ and $H^{\prime}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ is at most $4$. This in particular implies that $F_{v_{i}v_{i+1}}$ will satisfy the 4th bullet in the definition of a $8$-fortification of $(G,\mathcal{H})$. This in particular implies that no vertex incident to $F_{v_{i}v_{i+1}}$ is incident to the outerface of $G$. Then observe that $R_{F,H}$ is a connected set in $G\cup E_{F,H}$. For an illustration of such a $R_{F,H},E_{F,H}$, see Figure˜8. Let $G^{*}$ be the planar graph obtained from adding these edges from the union of all such $E_{F,H}$. Let $\mathcal{R}$ be the collection of all such $R_{F,H}$. Then every element of $\mathcal{R}$ is a connected set of $G^{*}$.
Figure 8: The (orange) vertex set $R_{F,H}$ given the (blue) encasing $\mathcal{B}$, the (red) $H\in\mathcal{H}$, and $F\in\mathcal{F}$. The orange vertices are the vertices of $R_{F,H}^{\prime}$. The pink edges are $E_{F,H}$, and the arrows orthogonal to them indicate their sides.
Let $\mathcal{H}_{C}$ be the elements $H$ of $\mathcal{H}$ that are contained in $\bigcup_{B\in\mathcal{B}}B$, and let $\mathcal{H}_{I}$ be the elements $H$ of $\mathcal{H}$ that intersect but are not contained in $\bigcup_{B\in\mathcal{B}}B$. Let $\mathcal{R}^{*}=(\mathcal{H}\backslash(\mathcal{H}_{C}\cup\mathcal{H}_{I}))\cup\mathcal{R}$. Let $G_{0}^{*}=G^{*}\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)$, and let $\mathcal{R}_{0}^{*}$ be the elements of $\mathcal{R}^{*}$ that contain a vertex incident to the outer face of $G_{0}$. Note that $\mathcal{R}^{*}$ is a spanning collection of connected vertex sets of $G_{0}$ and that $\mathcal{R}\subseteq\mathcal{R}^{*}_{0}$.
[…]
If $M\in\mathcal{M}^{\prime}\subseteq\mathcal{M}$ does not contain a vertex incident to the outer face of $G_{0}^{\prime}$, then $I_{\mathcal{H}^{\prime}}(M)=I_{\mathcal{R}^{\prime}}(M)$ has weak diameter at most 9242 in $\textbf{RIG}(G_{0}^{\prime},\mathcal{R}^{\prime})$, and therefore also in $\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})$. If $M\in\mathcal{M}^{\prime}\subseteq\mathcal{M}$ contains a vertex incident to the outer face of $G_{0}^{\prime}$, then $I_{\mathcal{R}^{\prime}}(M)$ has weak diameter at most 9240 in $\textbf{RIG}(G_{0}^{\prime},\mathcal{R}^{\prime})$. Since $I_{\mathcal{H}^{\prime}}(M)\subseteq I_{\mathcal{R}^{\prime}}(M)\cup N_{\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})}(I_{\mathcal{R}^{\prime}}(M))$, it follows that $I_{\mathcal{H}^{\prime}}(M)$ has weak diameter at most $9240+2=9242$ in $\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})$. So, $(G^{\prime},\mathcal{M})$ is a $(9242,\infty)$-impression of $(G^{\prime},\mathcal{H}^{\prime})$.
If $H\in\mathcal{R}^{\prime}\subseteq\mathcal{H}^{\prime}$, then $I_{\mathcal{M}^{\prime}}(H)=I_{\mathcal{M}}(H)$ has weak diameter at most 80 in $\textbf{IM}(G_{0}^{\prime},\mathcal{M}^{\prime})$ and therefore also in $\textbf{IM}(G^{\prime},\mathcal{M})$. To show that $(G^{\prime},\mathcal{M}^{\prime})$ is a $(9242,80)$-impression of $(G^{\prime},\mathcal{H}^{\prime})$, it now just remains to show that if $H\in\mathcal{H}_{I}$, then $I_{\mathcal{M}}(H)$ has weak diameter at most $80$ in $\textbf{IM}(G^{\prime},\mathcal{M})$.
Fix some $H\in\mathcal{H}_{I}$. Consider some $u,v\in H$ and let $M_{u},M_{v}\in\mathcal{M}$ be such that $u\in M_{u}$ and $v\in M_{v}$. Since $\mathcal{B}$ is a 7-cage of $(G,\mathcal{H})$, there exists some sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant 7$ such that
  • each $A_{i}$ is either a single element of $\mathcal{B}$ or a single element of $\mathcal{F}$ or the union of a subcollection $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ such that some connected set $H^{*}\subseteq H$ of $G$ (or equivalently $G^{\prime}$) intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]=G^{\prime}[\bigcup_{B\in\mathcal{B}^{\prime}}B]$,
  • $u\in A_{0}$,
  • $v\in A_{t}$, and
  • for every $1\leqslant i<t$, we have that either $H$ intersects both $A_{i-1}$ and $A_{i}$, or neither $A_{i-1}$ or $A_{i}$ is an element of $\mathcal{F}$ and they touch.
If $A_{i}\in\mathcal{B}\subseteq\mathcal{M}$, then clearly $\{A_{i}\}$ has weak diameter 0 in $\textbf{IM}(G^{\prime},\mathcal{M})$. If $A_{i}$ is the union of a subcollection $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$, then it follows from $\mathcal{B}$ being a $(9240,9,4,7)$-encasing of $(G,\mathcal{H})$ that $A_{i}$ has weak diameter at most $9$ in $\textbf{IM}(G,\mathcal{B})$ and therefore in $\textbf{IM}(G^{\prime},\mathcal{M})$. If $A_{i}$ is a single element $F$ of $\mathcal{F}$, then $R_{F,H}$ contains $H\cap F$. As $R_{F,H}\in\mathcal{R}_{0}^{*}$, we have that $I_{\mathcal{M^{\prime}}}(R_{F,H})$ has weak diameter at most 9 in $\textbf{IM}(G_{0},\mathcal{M}^{\prime})$, and therefore $I_{\mathcal{M}}(R_{F,H})$ has weak diameter at most 9 in $\textbf{IM}(G^{\prime},\mathcal{M})$. As $t\leqslant 7$, it now follows that $d_{\textbf{IM}(G^{\prime},\mathcal{M})}(M_{u},M_{v})\leqslant 7+8\cdot 9=80$. Therefore, $I_{\mathcal{M}}(H)$ has weak diameter at most $80$ in $\textbf{IM}(G^{\prime},\mathcal{M})$.
Hence $(G^{\prime},\mathcal{M})$ is a $(9242,80)$-impression of $(G^{\prime},\mathcal{H}^{\prime})$, as desired. ∎
By Theorem˜5.1 and Lemma˜2.2, we obtain the following.
Theorem 5.2.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(73936,80)$-impression $(G^{\prime},\mathcal{M})$ of $(G^{\prime},\mathcal{H})$, where $G^{\prime}$ is a fortification of $G$ and $\mathcal{H}$ 88-spans $G$.
Since for every finite string graph $S$ there is a finite planar graph $G$ and a spanning collection $\mathcal{H}$ of connected vertex sets of $G$ such that $\textbf{RIG}(G,\mathcal{H})=S$ [54], by Theorem˜5.2 and Proposition˜2.1, we finally get that string graphs are quasi-isometric to planar graphs.
Theorem 5.3.
Let $S$ be a finite string graph. Then there is a planar graph $G$ and a function $f:V(S)\to V(G)$ such that for all $u,v\in V(S)$,
$\frac{1}{73944}d_{S}(u,v)-\frac{73936}{73944}\leqslant d_{G}(f(u),f(v))\leqslant 160d_{S}(u,v),$
and for every $v\in V(G)$ there exists some $u\in V(S)$ such that $d_{G}(f(u),v)\leqslant 80$.
Theorem˜2.5 (and thus Theorem˜1.1) then finally follows from Theorem˜5.3 and Lemma˜2.3.
Acknowledgements
This work was initiated at the Workshop on Graph Product Structure Theory (BIRS21w5235) at the Banff International Research Station, 21–26 November 2021. We are grateful to the organisers and participants for providing a stimulating research environment. We thank a number of people for helpful discussions and encouragement on the project including Sandra Albrechtsen, Marthe Bonamy, Linda Cook, Louis Esperet, Agelos Georgakopoulos, Robert Hickingbotham, Rose McCarty, Harry Petyt, and Alexandra Wesolek.