### v2

# String graphs are quasi-isometric to planar graphs

String graphs are quasi-isometric to planar graphs
James Davies The author was supported by the Alexander von Humboldt Foundation in the framework of the Alexander von Humboldt Professorship of Daniel Kráľ endowed by the Federal Ministry of Education and Research.
Аннотация
We prove that for every countable string graph $S$, there is a planar graph $G$ with $V(G)=V(S)$ such that
$\frac{1}{23660800}d_{S}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{S}(u,v)$
for all $u,v\in V(S)$, where $d_{S}(u,v)$, $d_{G}(u,v)$ denotes the distance between $u$ and $v$ in $S$ and $G$ respectively. In other words, string graphs are quasi-isometric to planar graphs.
This theorem lifts a number of theorems from planar graphs to string graphs, we give some examples. String graphs have Assouad-Nagata (and asymptotic dimension) at most 2. Connected, locally finite, quasi-transitive string graphs are accessible. A finitely generated group $\Gamma$ is virtually a free product of free and surface groups if and only if $\Gamma$ is quasi-isometric to a string graph.
Two further corollaries are that countable planar metric graphs and complete Riemannian planes are also quasi-isometric to planar graphs, which answers a question of Georgakopoulos and Papasoglu. For finite string graphs and planar metric graphs, our proofs yield polynomial time (for string graphs, this is in terms of the size of a representation given in the input) algorithms for generating such quasi-isometric planar graphs.
We further extend our techniques to show that every complete Riemannian surfaces $\Sigma$ of bounded Euler genus has a triangulation $G\subset\Sigma$ such that $G^{(1)}\hookrightarrow\Sigma$ is a quasi-isometry, where $G^{(1)}$ is the simplicial 1-skeleton of $G$.
1 Introduction
A string graph is an intersection graph of a collection $\mathcal{S}$ of path-connected subsets of the plane, where the vertices are $\mathcal{S}$, and $S_{1},S_{2}\in\mathcal{S}$ are adjacent if they intersect. String graphs were first formally defined by Ehrlich, Even, and Tarjan [34] but were present in earlier works of Benzer [10] on the topology of genetic structures and of Sinden [89] on electrical networks realizable by printed circuit. String graphs are the most general class of intersection graphs in the plane and significantly generalize planar graphs, which by the Koebe-Andreev-Thurston circle packing theorem [59] can be viewed as intersection graphs of discs in the plane with disjoint interiors. String graphs can be thought of as a dense or induced minor analogue of planar graphs.
There are several important structural results for string graphs. For instance, Pach and Tóth [82] proved that recognizing string graphs is decidable. Fox and Pach [38, 39, 41] (see also [71]) proved a separator theorem for string graphs. Every incomparability graph is a string graph and Fox and Pach [40] proved a partial converse of this for dense string graphs. This was used by Tomon [91] to prove that string graphs have the Erdős-Hajanl property. Very recently, Abrishami, Briański, Davies, Du, Masaříková, Rzążewski, and Walczak [1] classified string graphs with large chromatic number.
Every planar graph is a string graph, although the structure of string graphs is much more complicated. For instance, while planar graphs can be recognized in linear time [55], recognition of string graphs is NP-hard [62, 88]. There are at most $c^{n}n!$ labelled planar graphs on $n$ vertices [28, 79], while the number of labelled string graphs on $n$ vertices is $2^{(\frac{3}{4}+o(1)){n\choose 2}}$ [83]. Every planar graph is 4-colourable [6, 86], however there are triangle-free string graphs with arbitrarily large chromatic number [84].
We prove a partial converse to the fact that planar graphs are string graphs. In a coarse sense, every string graph is equivalent to some planar graph. For vertices $u,v$ of a graph $G$, we let $d_{G}(u,v)$ denote the distance between $u$ and $v$ in $G$.
Theorem 1.1.
Let $S$ be a countable string graph.  [^1] We remark that Theorem 1.1 does not hold for uncountable string graphs as can be seen by Wagner’s [94] characterization of uncountable planar graphs and the fact that the graph obtained from $|\mathbb{R}|$ 1-ended infinite paths by identifying their ends and then for each vertex adding $|\mathbb{R}|$ new adjacent vertices of degree 1 is a string graph. Then there is a planar graph $G$ with $V(G)=V(S)$ such that
$\frac{1}{23660800}d_{S}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{S}(u,v)$
for all $u,v\in V(S)$.
In other words, string graphs are quasi-isometric (or even bi-Lipschitz equivalent) to planar graphs. For finite string graphs the proof of Theorem˜1.1 yields a polynomial time (with respect to the size of a given representation of $S$ in the input) algorithm for generating such a planar graph $G$. Unfortunately, there are string graphs that require exponentially sized representations [61]. However, natural subclasses such as segment intersection graphs or more generally intersection graphs of convex compact path-connected subsets of $\mathbb{R}^{2}$, do have polynomially sized representations. We remark that if the string graph $S$ is locally finite, then the resulting planar graph $G$ is also clearly locally finite.
Theorem˜1.1 has numerous applications. We discuss some direct corollaries for string graphs. Afterwards, we shall discuss further quasi-isometry results on metric planar graphs and complete Riemannian surfaces before giving a couple more applications.
Asymptotic dimension was introduced by Gromov [51] for metric spaces and groups (via their Cayley graphs). This notion has had significant impact in geometric group theory, for instance Yu [95] proved that finitely generated groups with finite asymptotic dimension satisfy the Novikov conjecture. For a survey of asymptotic dimension from this point of view, see [9]. In coarse graph theory, asymptotic dimension can be thought of as a coarse analogue of graph colouring. There has recently been significant progress on the study of asymptotic dimension in the intersection of structural graph theory and geometric group theory [1, 11, 12, 25, 29, 33, 64]. There are also applications of asymptotic dimension to graph colouring and $\chi$-boundedness, for instance hereditary classes of graphs with bounded asymptotic dimension are either $\chi$-bounded or contain all Burling graphs [1].
One of the most natural classes of graphs to consider is of course planar graphs. Ostrovskii and Rosenthal [81] proved that planar graphs have bounded asymptotic dimension. Fujiwara and Papasoglu [42] improved this bound to 3 and Jørgensen and Lang [58] proved that planar graphs have asymptotic dimension and furthermore Assouad–Nagata dimension at most 2, which is tight due to the infinite planar grid. Asymptotic dimension (and Assouad–Nagata dimension in the case of graphs) are persevered by quasi-isometries. Thus, by Theorem˜1.1, we obtain the following.
Theorem 1.2.
String graphs have Assouad–Nagata dimension at most 2.
Theorem˜1.2 settles the 2-dimensional case of a conjecture of Davies, Georgakopoulos, Hatzel, and McCarty [25] that sphere intersection graphs in $\mathbb{R}^{d}$ have asymptotic dimension at most $d$. For $d\geqslant 3$, such graphs have unbounded Assouad–Nagata dimension [25].
Bonamy, Bousquet, Esperet, Groenland, Liu, Pirot, and Scott [12] gave a proof that planar graphs have Assouad–Nagata dimension at most 2, which yields a polynomial time algorithm for computing the corresponding covers. Thus, the corresponding covers in Theorem˜1.2 can also be computed in polynomial time (in terms of the size of the provided representation of the string graph given in the input). For some algorithmic applications of such covers, see [8, 13, 14, 35, 37, 57, 63, 65].
Dunwoody [32] showed that every connected, locally finite, quasi-transitive, planar graph is accessible (see also [36, 53, 54]). MacManus [66] recently gave a remarkable extension of this theorem by showing that every connected, locally finite, quasi-transitive graph that is quasi-isometric to a planar graph is accessible. Thus, by Theorem˜1.1, Dunwoody’s [32] theorem extends from planar graphs to string graphs.
Theorem 1.3.
Every connected, locally finite, quasi-transitive string graphs is accessible.
There has been significant interest in planar Cayley graphs [7, 31, 30, 32, 49, 45, 46, 50, 70, 92] and Cayley graphs quasi-isometric to planar graphs [16, 19, 44, 66, 67, 68, 73, 93]. Closely related to this, there are several results characterizing certain finitely presented groups $\Gamma$ in terms of being quasi-isometric to certain Riemannian surfaces. For instance, $\Gamma$ is quasi-isometric to a complete Riemannian plane if and only if $\Gamma$ is a virtual surface group [16, 19, 44, 68, 73, 93]. Maillot also showed that $\Gamma$ is quasi-isometric to a complete, simply connected, planar Riemannian surface with non-empty geodesic boundary if and only if $\Gamma$ is virtually free. In both these cases, $\Gamma$ is quasi-isometric to a planar graph since both surface groups and free groups have planar Cayley graphs.
MacManus [66] used his aforementioned theorem on accessibility to extend and unify these two theorems on finitely generated groups. By Theorem˜1.1, we obtain the following further extension of MacManus’s [66] theorem (our contribution is the equivalence between the first two bullets, the rest is given in [66]).
Theorem 1.4.
Let $\Gamma$ be a finitely generated group. Then the following are equivalent.
  • $\Gamma$ is quasi-isometric to a string graph.
  • $\Gamma$ is quasi-isometric to a planar graph.
  • $\Gamma$ is quasi-isometric to a planar Cayley graph.
  • Some finite index subgroup of $\Gamma$ admits a planar Cayley graph.
  • $\Gamma$ is virtually a free product of free and surface groups.
We now turn our attention to some further quasi-isometry results on metric planar graphs and complete Riemannian surfaces.
A metric graph $H$ is a graph along with an assignment of non-negative edge lengths (in this paper, the length of edges in (non-metric) graphs is otherwise always uniformly equal to 1). We can view metric graphs as a length space in which the edges are replaces by intervals of the corresponding lengths. When the edge lengths are within $D\subseteq(0,\infty)$, we say that $H$ is an $D$-metric graph. The coarse geometry of metric graphs tends to be harder to work with than that of non-metric (simplicial) graphs [64].
Georgakopoulos and Papasoglu [47] asked whether every planar metric graph is quasi-isometric to a planar graph. As a straightforward application of Theorem˜1.1 (see Lemma˜2.4), we answer this in positive (note that every planar metric graph can be turned into a $(0,1]$-metric planar graph by subdividing edges appropriately). Such a planar graph $G$ can be generated in polynomial time.
Theorem 1.5.
Let $H$ be a countable $(0,1]$-metric planar graph. Then there is a planar graph $G$ with $V(G)=V(H)$ such that
$\frac{1}{47321600}d_{H}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{H}(u,v)+1$
for all $u,v\in V(H)$.
Beyond the general interest in metric planar graphs and also in groups acting on planar surfaces [50], one other motivation of Georgakopoulos and Papasoglu [47] was their coarse Kuratowski conjecture that graphs forbidding $K_{5}$ and $K_{3,3}$ as $K$-fat minors are quasi-isometric to planar graphs. Although Theorem˜1.5 and Theorem˜1.1 provide substantial progress towards this conjecture, we remark that the closely related graph minor analogue that graphs forbidding $H$ as a $K$-fat minor are quasi-isometric to $H$-minor-free graphs is false [2, 26], (although it is known to be true for some small graphs $H$ [3, 5, 22, 43, 47, 69, 75]). Such $K$-fat $H$-minor-free graphs need not even be quasi-isometric to $H^{\prime}$-minor-free graphs for any graph $H^{\prime}$ [2].
It is well known (see [12, 18, 23, 48, 80, 87]) that complete Riemannian surfaces are quasi-isometric to locally finite [^2] A $(0,1]$-metric graph is locally finite if every bounded radius ball contains at most a finite number of vertices. $(0,1]$-metric graphs embeddable in the surface. This was used for instance to show that complete Riemannian surfaces of bounded Euler genus (such as complete Riemannian planes) have Assouad–Nagata dimension at most 2 [12] and in Maillot’s [68] aforementioned classification of finitely generated virtually free groups as being those quasi-isometric to certain planar Riemannian surfaces. Ntalampekos and Romney [80] also used a strong version for length surfaces as a step in their proof that any length surface is the Gromov–Hausdorff limit of polyhedral surfaces with controlled geometry. They gave further applications of this, including a new proof of the Bonk-Kleiner theorem [15] characterizing Ahlfors 2-regular quasispheres.
By Theorem˜1.5, we obtain the following refinement, which significantly generalizes the previously discussed fact that finitely generated groups quasi-isometric to complete Riemannian planes are quasi-isometric to planar graphs.
Theorem 1.6.
Complete Riemannian planes are quasi-isometric to locally finite planar graphs.
In fact, we shall further extend our techniques to refine Theorem˜1.6 in three ways. Firstly, we obtain a quasi-isometry which is also an embedding of the graph into the Riemannian plane (as in [12, 18, 23, 48, 80, 87]). Secondly, we also extend this result to Riemannian surfaces of bounded Euler genus. Lastly, we show that such a graph and embedding can be chosen to be a triangulation. A key ingredient in this last refinement is a recent result of Georgakopoulos and Vigolo [48] that complete Riemannian surfaces admit triangulations by metric graphs that are quasi-isometric to the surface. For an embedded graph $G\subset\Sigma$, we denote by $G^{(1)}$ its simplicial 1-skeleton (in which all edges have length 1).
Theorem 1.7.
Let $\Sigma$ be a complete Riemannian surface (without boundary) of bounded Euler genus. Then $\Sigma$ has a triangulation $G\subset\Sigma$ such that $G^{(1)}\hookrightarrow\Sigma$ is a quasi-isometry.
The quasi-isometry in Theorem˜1.7 can be taken to be a $(10^{6},10^{6}+10^{4}g)$-quasi-isometry where $g$ is the Euler genus of the Riemannian surface. The dependence on the Euler genus is necessary since for each fixed $M\geqslant 1$, there are complete Riemannian surfaces of genus at most $g$ that admit no such $(M,\Omega(\sqrt{\log g}))$-quasi-isometric simplicial triangulation [27]. In particular, Theorem˜1.7 does not hold for complete Riemannian surfaces of unbounded genus [27]. Georgakopoulos and Vigolo [48] recently showed that complete Riemannian surfaces with uniform nets also admit quasi-isometric simplicial triangulations as in Theorem˜1.7, and Bowditch [17] showed that complete Riemannian manifolds of bounded geometry also admit such simplicial triangulations.
We remark that while we prove Theorem˜1.7 for Riemannian surfaces without boundary, it is possible to extend Theorem˜1.7 to Riemannian surfaces with boundary with more technical arguments and with the same resulting $(10^{6},10^{6}+10^{4}g)$-quasi-isometry bounds.
We expect our quasi-isometry results to have many further applications. We quickly give two more concerning $\ell_{1}$-embeddings and the coarse Menger conjecture.
A well-known conjecture (see [52]) states that planar graphs admit embeddings into $\ell_{1}$ with bounded distortion. Theorem˜1.1 implies that this is equivalent to string graphs having embeddings into $\ell_{1}$ with bounded distortion. Rao [85] showed that $n$-vertex planar graphs have embeddings into $\ell_{1}$ with distortion at most $O(\sqrt{\log n})$, and Theorem˜1.1 extends this to string graphs.
Theorem 1.8.
Every $n$-vertex string graph has a $O(\sqrt{\log n})$-distortion embedding into $\ell_{1}$.
For a more extensive survey and discussion of the applications of such embeddings, see [52].
Albrechtsen, Huynh, Jacobs, Knappe and Wollan [4], and independently Georgakopoulos and Papasoglu [47] conjectured a coarse analogue of Menger’s [72] theorem that for a graph $G$ containing two vertex sets $S,T$, either there are $k+1$ paths between $S$ and $T$ pairwise at distance at least $c$, or there is a set of at most $k$ balls in $G$ of radius at most $f(c,k)$ such that every path between $S$ and $T$ intersects one of these balls. This turns out to be false in general [74], and even in the weak version where up to $f(c,k)$ balls are allowed [76]. However, it is true for certain classes such as those of bounded path-width [77]. A particularly interesting open case is that of planar graphs. Nguyen, Scott, and Seymour [78] proved a version for paths between vertices on the outerface of a planar graph. By Theorem˜1.5, we obtain the following extension of their result to metric planar graphs [^3] We remark that there is a technicality of preserving vertices on the outerface, however, the proof of Theorem 1.5 gives this. (they proved the same theorem for (simplicial) graphs, just with a much better constant than $10^{10}$).
Theorem 1.9.
Let $c,k\geqslant 0$, and let $G$ be a metric planar graph, with vertex sets $S,T$ on the outerface. Then either:
  • there are $k+1$ paths between $S$ and $T$, pairwise at distance least $c$; or
  • there is a set of at most $k$ balls in $G$ of radius at most $10^{10}k^{3}c$ of G, such that every path between $S$ and $T$ intersects one of these balls.
One can also obtain versions of Theorem˜1.9 for string graphs and complete Riemannian surfaces homeomorphic to a disk.
In Section˜2 we introduce basic definitions, deduce Theorem˜1.5 from Theorem˜1.1, and also reduce proving Theorem˜1.1 to its finite case (Theorem˜2.5). We also introduce the necessary framework for our approach to proving Theorem˜1.1, namely through viewing string graphs as region intersection graphs, and the notion of ‘‘impressions’’ which provides a convenient way of finding quasi-isometries from region intersection graphs. The rest of the proof in Sections˜3, 4 and 5 is essentially a delicate technical induction. In Section˜3 we prove a weakening of Theorem˜1.1 for the subclass outerstring graphs (Theorem˜3.1). This is used in Section˜4 to find a technical structure tailored for use in the inductive argument that we call an ‘‘encasing’’ (see Lemma˜4.4). Finally, in Section˜5, we prove Theorem˜1.1 using Lemma˜4.4. Lastly, in Section˜6, we build on one version of our main technical result (Theorem˜5.2) to prove Theorem˜1.7.
Remark 1.10.
More recently, Chang, Conroy, Tan, and Zheng [21] have independently obtained another proof of a result very similar to Theorem˜1.1. While our proof is self contained and builds the quasi-isometry from first principles, theirs adapts a shortcut partition construction using gridtrees developed in [20]. We remark that their motivation and applications differs significantly from the present paper and we encourage the interested reader to also see [21] for further applications. For instance, they construct compact (approximate) distance oracles for string graphs and also prove that for any $\epsilon>0$, $n$-vertex string graphs have $(1+\epsilon,O(\epsilon^{-4}\log^{18}n))$-quasi-isometric metric planar graphs.
In [24], we will build on the work of the present paper and [21] to prove versions of Theorem˜1.5 and Theorem˜1.1 for minor-free graphs.
2 Preliminaries
For $M,A\geqslant 0$ with $M\geqslant 1$, an $(M,A)$-quasi-isometry from a metric space $G$ to another metric space $H$ is a map $f:G\to H$ such that
  1. 1.
    $M^{-1}d_{G}(u,v)-A\leqslant d_{H}(f(u),f(v))\leqslant Md_{G}(u,v)+A$ for every $u,v\in G$, and
  2. 2.
    for every $h\in H$, there exists some $v\in G$ with $d_{H}(h,f(v))\leqslant A$.
We say that $G$ is $(M,A)$-quasi-isometric to $H$ if there exists a $(M,A)$-quasi-isometry from $G$ to $H$. Two metric spaces $G$ and $H$ are quasi-isometric if they are $(M,A)$-quasi-isometry for some $M,A\geqslant 0$. Quasi-isometries for graphs are of course defined in the same expected way.
[…]
Given a graph $G$ and a collection of connected vertex sets $\mathcal{H}$ of $G$, we let $\textbf{RIG}(G,\mathcal{H})$ be the graph with vertex set $\mathcal{H}$ where, $H_{1},H_{2}\in\mathcal{H}$ are adjacent if they intersect. We call $\textbf{RIG}(G,\mathcal{H})$ a region intersection graph. Region intersection graphs were introduced by Lee [63] as a generalization of string graphs (usually we consider region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ in which $G$ forbids some graph as a minor, since otherwise the class contains all graphs). Finite string graphs are exactly finite region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ for which $G$ is planar [63] and this is how we will view string graphs to prove Theorem˜1.1 (or its finite version, Theorem˜2.5).
If $\mathcal{M}$ is a collection of disjoint connected vertex sets in $G$, then we let $\textbf{IM}(G,\mathcal{M})$ be the graph with vertex set $\mathcal{M}$ where $M_{1},M_{2}\in\mathcal{M}$ are adjacent if there is an edge in $G$ between $M_{1}$ and $M_{2}$ (which case, we say that they touch). This is one way to define induced minors of a graph $G$.
Let $G$ be a graph and $\mathcal{H}$ a collection of connected vertex sets of $G$. If every edge of $G$ is contained in $G[H]$ for some $H\in\mathcal{H}$, then we say that $\mathcal{H}$ spans $G$. More generally, we say that $\mathcal{H}$ $z$-spans $G$ if for every edge $uv$ of $G$, there exists $H_{u},H_{v}\in\mathcal{H}$ with $u\in H_{u}$ and $v\in H_{v}$, such that $d_{\textbf{RIG}(G,\mathcal{H})}(H_{u},H_{v})\leqslant z$. Note that $\mathcal{H}$ spans $G$ if and only if it 0-spans $G$.
For a graph $G$ and two collections of connected vertex sets $\mathcal{H},\mathcal{M}$ of $G$ with the sets in $\mathcal{M}$ being disjoint, we say that $(G,\mathcal{M})$ is a $(x,y)$-impression of $(G,\mathcal{H})$ if
  • for every $M\in\mathcal{M}$, $I_{\mathcal{H}}(M)$ has weak diameter at most $x$ in $\textbf{RIG}(G,\mathcal{H})$,
  • for every $H\in\mathcal{H}$, $I_{\mathcal{M}}(H)$ has weak diameter at most $y$ in $\textbf{IM}(G,\mathcal{M})$, and
  • $\bigcup_{M\in\mathcal{M}}M=V(G)$.
Impressions provide a convenient way of obtaining a quasi-isometry from a region intersection graph $\textbf{RIG}(G,\mathcal{H})$ to a minor of $G$.
Proposition 2.1.
Let $(G,\mathcal{M})$ be a $(x,y)$-impression of $(G,\mathcal{H})$ where $\mathcal{H}$ $z$-spans $G$. Then there is a function $f:\mathcal{H}\to\mathcal{M}$ such that for all $H_{1},H_{2}\in\mathcal{H}$, we have
$\frac{1}{x+z}d_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2})-\frac{x}{x+z}\leqslant d_{\textbf{IM}(G,\mathcal{M})}(f(H_{1}),f(H_{2}))\leqslant 2yd_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2}),$
and for every $M\in\mathcal{M}$ there exists some $H\in\mathcal{H}$ such that $d_{\textbf{IM}(G,\mathcal{M})}(f(H),M)\leqslant y$.
Furthermore, $f$ can be chosen to be any function $f:\mathcal{H}\to\mathcal{M}$ such that $f(H)\in I_{\mathcal{M}}(H)$ for all $H\in\mathcal{H}$.
Доказательство.
Let $f:\mathcal{H}\to\mathcal{M}$ be chosen so that $f(H)\in I_{\mathcal{M}}(H)$ for all $H\in\mathcal{H}$. Clearly such a function exists since $I_{\mathcal{M}}(H)$ is always non-empty.
If $H_{1}$ and $H_{2}$ are adjacent in $\textbf{RIG}(G,\mathcal{H})$, then they intersect. So then $I_{\mathcal{M}}(H_{1}\cup H_{2})$ is a connected set in $\textbf{IM}(G,\mathcal{M})$ and furthermore has weak diameter at most $2y$. It follows for (possibly non-adjacent $H_{1},H_{2}\in\mathcal{H}$) that $d_{\textbf{IM}(G,\mathcal{M})}(f(H_{1}),f(H_{2}))\leqslant 2yd_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2})$.
Suppose that $d_{\textbf{IM}(G,\mathcal{M})}(f(H_{0}),f(H_{t}))=t$ and let $M_{0},\ldots,M_{t}$ be a path in $\textbf{IM}(G,\mathcal{M})$ between $M_{0}=f(H_{0})$ and $M_{t}=f(H_{t})$. For each $1\leqslant i<t$, there exists some $u_{i}v_{i}\in E(G)$ with $u_{i}\in M_{i-1}$ and $v_{i}\in M_{i}$. For each $1\leqslant i<t$, let $H_{i}^{\prime},H_{i}\in\mathcal{H}$ be such that $u_{i}\in V(H_{i}^{\prime})$, $v_{i}\in V(H_{i})$, and the distance between $H_{i}^{\prime}$ and $H_{i}$ in $\textbf{RIG}(G,\mathcal{H})$ is at most $z$. For each $1\leqslant i\leqslant t$, we have that $H_{i-1},H_{i}^{\prime}\in I_{\mathcal{H}}(M_{i-1})$. Thus, there is a path in $\textbf{RIG}(G,\mathcal{H})$ between $H_{0}$ and $H_{t}$ of length at most $(t+1)x+tz$. It now follows that for all $H_{1},H_{2}\in\mathcal{H}$, we have
$\frac{1}{x+z}d_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2})-\frac{x}{x+z}\leqslant d_{\textbf{IM}(G,\mathcal{M})}(f(H_{1}),f(H_{2}))\leqslant 2yd_{\textbf{RIG}(G,\mathcal{H})}(H_{1},H_{2}).$
Lastly, consider some $M\in\mathcal{M}$. Choose some $H\in I_{\mathcal{H}}(M)$. Then $M,f(H)\in I_{\mathcal{M}}(H)$, and thus $M$ is at distance at most $y$ from $f(M)$ in $\textbf{IM}(G,\mathcal{M})$. ∎
In most but not all cases we consider, we will actually have that $\mathcal{H}$ spans $G$, which means that we simply have $z=0$ in Proposition˜2.1. Given $(G,\mathcal{H})$, we could simply reduce to the case the $\mathcal{H}$ spans $G$ by removing edges of $G$ not contained in $G[H]$ for any $H\in\mathcal{H}$ (which is something we often do). However, adding edges to $G$ can also make finding impressions easier. In fact, for $G$ planar, it is convenient for inductive arguments that we can also add new connected sets to $\mathcal{H}$ in a controlled way so that our final impression can be obtained from a related impression.
Lemma 2.2.
[…]
Доказательство.
[…]
It is straight forward to make the quasi-isometry $f$ in Proposition˜2.1 into a bijective map by taking a minor and possibly adding some pendent edges. Note in particular that for countable graphs this preserves planarity.
Lemma 2.3.
Let $G,H$ be graphs and let $f:V(G)\to V(H)$ be a function such that for all $u,v\in V(G)$, we have
$\frac{1}{x_{1}}d_{G}(u,v)-\frac{x_{2}}{x_{1}}\leqslant d_{H}(f(u),f(v))\leqslant x_{3}d_{G}(u,v),$
and for every $v\in V(H)$ there exists some $u\in V(G)$ such that $d_{H}(f(u),v)\leqslant x_{4}$.
Then there exists some graph $H^{\prime}$ obtained from $H$ by adding vertices adjacent to individual vertices of $V(H)$ and then taking a minor, and some bijection $f^{\prime}:V(G)\to V(H^{\prime})$ such that for all $u,v\in V(G)$, we have
$\frac{1}{2x_{4}(x_{1}+x_{2})}d_{G}(u,v)\leqslant d_{H^{\prime}}(f^{\prime}(u),f^{\prime}(v))\leqslant(x_{3}+2)d_{G}(u,v).$
Доказательство.
Let $A=f(V(G))$. Since for every $v\in V(H)$ there exists some $u\in V(G)$ such that $d_{H}(f(u),v)\leqslant x_{4}$, we can choose a forest $F$ contained in $H$ as a subgraph with $V(F)=V(H)$ such that every connected component of $F$ contains exactly one vertex of $A$ and every vertex of $F$ is at distance at most $x_{4}$ from $A$ in $F$. Let $H^{*}$ be obtained from $H$ by for each $a\in A$, adding $|f^{-1}(a)|-1$ vertices to $H$ that are adjacent to $a$ only. Choose $f^{\prime}:V(G)\to V(H^{*})$ so that for every $a\in A$, $f^{-1}(a)$ is mapped bijectivly to $a$ and the newly added $|f^{-1}(a)|-1$ vertices adjacent to $a$. Let $H^{\prime}$ be obtained from $H^{*}$ by contracting the edges of $F$ onto $A$. Then $f^{\prime}$ is a bijection between $G$ and $H^{\prime}$.
For every $u,v\in V(G)$, we have
$\frac{1}{2x_{4}}d_{H}(f(u),f(v))\leqslant d_{H^{\prime}}(f^{\prime}(u),f^{\prime}(v))\leqslant d_{H}(f(u),f(v))+2.$
For distinct $u,v\in V(G)$, we clearly have $d_{H}(f(u),f(v))+2\leqslant 3d_{H}(f(u),f(v))\leqslant x_{3}d_{G}(u,v)+2\leqslant(x_{3}+2)d_{G}(u,v)$. For $t\geqslant 0$, we have that $\frac{1}{x_{1}+x_{2}}t\leqslant\max\{\frac{1}{x_{1}}t-\frac{x_{2}}{x_{1}},1\}$. Since $f^{\prime}$ is bijective, it therefore follows that $\frac{1}{2x_{4}(x_{1}+x_{2})}d_{G}(u,v)\leqslant d_{H^{\prime}}(f^{\prime}(u),f^{\prime}(v))$. ∎
Theorem˜1.1 for finite string graphs will be quickly deduced in Section˜5 from our main technical result (Theorem˜5.1) by applying Lemma˜2.2, Proposition˜2.1, and Lemma˜2.3 in order.
Next we show that Theorem˜1.1 implies Theorem˜1.5. This is immediate from the following lemma.
Lemma 2.4.
Let $H$ be a $(0,1]$-metric planar graph. Then there is a string graph $S$ with $V(S)=V(H)$ such that
$\frac{1}{2}d_{H}(u,v)\leqslant d_{S}(u,v)\leqslant d_{H}(u,v)+1$
for all $u,v\in V(H)$.
Доказательство.
Consider $H$ embedded in the plane. For each $u\in V(H)$, let $S_{u}$ be the subset of the plane consisting of all points of $H$ at distance at most 1 in $H$ from $u$. Let $S$ be the string graph with vertex set $V(H)$ in which each $u\in V(S)=V(H)$ is represented by $S_{u}$. Consider $u,v\in V(H)$ and let $P$ be a geodesic path from $u$ to $v$ in $H$. Choose vertices $x_{0},\ldots,x_{t}$ of $H$ along $P$ so that $x_{0}=u$, $x_{t}=v$ and for each $0\leqslant i<t$, the distance along $P$ between $x_{i}$ and $x_{i+1}$ is in $[1,2]$ (or if $d_{H}(u,v)<1$, we just take $x_{0}=u$, $x_{1}=v$). So, $\lceil\frac{1}{2}d_{H}(u,v)\rceil\leqslant t\leqslant\lfloor d_{H}(u,v)\rfloor$. The lemma now follows since $x_{i}$ is adjacent to $x_{i+1}$ in $S$ for each $0\leqslant i<t$. ∎
We remark that with a very similar proof to Lemma˜2.4, it is easy to extend Theorem˜1.1 to $(0,1]$-metric string graphs (again with slightly worse bounds).
For our inductive argument, it is more convenient to prove Theorem˜1.1 for finite string graphs. Section˜3, Section˜4, and Section˜5 are dedicated to proving Theorem˜1.1 for finite string graphs (Theorem˜2.5 below), which in particular, will be quickly deduced in Section˜5 from Theorem˜5.1 (which is a more technical result tailored for an inductive argument) by applying Lemma˜2.2, Proposition˜2.1, and Lemma˜2.3 in order.
Theorem 2.5.
Let $S$ be a finite string graph. Then there is a planar graph $G$ with $V(G)=V(S)$ such that
$\frac{1}{23660800}d_{S}(u,v)\leqslant d_{G}(u,v)\leqslant 162d_{S}(u,v)$
for all $u,v\in V(S)$.
We finish this section by showing that Theorem˜2.5 implies Theorem˜1.1.
Proof of Theorem˜1.1 assuming Theorem˜2.5..
Clearly it is enough to consider the case that $S$ is connected. Let $V(S)=\{x_{k}:k\geqslant 1\}$ and for each $i\leqslant 1$, let $X_{i}=\{x_{k}:1\leqslant k\leqslant i\}$. For each $i\geqslant 1$, let $\mathcal{G}_{i}$ be the collection of embedded planar graphs $G^{\prime}$ that are topologically inequivelant after possibly relabeling vertices of $V(G^{\prime})\backslash X_{i}$, such that $X_{i}\subseteq V(G^{\prime})\subseteq V(S)$, $|V(G^{\prime})|\leqslant i+\sum_{x,y\in X_{i}}162d_{S}(x,y)$, and for every $x,y\in X_{i}$, we have that
$\frac{1}{23660800}d_{S}(x,y)\leqslant d_{G^{\prime}}(x,y)\leqslant 162d_{S}(x,y).$
Note that each $\mathcal{G}_{i}$ is non-empty by Theorem˜2.5 (applied to some $S[Y]$ for some $Y$ with $X_{i}\subseteq Y$, $|Y|\leqslant|X_{i}|+\sum_{x,y\in X_{i}}d_{S}(x,y)$ such that $d_{S[Y]}(x,y)=d_{S}(x,y)$ for all $x,y\in X_{i}$). Also, each $|\mathcal{G}_{i}|$ is finite. Let $T$ be the graph with vertex set $\bigcup_{i=1}^{\infty}\mathcal{G}_{i}$ where each $G^{\prime\prime}\in\mathcal{G}_{i+1}$ is adjacent to some $G^{\prime}\in\mathcal{G}_{i}$ such that $G^{\prime}$ is topologically equivalent to a restriction of $G^{\prime\prime}$ after possibly relabeling vertices of $V(G^{\prime\prime})\backslash X_{i}$. Observe that by considering shortest paths between vertices of $X_{i}$, such a $G^{\prime}$ always exists.
By Kőnig’s infinity lemma [60], $T$ has an infinite path $G_{1},G_{2},\ldots$ in which $G_{i}\in\mathcal{G}_{i}$ for each $i\geqslant 1$. Note that for each $i\geqslant 1$, there exists some $R$ so that $V(G_{i})\subseteq V(G_{r})$ for all $r\geqslant R$. We obtain the desired countable planar graph $G$ by iteratively extending an embedded planar graph topically equivalent to $G_{i}$ to one topologically equivalent to $G_{i+1}$ for each $i\geqslant 1$ in order. ∎
3 Outerstring graphs
Outerstring graphs are equivalent to region intersection graphs $\textbf{RIG}(G,\mathcal{H})$ for which $G$ is planar and every $H\in\mathcal{H}$ contains a vertex of the outer face of $G$. They can be thought of as an ‘‘outerplanar’’ analogue of string graphs, and are thus much simpler. The main aim of this section is to prove the following theorem on finding certain restricted impressions of such region intersection graphs. This will be crucial to our arguments on string graphs (or strictly speaking, Lemma˜3.4 and Lemma˜3.2 which easilly imply the following theorem will be).
Theorem 3.1.
Let $G$ be a finite planar graph and $\mathcal{H}$ be a spanning collection of connected vertex sets in $G$, each of which contains a vertex incident to the outer face of $G$. Then there is a $(770,9)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ such that each set in $\mathcal{M}$ contains a vertex incident to the outer face of $G$.
Note that if $(G,\mathcal{M})$ is an impression of $(G,\mathcal{H})$ where each graph in $\mathcal{M}$ contains a vertex incident to the outer face of $G$, then $\textbf{IM}(G,\mathcal{M})$ is an outerplanar graph.
Рис. 1: The graph $F_{6}$.
By Proposition˜2.1, Theorem˜3.1 implies that outerstring graphs are quasi-isometric to outerplanar graphs. This result is already easily implied by a theorem that graphs forbidding $K_{2,3}$ as a fat minor are quasi-isometric to cacti [22, 43], where a cacti is a planar graph in which every edge is on the outer face, or equivalently, a graph in which each maximal 2-connected subgraph is either an edge of a cycle. One might hope that Theorem˜3.1 could be improved so that $\textbf{IM}(G,\mathcal{M})$ is also a cacti. However this is not possible. For every $\ell\geqslant 3$, let $F_{\ell}$ be the outerplanar graph consisting of a cycle of length $\ell$ where for every edge of this cycle, we identify the edge with an edge of another cycle of length $\ell$ (see Fig.˜1 for an illustration of the graph $F_{6}$). Then it can easily be seen that there is no $x,y$ such that every $(F_{\ell},\{\{u,v\}:uv\in E(F_{\ell})\})$ has a $(x,y)$-impression $(F_{\ell},\mathcal{M}_{\ell})$ in which $\textbf{IM}(F_{\ell},\mathcal{M}_{\ell})$ is a cacti. It is however a nice exercise to show that outerplanar graphs are quasi-isometric to cacti. With this, Theorem˜3.1 and Proposition˜2.1 then yeild another proof that outerstring graphs are quasi-isometric to cacti.
[…]
Lemma 3.2.
Let $G$ be a finite outerplanar graph and $\mathcal{H}$ a spanning collection of connected vertex sets in $G$. Then there is a $(11,9)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$.
Доказательство.
Clearly it is enough to handle the case that $G$ is connected. Choose some $H_{0}\in\mathcal{H}$ and for every non-negative integer $i$, let $L_{i}$ be the vertices of $\textbf{RIG}(G,\mathcal{H})$ at distance $i$ from $H_{0}$ in $\textbf{RIG}(G,\mathcal{H})$. For each $i$ we define the following two vertex sets in $G$. Let $D_{i}=\bigcup_{H\in L_{i}}H$, and let $A_{i}=D_{i}\backslash D_{i-1}$ (where $D_{-1}=\emptyset$). For each $A_{i}$, let $\mathcal{M}_{i}$ be the maximal connected vertex sets of $G[A_{i}]$. Let $\mathcal{M}=\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$. See Fig.˜2 for an illustration. We must show that $(G,\mathcal{M})$ is a $(11,9)$-impression of $(G,\mathcal{H})$.
Рис. 2: An illustration of the resulting collection of disjoint connected sets $\mathcal{M}$. The connected sets in $\mathcal{H}$ are the red sets and the sets in $\mathcal{M}$ are the blue ones. The dashed purple lines indicate the partition of $\mathcal{M}$ into $\bigcup_{i=0}^{\infty}\mathcal{M}_{i}$.
Next, we show for every $M\in\mathcal{M}$ that $I_{\mathcal{H}}(M)$ has weak diameter at most $11$ in $\textbf{RIG}(G,\mathcal{H})$. Let $M\in\mathcal{M}$. Then, for some non-negative integer $i$, we have that $M\in\mathcal{M}_{i}$. We may assume that $i\geqslant 2$, since otherwise we clearly have that $I_{\mathcal{H}}(M)$ has weak diameter at most $4$ in $\textbf{RIG}(G,\mathcal{H})$. Since $G$ is outerplanar, observe that there exists a $2$-seperation $(X,Y)$ of $G$ with $A_{i-2}\subseteq X$, $X\cap Y\subseteq A_{i-1}$, and $M\subseteq Y$. To see this one can apply Menger’s theorem between $\bigcup_{0\leqslant j\leqslant i-2}A_{j}$ and the connected component of $\bigcup_{j\geqslant i}A_{j}$ that contains $M$. For each $x\in X\cap Y$, let $\mathcal{H}_{x}=\{H\in L_{i-1}:x\in V(H)\}$. So, $\mathcal{H}_{x}$ is a clique in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, each $H\in I_{\mathcal{H}}(M)$ is at distance at most $2$ from $\bigcup_{x\in X\cap Y}\mathcal{H}_{x}$ in $\textbf{RIG}(G,\mathcal{H})$. Since $|X\cap Y|\leqslant 2$ and $\textbf{RIG}(G,\mathcal{H})[I_{\mathcal{H}}(M)]$ is connected, it now follows that $I_{\mathcal{H}}(M)$ has weak diameter at most $(2+1+2)+1+(2+1+2)=11$ in $\textbf{RIG}(G,\mathcal{H})$.
Now, we show for every $H\in\mathcal{H}$ that $I_{\mathcal{M}}(H)$ has weak diameter at most 9 in $\textbf{IM}(G,\mathcal{M})$. This is similar to the previous argument. Let $H\in\mathcal{H}$. Then for some $i\geqslant 0$, we have that $H\in L_{i}$. We may assume that $i\geqslant 3$, since otherwise we clearly have that $I_{\mathcal{M}}(H)$ has weak diameter at most $6$ in $\textbf{IM}(G,\mathcal{M})$. Similarly to before, by applying Menger’s theorem between $\bigcup_{0\leqslant j\leqslant i-3}A_{j}$ and the connected component of $\bigcup_{j\geqslant i-1}A_{j}$ that contains $H$, there exists a $2$-separation $(X,Y)$ of $G$ with $A_{i-3}\subseteq X$, $X\cap Y\subseteq A_{i-2}$, and $H\subseteq Y$. For each $x\in X\cap Y$, choose some $M_{x}\in\mathcal{M}_{i-2}$ with $x\in M_{x}$. Then in $\textbf{IM}(G,\mathcal{M})$, each vertex of $I_{\mathcal{M}}(H)$ is at distance at most 2 from some vertex of $\bigcup_{x\in X\cap Y}M_{x}$. Since $|X\cap Y|\leqslant 2$ and $\textbf{IM}(G,\mathcal{M})[I_{\mathcal{M}}(H)]$ is connected, it now follows that $I_{\mathcal{M}}(H)$ has weak diameter at most $(2+2)+1+(2+2)=9$ in $\textbf{IM}(G,\mathcal{M})$.
Hence $(G,\mathcal{M})$ is indeed a $(11,9)$-impression of $(G,\mathcal{H})$, as desired. ∎
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets in $G$, each of which contains a vertex incident to the outer face of $G$. For a collection disjoint connected sets $\mathcal{B}$ of $G$, we say that $(G,\mathcal{B})$ is a partial $(x,\infty)$-impression of $(G,\mathcal{H})$ if for every $B\in\mathcal{B}$, $I_{\mathcal{H}}(B)$ has weak diameter at most $x$ in $\textbf{RIG}(G,\mathcal{H})$.
We aim to find a $(x,y)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ in which each $M\in\mathcal{M}$ contains a vertex incident to the outer face of $G$. In light of Lemma˜3.2, we can find such a $(x,y)$-impression by first finding such a $(x^{\prime},\infty)$-impression and then refining this impression. The following theorem comes close to giving us such a $(x^{\prime},\infty)$-impression (not every $M\in\mathcal{M}$ contains a vertex incident to the outer face of $G$, but they are at distance at most $1$ from such a $M^{\prime}\in\mathcal{M}$ in $\textbf{IM}(G,\mathcal{M})$). The exact technical statement of this theorem is tailored to allow for an inductive argument on extending partial $(x^{\prime},\infty)$-impressions.
Theorem 3.3.
Let $G$ be a finite planar graph and $\mathcal{H}$ be a spanning collection of connected vertex sets in $G$, each of which contains a vertex incident to the outer face of $G$. Let $(G,\mathcal{B})$ be a partial $(30,\infty)$-impression of $(G,\mathcal{H})$ in which for every $B\in\mathcal{B}$, either $B$ contains a vertex incident to the outer face of $G$ and $I_{\mathcal{H}}(B)$ has weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$, or $B$ touches some $B^{\prime}\in\mathcal{B}$ where $B^{\prime}$ contains a vertex incident to the outer face of $G$, and for every maximal connected vertex set $K$ of $G\backslash(\bigcup_{B\in\mathcal{B}}B)$, either
  1. 1.
    $I_{\mathcal{H}}(K)$ has weak diameter at most $10$ in $\textbf{RIG}(G,\mathcal{H})$ and contains a vertex incident to the outer face of $G$, or $I_{\mathcal{H}}(K)$ has weak diameter at most $30$ in $\textbf{RIG}(G,\mathcal{H})$ and touches some $B\in\mathcal{B}$ that does contain a vertex incident incident to the outer face of $G$, or
  2. 2.
    $K$ contains a vertex incident to the outer face of $G$ and touches no $B\in\mathcal{B}$, or
  3. 3.
    $K$ contains a vertex incident to the outer face of $G$ and touches a unique $B\in\mathcal{B}$, or
  4. 4.
    $K$ contains a vertex incident to the outer face of $G$ and there exists distinct $B_{1},B_{2}\in\mathcal{B}$ satisfying the following
    • $K$ touches $B_{1}$ and $B_{2}$ and no other $\{B$$\in\mathcal{B}\backslash\{B_{1},B_{2}\}$,
    • $B_{1}$ and $B_{2}$ both contain a vertex incident to the outer face of $G$, and
    • $B_{1}$ and $B_{2}$ both contain a vertex incident to a common face $F$ distinct from the outer-face.
Then there is a $(30,\infty)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ such that for every $M\in\mathcal{M}$, either $M$ contains a vertex incident to the outer face of $G$ and $I_{\mathcal{H}}(B)$ has weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$, or $M$ touches some $M^{\prime}\in\mathcal{M}$ where $M^{\prime}$ contains a vertex incident to the outer face of $G$.
Доказательство.
We shall argue inductively on $|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$. Clearly the theorem holds if we have $V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)=\emptyset$. So, we may assume that there is some connected component of $G\backslash(\bigcup_{B\in\mathcal{B}}B)$, and we let $K$ be its vertex set.
Suppose first that $K$ satisfies Item˜1. Let $B^{*}=K$ and $\mathcal{B}^{*}=\mathcal{B}\cup\{B^{*}\}$. Clearly, $(G,\mathcal{B}^{*})$ satisfies the inductive hypothesis and $|V(G)\backslash(\bigcup_{B\in\mathcal{B}^{*}}B)|<|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$, so the theorem follows.
Now, suppose that $K$ satisfies Item˜2. Then simply choose some $b\in K$ incident to the outer face of $G$, let $B^{*}=\{b\}$ and $\mathcal{B}^{*}=\mathcal{B}\cup\{B^{*}\}$. Clearly, $(G,\mathcal{B}^{*})$ satisfies the inductive hypothesis and $|V(G)\backslash(\bigcup_{B\in\mathcal{B}^{*}}B)|<|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$, so the theorem follows.
[…]
Lastly, we can now assume that $K$ satisfies Item˜4. Let $B_{1},B_{2},F$ be as in Item˜4. Note that the intersection of $K$ with the vertices of the outer face of $G$ induces a connected graph since $B_{1}$ and $B_{2}$ both contain a vertex on the outer face of $G$ and the face $F$ (which is distinct from the outer face). Let $Q$ be the vertex set of this connected induced subgraph. Let $p_{1}p_{2}\ldots p_{t}$ be the walk in $G[Q]$ along the outer face of $G$ (and containing all of) $Q$, where $p_{1}$ touches $B_{1}$ and $p_{t}$ touches $B_{2}$. Choose $1\leqslant\ell\leqslant t$ maximum such that that there exists a path with vertex set $L$ between $p_{\ell}$ and $N_{G}(B_{1})$ such that $L$ is contained in $K\cap(H_{\ell}\cup H^{\prime}_{\ell})$ for some intersecting (but not necessarily distinct) $H_{\ell},H^{\prime}_{\ell}\in\mathcal{H}$. Similarly, choose $1\leqslant r\leqslant t$ minimum such that that there exists a path with vertex set $R$ between $p_{r}$ and $N_{G}(B_{2})$ such that $R$ is contained in $K\cap(H_{r}\cup H^{\prime}_{r})$ for some intersecting (but not necessarily distinct) $H_{r},H^{\prime}_{r}\in\mathcal{H}$. Note that such $\ell,r$ exist since $K$ touches exactly $B_{1},B_{2}$ from $\mathcal{B}$, and $K$ (and every $H\in\mathcal{H}$) contains a vertex incident to the outer face of $G$. Let $L^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(L)}H\right)\right]$ containing $L$, and similarly, let $R^{*}$ be the vertex set of the connected component of $G\left[K\cap\left(\bigcup_{H\in I_{\mathcal{H}}(R)}H\right)\right]$ containing $R$. Clearly $I_{\mathcal{H}}(L^{*})$ and $I_{\mathcal{H}}(R^{*})$ both have weak diameter at most 5 in $\textbf{RIG}(G,\mathcal{H})$. See Fig.˜3 for an illustration (in a case where $L^{*}$ and $R^{*}$ do not intersect). We split into two cases depending on whether or not $L^{*}$ and $R^{*}$ intersect.
Рис. 3: An illustration of the constructed connected sets $L^{*}$ and $R^{*}$ (coloured purple) in a case where they do not intersect. Both $B_{1}$ and $B_{2}$ are coloured blue. The connected sets in $\mathcal{H}$ are coloured red or orange, with the orange ones being $H_{\ell},H_{\ell}^{\prime},H_{r},H_{r}^{\prime}$. The two orange paths indicate $L$ and $R$.
[…]
We may now assume that $L^{*}$ and $R^{*}$ do not intersect. This in particular implies that $\ell<r$. Let $\mathcal{B}^{*}=\mathcal{B}\cup\{L^{*},R^{*}\}$. Consider the vertex set $K^{\prime}$ of a connected component of $G[K]\backslash(L^{*}\cup R^{*})$. Suppose that $K^{\prime}$ does not contain a vertex incident to the outer face of $G$. Then $K$ touches at least one of $B_{1},B_{2},L^{*},R^{*}$, and for every $H\in I_{\mathcal{H}}(K^{\prime})$, we have that $H$ intersects at least one of $B_{1},B_{2},L^{*},R^{*}$. Since $I_{\mathcal{H}}(B_{1}),I_{\mathcal{H}}(B_{2})$ both have weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$ and $I_{\mathcal{H}}(L^{*}),I_{\mathcal{H}}(R^{*})$ both have weak diameter at most 5 in $\textbf{RIG}(G,\mathcal{H})$, it follows that $I_{\mathcal{H}}(K^{\prime})$ has weak diameter at most 30 in $\textbf{RIG}(G,\mathcal{H})$. Thus in this case, $K^{\prime}$ satisfies Item˜1. So, we may now assume that $K^{\prime}$ does contain a vertex incident to the outer face of $G$. If $K^{\prime}$ does not touch both $L^{*}$ and $R^{*}$, then as before, $K^{\prime}$ satisfies either Item˜3 or Item˜4. So we assume now that $K^{\prime}$ touches both $L^{*}$ and $R^{*}$. Showing that $K^{\prime}$ satisfies Item˜4 will complete the proof.
Let $Q^{\prime}$ be the vertices $p_{i}$ of the walk $Q$ with $\ell<i<r$. Note that $Q^{\prime}$ is non-empty since $L^{*}$ and $R^{*}$ do not intersect (and $L^{*}$ contains both $p_{i}$ and $p_{i+1}$). Let $\mathcal{H}_{Q^{\prime}}=I_{\mathcal{H}}(Q^{\prime})$. Observe that by our choice of $L$ and $R$, there is no intersecting pair $H_{1}\in\mathcal{H}_{Q^{\prime}}$, $H_{2}\in\mathcal{H}$ such that there exists a path in $G[K\cap(H_{1}\cup H_{2})]$ between $Q^{\prime}$ and $N_{G}(B_{1}\cup B_{2})$. By our choice of $L^{*}$ and $R^{*}$, we therefore have that $I_{\mathcal{H}}(K^{\prime})\subseteq\mathcal{H}_{Q^{\prime}}$. In particular, $K^{\prime}$ does not touch $B_{1}$ or $B_{2}$. It now follows that there is some face $F$ of $G$ distinct from the outer face containing a vertex of $L^{*}$ and a vertex of $R^{*}$. Thus $K^{\prime}$ satisfies Item˜4.
So, $(G,\mathcal{B}^{*})$ satisfies the inductive hypothesis and $|V(G)\backslash(\bigcup_{B\in\mathcal{B}^{*}}B)|<|V(G)\backslash(\bigcup_{B\in\mathcal{B}}B)|$, so the theorem follows. ∎
Now we quickly refine the previous theorem so that every graph in $\mathcal{M}$ contains a vertex incident to the outer face of $G$.
Lemma 3.4.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$, each of which contains a vertex incident to the outer face of $G$. Then there is a $(70,\infty)$-impression $(G,\mathcal{M})$ of $(G,\mathcal{H})$ such that for every $\in\mathcal{M}, MM$ contains a vertex incident to the outer face of $G$.
Доказательство.
By Theorem˜3.3 (applied with $\mathcal{B}=\emptyset$), there is a $(30,\infty)$-impression $(G,\mathcal{M}^{\prime})$ of $(G,\mathcal{H})$ such that for every $M\in\mathcal{M}^{\prime}$, either $M$ contains a vertex incident to the outer-face of $G$ and $I_{\mathcal{H}}(B)$ has weak diameter at most 10 in $\textbf{RIG}(G,\mathcal{H})$, or $M$ touches some $M^{\prime}\in\mathcal{M}^{\prime}$ where $M^{\prime}$ contains a vertex incident to the outer face of $G$.
Let $\mathcal{M}_{1}$ be the elements $M$ of $\mathcal{M}^{\prime}$ that contain a vertex incident to the outer face of $G$, and let $\mathcal{M}_{2}$ be the elements $M$ of $\mathcal{M}^{\prime}$ that do not contain a vertex incident to the outer face of $G$. Then $\mathcal{M}^{\prime}=\mathcal{M}_{1}\cup\mathcal{M}_{2}$. Choose some function $f:\mathcal{M}_{2}\to\mathcal{M}_{1}$ such that $M$ touches $f(M)$ for every $M\in\mathcal{M}_{2}$. Let $\mathcal{M}=\{M\cup f^{-1}(M):M\in\mathcal{M}_{1})\}$. Then for every $M\in\mathcal{M}$, we have that $I_{\mathcal{H}}(M)$ has weak diameter at most $30+10+30=70$ in $G$. Then $(G,\mathcal{M})$ is a $(70,\infty)$-impression of $(G,\mathcal{H})$ such that for every $\in\mathcal{M}, MM$ contains a vertex incident to the outer face of $G$, as desired. ∎
We can now quickly prove Theorem˜3.1 using Lemma˜3.4 and Lemma˜3.2. We remark that in our proof of Theorem˜2.5, it actually turns out to be more convenient to use Lemma˜3.4 and Lemma˜3.2 separately as below instead of Theorem˜3.1. This is essentially because we end up needing to apply Lemma˜3.2 another time later anyway. However for the sake of completeness for outerstring graphs, we still complete the proof of Theorem˜3.1 here.
Proof of Theorem˜3.1..
By Lemma˜3.4, there is a $(70,\infty)$-impression $(G,\mathcal{M}^{\prime})$ of $(G,\mathcal{H})$ such that for every $\in\mathcal{M}^{\prime}, MM$ contains a vertex incident to the outer face of $G$. Let $G^{\prime}=\textbf{IM}(G,\mathcal{M}^{\prime})$. Then $G^{\prime}$ is outerplanar. For each $H\in\mathcal{H}$, let $H^{\prime}$ be the connected vertex set of $G^{\prime}$ that is exactly the union of all $M^{\prime}\in\mathcal{M}^{\prime}$ such that $M^{\prime}$ intersects $H$. Let $\mathcal{H}^{\prime}=\{H^{\prime}:H\in\mathcal{H}\}$.
By Lemma˜3.2, there is a $(11,9)$-impression $(G^{\prime},\mathcal{M}^{*})$ of $(G^{\prime},\mathcal{H}^{\prime})$. Let $\mathcal{M}=\{\bigcup_{M^{\prime}\in M^{*}}M^{\prime}:M^{*}\in\mathcal{M}\}$. Clearly for every $H\in\mathcal{H}$, $I_{\mathcal{M}}(H)$ has weak diameter at most $9$ in $\textbf{IM}(G,\mathcal{M})$. Also, for every $M\in\mathcal{M}$, we have that $I_{\mathcal{H}}(M)$ has weak diameter at most $11\cdot 70=770$ in $\textbf{RIG}(G,\mathcal{H})$.
Hence, $(G,\mathcal{M})$ is a $(770,9)$-impression of $(G,\mathcal{H})$, as desired. ∎
4 Encasings
The key to carrying out our inductive proof of Theorem˜2.5 in Section˜5 will be that we can find what we shall call a $(a,b,c,d)$-encasing. The aim of this section is to find these encasings. We show that we can find $(a,b,c,d)$-encasing by first finding a $(a,\infty,\infty,\infty)$-encasing, then refining it to a $(a,\infty,c,\infty)$-encasing, then further refining it to a $(a,\infty,c,d)$-encasing, and then finally refining it to a $(a,b,c,d)$-encasing. To ease into things, we delay the full definition of an $(a,b,c,d)$-encasing for now, and start simply with $(a,\infty,\infty,\infty)$-encasings.
Let $G$ be a finite planar graph and let $\mathcal{H}$ be a spanning collection of connected vertex sets in $G$. Let $\mathcal{B}$ be a collection of disjoint connected vertex sets in $G$. We say that $\mathcal{B}$ encircles $(G,\mathcal{H})$ if every $B\in\mathcal{B}$ contains a vertex incident to the outer face of $G$, and every $H\in\mathcal{H}$ that contains a vertex incident to the outer face of $G$ is contained in $\bigcup_{B\in\mathcal{B}}B$. We say that $\mathcal{B}$ is a $(a,\infty,\infty,\infty)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$.
We begin by finding a $(a,\infty,\infty,\infty)$-encasing. The lemma has an extra condition on the vertices contained in $\bigcup_{B\in\mathcal{B}}B$ that we will need later.
Lemma 4.1.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(210,\infty,\infty,\infty)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ with $\bigcup_{B\in\mathcal{B}}B=\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$, where $\mathcal{H}_{0}$ are the elements of $\mathcal{H}$ that contain a vertex incident to the outer face of $G$ and $\mathcal{H}_{1}$ are the elements of $\mathcal{H}\backslash\mathcal{H}_{0}$ that intersect a element of $\mathcal{H}_{0}$.
Доказательство.
Let $G^{*}$ be the subgraph of $G$ with edges $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}E(G[H])$ and no isolated vertices. Note that $G^{*}$ has the same outer face as $G$. For each $H\in\mathcal{H}_{1}$, choose some $H^{\prime}\in\mathcal{H}_{0}$ that intersects $H$ and let $H^{*}=H\cup H^{\prime}$. Let $\mathcal{H}^{*}=\mathcal{H}_{0}\cup\{H^{*}:H\in\mathcal{H}_{1}\}$. Then every vertex on the outer face of $G$ is contained in some element of $\mathcal{H}^{*}$. By Lemma˜3.4, there is a $(70,\infty)$-impression $(G^{*},\mathcal{B})$ of $(G^{*},\mathcal{H}^{*})$ such that every vertex of $G^{*}$ incident to the outer face of $G^{*}$ is contained in some $B\in\mathcal{B}$. In particular, every vertex of $G$ incident to the outer face of $G$ is contained in some $B\in\mathcal{B}$. Furthermore, observe that as $I_{\mathcal{H}^{*}}(B)$ has weak diameter at most 70 in $\textbf{RIG}(G^{*},\mathcal{H}^{*})$ for each $B\in\mathcal{B}$, we also have that $I_{\mathcal{H}^{*}}(B)$ has weak diameter at most $3\cdot 70=210$ in $\textbf{RIG}(G,\mathcal{H})$ for each $B\in\mathcal{B}$. ∎
Let $G$ be a finite planar graph, $\mathcal{H}$ a spanning collection of connected vertex sets of $G$, and let $\mathcal{B}$ be a collection of disjoint connected sets of $G$ that encircles $(G,\mathcal{H})$. Let $\mathcal{F}$ be the collection of non-empty vertex sets of $G$ that are maximal with respect to being contained in the interior of a face of $G[\bigcup_{B\in\mathcal{B}}B]$. We call these facial vertex sets. We remark that these facial vertex sets are not necessarily connected sets of $G$. Note that $\mathcal{B}\cup\mathcal{F}$ is a partition of $V(G)$. We denote by $G(\mathcal{B})$ the graph obtained from $\textbf{IM}(G,\mathcal{B})$ by for every facial vertex set $\mathcal{F}$ of $G[\bigcup_{B\in\mathcal{B}}B]$, we add $F$ as a vertex and make it adjacent to exactly the elements $B\in\mathcal{B}$ that contain a vertex incident to the face of $G[\bigcup_{B\in\mathcal{B}}B]$ that contains $F$ (we remark that $B$ and $F$ need not touch).
We say that some $H\in\mathcal{H}$ is $d$-caged by $\mathcal{B}$ if for every $u,v\in H$, there exists some sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant d$ such that
  • each $A_{i}$ is either a single element of $\mathcal{B}$ or a single element of $\mathcal{F}$ or the union of a subcollection $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$,
  • $u\in A_{0}$,
  • $v\in A_{t}$, and
  • for every $1\leqslant i<t$, we have that either $H$ intersects both $A_{i-1}$ and $A_{i}$, or neither $A_{i-1}$ or $A_{i}$ is an element of $\mathcal{F}$ and they touch.
See Fig.˜4 for an illustration of some $7$-caged $H$. We say that $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H)}$ if every $H\in\mathcal{H}$ is $d$-caged by $\mathcal{B}$.
Рис. 4: An illustration of some $H\in\mathcal{H}$ (coloured red) that is 7-caged by $\mathcal{B}$ (coloured blue). For $u,v\in H$, the vertex sets $A_{0},\ldots,A_{7}$ (coloured green) are as in the definition of being 7-caged. As an example illustrating the end of the proof of Lemma˜4.2, we would have that $A_{0}=A_{u}$, $A_{1}\in\mathcal{F}_{u}$, $A_{2},A_{5}\in\mathcal{A}_{u,v}$, $\{A_{3},A_{4}\}=\mathcal{V}_{(X,Y)}$, $A_{6}\in\mathcal{F}_{v}$, and $A_{7}=A_{v}$.
The motivation behind the notion of a $d$-cage is that it will allow us to ‘‘cut up’’ connected sets $H$ in such a way that each part is in some sense still close via the encircling $\mathcal{B}$. Being able to do this is crucial to the inductive argument of Theorem˜2.5.
We say that $\mathcal{B}$ is a $(a,\infty,\infty,d)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$, and
  • $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H})$.
In other words, $\mathcal{B}$ is a $(a,\infty,\infty,d)$-encasing of $(G,\mathcal{H})$ if it is both a $(a,\infty,\infty,\infty)$-encasing and a $d$-cage of $(G,\mathcal{H})$.
We will use our $(a,\infty,\infty,\infty)$-encasing to get a $(a,\infty,\infty,d)$-encasing. However, we again need this $(a,\infty,\infty,d)$-encasing to have some stronger structure on the vertices of $G$ that $\bigcup_{B\in\mathcal{B}}B$ contains. Let $\mathcal{H}_{0}$ be the elements of $\mathcal{H}$ that contain a vertex incident to the outer face of $G$, let $\mathcal{H}_{1}$ be the elements of $\mathcal{H}\backslash\mathcal{H}_{0}$ that intersect a element of $\mathcal{H}_{0}$, and similarly, let $\mathcal{H}_{2}$ be the elements of $\mathcal{H}\backslash(\mathcal{H}_{0}\cup\mathcal{H}_{1})$ that intersect a element of $\mathcal{H}_{1}$. We say that a $(a,\infty,\infty,d)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ is enforced if $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\subseteq\bigcup_{B\in\mathcal{B}}B\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2}}H$, and for every $v\in\left(\bigcup_{B\in\mathcal{B}}B\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, there is some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $\bigcup_{B\in\mathcal{B}}B$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$.
Lemma 4.2.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is an enforced $(840,\infty,\infty,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Доказательство.
We define $\mathcal{H}_{0},\mathcal{H}_{1},\mathcal{H}_{2}$ as above in the definition of an enforced $(a,\infty,\infty,d)$-encasing of $(G,\mathcal{H})$. By Lemma˜4.1, there is a $(210,\infty,\infty,\infty)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ with $\bigcup_{B\in\mathcal{B}}B=\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$. We will adjust this encasing to obtain the desired enforced $(840,\infty,\infty,7)$-encasing.
Let $G^{*}=G(\mathcal{B})$ as above, where $\mathcal{F}$ are the additional facial vertex sets. Choose some $B_{0}\in\mathcal{B}$. We say that a separation $(X,Y)$ of $G^{*}$ is 1-clean if $B_{0}\in X$, $X\cup Y=V(G^{*})$, $|X\cap Y|=1$, and subject to this, there is no other such $(X^{\prime},Y^{\prime})$ with $X^{\prime}\cap Y^{\prime}=X\cap Y$ and $|Y^{\prime}|>|Y|$. A separation $(X,Y)$ of $G^{*}$ is 2-clean if $B_{0}\in X$, $X\cup Y=V(G^{*})$, $X\cap Y$ has two elements and they are adjacent in $G^{*}$, and there is no 1-clean separation $(X^{*},Y^{*})$ with $X^{*}\cap Y^{*}\subset X\cap Y$ and $Y^{*}\subseteq Y$, and there is no other such separation $(X^{\prime},Y^{\prime})$ of $G^{*}$ with $X^{\prime}\cap Y^{\prime}=X\cap Y$ and $|Y^{\prime}|>|Y|$. By abuse of notation, we shall also call $(\{B_{0}\},V(G^{*}))$ a 1-clean separation. A separation $(X,Y)$ of $G^{*}$ is clean if it is 1-clean or 2-clean. Observe that if $(X,Y)$ is a clean separation, then $X\cap Y\subseteq\mathcal{B}$.
[…]
Рис. 5: A choice of $V^{1}_{(X,Y)},V^{2}_{(X,Y)}$ (coloured purple) as in the proof of Lemma˜4.2. Elements of $\mathcal{B}$ are coloured blue and elements of $\mathcal{H}$ are coloured red.
For a clean separation $(X,Y)$ of $G^{*}$, we let $H_{(X,Y)}$ be the vertices $v$ of $G$ contained in some vertex set of $\mathcal{F}\cap Y$ such that there exists some $H\in\mathcal{H}_{2}$ containing both $v$ and intersecting $\bigcup_{A\in X\cap Y}A$. If $(X,Y)$ is 1-clean, then we let $V_{(X,Y)}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$. Note that in this case, $V_{(X,Y)}$ is a connected set of $G$ and $I_{\mathcal{H}}(V_{X,Y})$ has weak diameter at most $3\cdot 210=630$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in V_{(X,Y)}\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=V_{(X,Y)}\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$ (or more precisely, the unique vertex set of $X\cap Y$). Otherwise, if $(X,Y)$ is 2-clean with $X\cap Y=\{B^{1},B^{2}\}$, then we choose disjoint $V_{(X,Y)}^{1}$, $V_{(X,Y)}^{2}$ such that $B^{1}\subseteq V_{(X,Y)}^{1}$, $B^{2}\subseteq V_{(X,Y)}^{2}$, $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}=H_{(X,Y)}\cup\bigcup_{R\in R_{(X,Y)}}R$, and both $V_{(X,Y)}^{1}$ and $V_{(X,Y)}^{2}$ are connected sets of $G$ (see Fig.˜5 for an illustration). Then $I_{\mathcal{H}}(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})$, and in particular both of $I_{\mathcal{H}}(V_{(X,Y)}^{1})$ and $I_{\mathcal{H}}(V_{(X,Y)}^{2})$ all have weak diameter at most $4\cdot 210=840$ in $\textbf{RIG}(G,\mathcal{H})$. Furthermore, if $v\in(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)=(V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2})\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, then there exists some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $V_{(X,Y)}^{1}\cup V_{(X,Y)}^{2}$, and intersects $B^{1}\cup B^{2}\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$.
Let $\mathcal{B}^{\prime}$ be the elements of $\mathcal{B}$ that do not intersect $V_{(X,Y)}$ for any $(X,Y)\in\mathcal{T}_{1}$ and do not intersect $V_{(X,Y)}^{1}$ or $V_{(X,Y)}^{2}$ for any $(X,Y)\in\mathcal{T}_{2}$. At last, let $\mathcal{B}^{*}=\mathcal{B}^{\prime}\cup\{V_{(X,Y)}:(X,Y)\in\mathcal{T}_{1}\}\cup\{V_{(X,Y)}^{1},V_{(X,Y)}^{2}:(X,Y)\in\mathcal{T}_{2}\}$.
Clearly, we have that $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\subseteq\bigcup_{B\in\mathcal{B}^{*}}B\subseteq\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2}}H$, and for every $B\in\mathcal{B}^{*}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $840$ in $\textbf{RIG}(G,\mathcal{H})$. We also have that for every $v\in\left(\bigcup_{B\in\mathcal{B}^{*}}B\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$, there is some connected set $H^{\prime}\subseteq H\in\mathcal{H}_{2}$ of $G$ that contains $v$, is contained in $\bigcup_{B\in\mathcal{B}^{*}}B$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$. To show that $\mathcal{B}^{*}$ provides the desired $(840,\infty,\infty,7)$-encasing of $(G,\mathcal{H})$, it now just remains to show that $\mathcal{B}^{*}$ is a $7$-cage of $(G,\mathcal{H})$.
If $H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}$, then $H$ is 0-caged by $\mathcal{B}^{*}$ since $H\subseteq\bigcup_{B\in\mathcal{B}^{*}}B$. If $H\in\mathcal{H}\backslash(\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2})$, then $H$ lies in some face of $G(\mathcal{B})$. In particular, $H$ intersects at most two elements of $\mathcal{B}^{*}$ and furthermore if $H$ does intersect two elements of $\mathcal{B^{*}}$ then they are touching, so it follows that $H$ is 3-caged by $\mathcal{B}^{*}$.
It remains to consider $H\in\mathcal{H}_{2}$. To do so, we must first introduce certain subsets $\mathcal{B}^{*}_{(X,Y)}$ and $\mathcal{F}^{*}_{(X,Y)}$ of $V(G(\mathcal{B}^{*}))$ according to clean separations $(X,Y)$ of $G^{*}$. Fix some $H\in\mathcal{H}_{2}$. For each 1-clean separation $(X,Y)\in\mathcal{T}_{1}$, let $\mathcal{V}_{(X,Y)}=\{V_{(X,Y)}\}$, and for each 2-clean separation $(X,Y)\in\mathcal{T}_{2}$, let $\mathcal{V}_{(X,Y)}=\{V^{1}_{(X,Y)},V_{(X,Y)}^{2}\}$. Let $\mathcal{F}^{*}$ be the facial vertex sets of $G(\mathcal{B}^{*})$. For each clean separation $(X,Y)\in\mathcal{T}$, let $\mathcal{F}^{*}_{(X,Y)}$ be the elements $F$ of $\mathcal{F}^{*}$ adjacent in $G(\mathcal{B}^{*})$ to a element of $\mathcal{V}_{(X,Y)}$. Now, let $\mathcal{B}^{*}_{(X,Y)}=N_{G(\mathcal{B}^{*})}(\mathcal{F}^{*}_{(X,Y)})$.
Consider some $H\in\mathcal{H}_{2}$. By our choice of $\mathcal{B}^{*}$, we have that there exists some clean separation $(X,Y)\in\mathcal{T}$ such that $H$ is entirely contained in $\bigcup_{B\in\mathcal{B}^{*}_{(X,Y)}}B\cup\bigcup_{F\in\mathcal{F}^{*}_{(X,Y)}}F$. Observe that $G(\mathcal{B}^{*})[\mathcal{B}^{*}_{(X,Y)}\cup\mathcal{F}^{*}_{(X,Y)}]$ is a planar graph in which the vertices $\mathcal{B}^{*}_{(X,Y)}$ are incident to the outer face and every $F\in\mathcal{F}^{*}_{(X,Y)}$ is adjacent to some $V\in\mathcal{V}_{(X,Y)}$, where $|\mathcal{V}_{(X,Y)}|\leqslant 2$ and if $|\mathcal{V}_{(X,Y)}|=2$, then the two vertices of $\mathcal{V}_{(X,Y)}$ are adjacent.
Consider some $u,v\in H$. Let $H_{u}$ be a subset of $H$ that contains $u$ and is maximal with respect to being a connected set of $G$ and being contained in $G[\bigcup_{B\in\mathcal{B}}B]$ (or equivalently $G[\bigcup_{B\in\mathcal{B}^{*}}B]$). Let $A_{u}$ be the union of all $B\in\mathcal{B}^{*}$ that intersect $H_{u}$. Let $\mathcal{F}_{u}$ be elements $F$ of $\mathcal{F}^{*}$ such that $H$ intersects both $F$ and $A_{u}$. We define $A_{v}$ and $\mathcal{F}_{v}$ similarly. Crucially, note that if $H$ intersects distinct $F_{1},F_{2}\in\mathcal{F}^{*}$, then $H$ intersects some $V\in\mathcal{B}$ incident to the face of $G(\mathcal{B}^{*})$ corresponding to $F_{1}$ that is also either contained in $\mathcal{V}_{(X,Y)}$ or adjacent to some element of $\mathcal{V}_{(X,Y)}$ in $G(\mathcal{B}^{*})$. Let $\mathcal{A}_{u,v}$ be the collection of unions of subcollections $\mathcal{B}^{\prime}\subseteq\mathcal{B}^{*}\backslash\mathcal{V}_{(X,Y)}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$, and some $B\in\mathcal{B^{\prime}}$ touches some $V\in\mathcal{V}_{(X,Y)}$. To see that $H$ is 7-caged by $\mathcal{B}^{*}$, it is now straight forward to choose the required sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant 7$ from the collection $\{A_{u}\}\cup\mathcal{F}_{u}\cup(\mathcal{A}_{u,v}\cup\mathcal{V}_{(X,Y)})\cup\mathcal{F}_{v}\cup\{A_{v}\}$ (here we take $A_{0}=A_{u}$, $A_{t}=A_{v}$, at most one $A_{i}$ is contained in $\mathcal{F}_{u}$ (this would be $A_{1}$ if so), at most one is contained in $\mathcal{F}_{v}$ (this would be $A_{t-1}$ if so), at most two are contained in $\mathcal{A}_{u,v}$, and at most two more are contained in $\mathcal{V}_{(X,Y)}$). For an illustrative example, see Fig.˜4.
So, $\mathcal{B}^{*}$ is indeed a 7-cage of $(G,\mathcal{H})$, and thus $\mathcal{B}^{*}$ is an enforced $(840,\infty,\infty,7)$-encasing of $(G,\mathcal{H})$, as desired. ∎
Let $G$ be a finite planar graph, $\mathcal{H}$ a spanning collection of connected vertex sets of $G$, and let $\mathcal{B}$ be a collection of disjoint connected sets of $G$ such that $\mathcal{B}$ encircles $(G,\mathcal{H})$. Let $\mathcal{R}$ be the interiors of the internal faces of $G[\bigcup_{B\in\mathcal{B}}B]$. For $u\in\bigcup_{B\in\mathcal{B}}B$ and $H\in\mathcal{H}$, we say that $u$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ if either $u\in H$ or $u$ is contained in one of the bounded regions of $\mathbb{R}^{2}\backslash(G[H]\cup\bigcup_{R\in\mathcal{R}}R)$. We say that $U\subseteq\bigcup_{B\in\mathcal{B}}B$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ if each $u\in U$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$. Let $F$ be the vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$. We say that a vertex $v\in F$ is $\mathcal{B}$-encroached by some $H\in\mathcal{H}$ with $v\notin H$ if there exists some connected set $U\subseteq\bigcup_{B\in\mathcal{B}}B$ of $G$ and edges $u_{1}v_{1},\ldots,u_{r}v_{r}$ between $F$ and $\bigcup_{B\in\mathcal{B}}B$ that appear consecutively along the boundary of the face of $G[\bigcup_{B\in\mathcal{B}}B]$ corresponding to $F$, such that
  • $U$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$,
  • for each $1\leqslant i<r$, there is a facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$,
  • $u_{1},v_{1},u_{r},v_{r}\in H$, and
  • $v\in\{v_{1},\ldots,v_{r}\}$.
See Fig.˜6 for an illustrative example. We further say that some $H^{\prime}\in\mathcal{H}$ is $\mathcal{B}$-encroached by some $H\in\mathcal{H}$ if there is some set $F$ of vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$ and some vertex $v\in H^{\prime}\cap F$ such that $v$ is $\mathcal{B}$-encroached by $H$.
Рис. 6: An encircling $\mathcal{B}$ (in blue) and some (red) $H\in\mathcal{H}$. The vertices of the facial vertex set $F$ that are encroached by $H$ are coloured orange.
Let $\mathcal{H}_{0}$ be the connected sets in $\mathcal{H}$ that contain a vertex incident to the outer face of $G$. We say that a collection of disjoint connected sets $\mathcal{B}$ of $G$ that encircles $(G,\mathcal{H})$ is a $c$-surround of $(G,\mathcal{H})$ if
  • for every pair $H_{u},H\in\mathcal{H}$ such that $H$ intersects but is not contained in $\bigcup_{B\in\mathcal{B}}B$, and $H_{u}$ contains a vertex $u\in\bigcup_{B\in\mathcal{B}}B$ that is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$, we have that $H_{u}$ is at distance at most c1c-1 from $H$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$, and
  • for every pair $H^{\prime},H\in\mathcal{H}$ such that $H$ encroaches $H^{\prime}$, we have that $H^{\prime}$ is at distance at most $c$ from $H$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$.
Note that since $\mathcal{B}$ encircles $(G,\mathcal{H})$, if $H$ encroaches $H^{\prime}$, then $H\not\in\mathcal{H}_{0}$.
The motivation behind the notion of a $c$-surround is that it will allow us to add some (not necessarily yet connected until more edges are added to $G$) sets $H^{\prime}$ to $\mathcal{H}$ such that $I_{\mathcal{H}}(H^{\prime})$ has bounded weak diameter in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ and $H^{\prime}$ is disjoint from $\bigcup_{B\in\mathcal{B}}B$. Such $H^{\prime}$ will contain for instance the vertices encroached by a given $H\in\mathcal{H}$ and will be crucial for the inductive argument of Theorem˜2.5 in which we remove the vertices of $\bigcup_{B\in\mathcal{B}}B$.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Let $\mathcal{B}$ be a collection of disjoint connected subgraphs of $G$. We say that $\mathcal{B}$ is a $(a,\infty,c,d)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$,
  • $\mathcal{B}$ is a $c$-surround of $(G,\mathcal{H})$, and
  • $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H})$.
The enforced encasing found in Lemma˜4.2 is actually a 4-surround (which was the purpose of the technical definition of enforced encasings).
Lemma 4.3.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(840,\infty,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Доказательство.
By Lemma˜4.2, there is an enforced $(840,\infty,\infty,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$. Let $\mathcal{H}_{0},\mathcal{H}_{1},\mathcal{H}_{2}$ be as in the definition of an enforced encasing, and furthermore, let $\mathcal{H}_{3}$ be the elements of $\mathcal{H}\backslash(\mathcal{H}_{0}\cup\mathcal{H}_{1}\cup\mathcal{H}_{2})$ that intersect an element of $\mathcal{H}_{2}$. We will show that $\mathcal{B}$ is in fact a $(840,\infty,4,7)$-encasing of $(G,\mathcal{H})$.
First, we show the first bullet in the definition of being a 4-surround. Consider some pair $H_{u},H^{*}\in\mathcal{H}$ such that $H^{*}$ intersects but is not contained in $\bigcup_{B\in\mathcal{B}}B$, and $H_{u}$ contains a vertex $u\in\bigcup_{B\in\mathcal{B}}B$ that is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$. Clearly $H^{*}\in\mathcal{H}_{2}\cup\mathcal{H}_{3}$, and therefore $u\not\in\bigcup_{H\in{\mathcal{H}_{0}}}H$. In particular, $H_{u},H^{*}\not\in\mathcal{H}_{0}$.
Suppose first that $u\in\left(\bigcup_{H\in\mathcal{H}_{1}}H\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}}H\right)$. Then, there is some $H_{u}^{\prime}\in\mathcal{H}_{1}$ that contains $u$ (possibly $H_{u}^{\prime}$ is not distinct from $H_{u}$). Observe that $H_{u}^{\prime}$ must intersect $H^{*}$ since $u$ is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$, $H_{u}\in\mathcal{H}_{1}$, and $H^{*}\in\mathcal{H}_{2}\cup\mathcal{H}_{3}$. Therefore, $H_{u}$ is at distance at most 2 from $H^{*}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$.
So, we may assume now that $u\in\left(\bigcup_{B\in\mathcal{B}}B\right)\backslash\left(\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H\right)$. Then, by the definition of an enforced encasing, there is some connected set $P_{u}\subseteq H_{u}^{\prime}\in\mathcal{H}_{2}$ of $G$ that contains $u$, is contained in $\bigcup_{B\in\mathcal{B}}B$, and intersects $\bigcup_{H\in\mathcal{H}_{0}\cup\mathcal{H}_{1}}H$. Choose $H_{u}^{\prime\prime}\in\mathcal{H}_{1}$ that intersects $P_{u}$ (and thus also $H_{u}^{\prime}$). Similarly to before, $H_{u}^{\prime\prime}\cup P_{u}$ must intersect $H^{*}$ since $u$ is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$, $u\in H_{u}^{\prime\prime}\cup P_{u}\subseteq\bigcup_{B\in\mathcal{B}}B$, $H_{u}^{\prime\prime}\in\mathcal{H}_{1}$, and $H^{*}\in\mathcal{H}_{2}\cup\mathcal{H}_{3}$. In particular, $H^{*}$ intersects at least one of $H_{u}^{\prime\prime}$ or $H_{u}^{\prime}$. Therefore, $H_{u}$ is at distance at most 3 from $H^{*}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$. Thus, $\mathcal{B}$ satisfies the first bullet of being a 4-surround.
Suppose that some $H^{\prime}\in\mathcal{H}$ is $\mathcal{B}$-encroached by some $H^{*}\in\mathcal{H}$. Clearly we have that $H^{\prime},H^{*}\not\in\mathcal{H}_{0}$. Note that $H^{*}$ intersects but is not contained in $\bigcup_{B\in\mathcal{B}}B$. By definition, there is some set $F$ of vertices of $G$ contained in some face of $G[\bigcup_{B\in\mathcal{B}}B]$ and some vertex $v\in H^{\prime}\cap F$ such that $v$ is $\mathcal{B}$-encroached by $H^{*}$. Let $u$ be a neighbour of $v$ in $\bigcup_{B\in\mathcal{B}}B$ that is bounded by $H^{*}$ in $G[\bigcup_{B\in\mathcal{B}}B]$, and let $H_{u}$ be some connected set in $\mathcal{H}$ that contains $\{u,v\}$ (possibly $H_{u}$ is not distinct from $H^{\prime}$). Then, $H^{\prime}$ is at distance at most $1+3=4$ from $H^{*}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$.
Thus, $\mathcal{B}$ is a 4-surround of $(G,\mathcal{H})$, and in particular, a $(840,\infty,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$ as desired. ∎
We are finally ready to give the entire definition of a $(a,b,c,d)$-encasing. We say that $\mathcal{B}$ is a $(a,b,c,d)$-encasing of $(G,\mathcal{H})$ if $\mathcal{B}$ encircles $(G,\mathcal{H})$, and
  • for every $B\in\mathcal{B}$, we have that $I_{\mathcal{H}}(B)$ has weak diameter at most $a$ in $\textbf{RIG}(G,\mathcal{H})$,
  • for every connected set $H^{*}\subseteq H\in\mathcal{H}$ of $G$ such that $H^{*}$ is contained in $\bigcup_{B\in\mathcal{B}}B$, we have that $I_{\mathcal{B}}(H^{*})$ has weak diameter at most $b$ in $\textbf{IM}(G,\mathcal{B})$,
  • $\mathcal{B}$ is a $c$-surround of $(G,\mathcal{H})$, and
  • $\mathcal{B}$ is a $d$-cage of $(G,\mathcal{H})$.
We can at last complete the refinement of the encasing with the use of Lemma˜3.2.
Lemma 4.4.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(9240,9,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Доказательство.
By Lemma˜4.3, there is a $(840,\infty,4,7)$-encasing $\mathcal{B}^{\prime}$ of $(G,\mathcal{H})$. Let $G^{\prime}=\textbf{IM}(G,\mathcal{B}^{\prime})$. Then $G^{\prime}$ is outerplanar. For every connected set $H^{*}\subseteq H\in\mathcal{H}$ of $G$ such that $H^{*}$ is contained in $\bigcup_{B\in\mathcal{B}^{\prime}}B$, let $H^{\prime}$ be the connected set of $G^{\prime}$ whose vertex set is exactly the $B\in\mathcal{B}^{\prime}$ such that $H^{*}$ intersects $B$. Let $\mathcal{H}^{\prime}$ be the collection of all of these connected vertex sets of $G^{\prime}$.
By Lemma˜3.2, there is a $(11,9)$-impression $(G^{\prime},\mathcal{M})$ of $(G^{\prime},\mathcal{B}^{\prime})$. Let $\mathcal{B}=\{\bigcup_{B\in M}B:M\in\mathcal{M}\}$. Clearly for every $H\in\mathcal{H}$, $I_{\mathcal{M}}(H)$ has weak diameter at most $9$ in $\textbf{IM}(G,\mathcal{M})$. Also, for every $M\in\mathcal{M}$, we have that $I_{\mathcal{H}}(M)$ has weak diameter at most $11\cdot 840=9240$ in $\textbf{RIG}(G,\mathcal{H})$.
Hence, $\mathcal{B}$ is a $(9240,9,4,7)$-encasing of $(G,\mathcal{H})$, as desired. ∎
5 String graphs
Lemma˜4.4 is our main technical tool for proving Theorem˜2.5. We still need to introduce the notion of fortifications in order to give the precise technical statement that will allow for the inductive proof using Lemma˜4.4.
[…]
Рис. 7: A fortification $G^{\prime}$ of planar graph $G$ (which is the subgraph on the black edges. The arrows orthogonal to the additional red edges indicate their sides.
Let $G$ be a planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets $\mathcal{H}$ of $G$. Let $\mathcal{H}_{0}$ be the collection of connected sets $H\in\mathcal{H}$ of $G$ that contain a vertex incident to the outer face of $G$. we say that $(G^{\prime},\mathcal{H^{\prime}})$ is a $k$-fortification of $(G,\mathcal{H})$ if
  • $G^{\prime}$ is a fortification of $G$,
  • $\mathcal{H}^{\prime}$ spans $G^{\prime}$,
  • $\mathcal{H}^{\prime}$ can be obtained from $\mathcal{H}$ by adding connected sets $H^{\prime}$ of $G$ such that each $I_{\mathcal{H}}(H^{\prime})$ has weak diameter at most $k$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$, and
[…]
We use the following technical statement to facilitate an inductive argument making use of Lemma˜4.4.
Theorem 5.1.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(9242,80)$-impression $(G^{\prime},\mathcal{M})$ of a $8$-fortification $(G^{\prime},\mathcal{H}^{\prime})$ of $(G,\mathcal{H})$ such that for every $M\in\mathcal{M}$ containing a vertex incident to the outer face of $G$, we have that $I_{\mathcal{H^{\prime}}}(M)$ has weak diameter at most $9240$ in $\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})$, and for every $H\in\mathcal{H}$ containing a vertex incident to the outer face of $G$, we have that $I_{\mathcal{M}}(H)$ has weak diameter at most $9$ in $\textbf{IM}(G^{\prime},\mathcal{M})$.
Доказательство.
We shall argue inductively on $|V(G)|$ (trivially the theorem holds for small $|V(G)|$). Let $\mathcal{H}_{0}$ be the collection of connected sets $H\in\mathcal{H}$ that contain a vertex incident to the outer face of $G$. By Lemma˜4.4, there is a $(9240,9,4,7)$-encasing $\mathcal{B}$ of $(G,\mathcal{H})$.
Let $\mathcal{F}$ be the facial vertex sets of $G(\mathcal{B})$. Consider some $F\in\mathcal{F}$ and $H\in\mathcal{H}$ such that $H$ intersects both $\bigcup_{B\in\mathcal{B}}B$ and $F$. Let $R_{F,H}^{\prime}$ be the vertices of $F$ that are encroached by $H$, and let $R_{F,H}=R_{F,H}^{\prime}\cup(F\cap H)$. Consider some $H^{\prime}\in\mathcal{H}$ intersecting $R_{F,H}$ but not $H$. Then, $H^{\prime}$ intersects $R_{F,H}^{\prime}$. Since $\mathcal{B}$ is a $4$-surround of $(G,\mathcal{H})$, the distance between $H^{\prime}$ and $H$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ is at most 4. Therefore, $I_{\mathcal{H}}(R_{F,H})$ has weak diameter at most $2\cdot 4=8$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$. We wish to add edges to $G$ so that $R_{F,H}$ is a connected set.
Consider some $v\in R_{F,H}$. By definition of being encroached, there exists connected set $U\subseteq\bigcup_{B\in\mathcal{B}}B$ of $G$ and edges $u_{1}v_{1},\ldots,u_{r}v_{r}$ between $F$ and $\bigcup_{B\in\mathcal{B}}B$ that appear consecutively along the boundary of the face of $G[\bigcup_{B\in\mathcal{B}}B]$ corresponding to $F$, such that
  • $U$ is bounded by $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$,
  • for each $1\leqslant i<r$, there is a facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$,
  • $u_{1},v_{1},u_{r},v_{r}\in H$, and
  • $v\in\{v_{1},\ldots,v_{r}\}$.
Clearly $\{v_{1},\ldots,v_{r}\}\subseteq R_{F,H}$. Let $E_{F,H}$ be the collection of such edges $v_{i}v_{i+1}$ (across each $v\in R_{F,H}$) that are not already an edge of $G$. For each such $v_{i}v_{i+1}\in E_{F,H}$, let $F_{v_{i}v_{i+1}}$ be the side of $v_{i}v_{i+1}$ drawn in the corresponding face of $G[\bigcup_{B\in\mathcal{B}}B]$ that contains the facial walk of $G$ between $u_{i}$ and $u_{i+1}$ that is contained in $U$ (note that the vertices of this facial walk are bounded $H$ in $G[\bigcup_{B\in\mathcal{B}}B]$ and the only two other vertices incident to this face are $v_{i},v_{i+1}$). Since $\mathcal{B}$ is a 4-surround of $(G,\mathcal{H})$, it follows that for every $H^{\prime}\in\mathcal{H}$ that contains a vertex incident to $F_{v_{i}v_{i+1}}$, the distance between $H$ and $H^{\prime}$ in $\textbf{RIG}(G,\mathcal{H}\backslash\mathcal{H}_{0})$ is at most $4$. This in particular implies that $F_{v_{i}v_{i+1}}$ will satisfy the 4th bullet in the definition of a $8$-fortification of $(G,\mathcal{H})$. This in particular implies that no vertex incident to $F_{v_{i}v_{i+1}}$ is incident to the outerface of $G$. Then observe that $R_{F,H}$ is a connected set in $G\cup E_{F,H}$. For an illustration of such a $R_{F,H},E_{F,H}$, see Fig.˜8. Let $G^{*}$ be the planar graph obtained from adding these edges from the union of all such $E_{F,H}$. Let $\mathcal{R}$ be the collection of all such $R_{F,H}$. Then every element of $\mathcal{R}$ is a connected set of $G^{*}$.
Рис. 8: The (orange) vertex set $R_{F,H}$ given the (blue) encasing $\mathcal{B}$, the (red) $H\in\mathcal{H}$, and $F\in\mathcal{F}$. The orange vertices are the vertices of $R_{F,H}^{\prime}$. The pink edges are $E_{F,H}$, and the arrows orthogonal to them indicate their sides.
Let $\mathcal{H}_{C}$ be the elements $H$ of $\mathcal{H}$ that are contained in $\bigcup_{B\in\mathcal{B}}B$, and let $\mathcal{H}_{I}$ be the elements $H$ of $\mathcal{H}$ that intersect but are not contained in $\bigcup_{B\in\mathcal{B}}B$. Let $\mathcal{R}^{*}=(\mathcal{H}\backslash(\mathcal{H}_{C}\cup\mathcal{H}_{I}))\cup\mathcal{R}$. Let $G_{0}^{*}=G^{*}\backslash\left(\bigcup_{B\in\mathcal{B}}B\right)$, and let $\mathcal{R}_{0}^{*}$ be the elements of $\mathcal{R}^{*}$ that contain a vertex incident to the outer face of $G_{0}$. Note that $\mathcal{R}^{*}$ is a spanning collection of connected vertex sets of $G_{0}$ and that $\mathcal{R}\subseteq\mathcal{R}^{*}_{0}$.
[…]
If $M\in\mathcal{M}^{\prime}\subseteq\mathcal{M}$ does not contain a vertex incident to the outer face of $G_{0}^{\prime}$, then $I_{\mathcal{H}^{\prime}}(M)=I_{\mathcal{R}^{\prime}}(M)$ has weak diameter at most 9242 in $\textbf{RIG}(G_{0}^{\prime},\mathcal{R}^{\prime})$, and therefore also in $\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})$. If $M\in\mathcal{M}^{\prime}\subseteq\mathcal{M}$ contains a vertex incident to the outer face of $G_{0}^{\prime}$, then $I_{\mathcal{R}^{\prime}}(M)$ has weak diameter at most 9240 in $\textbf{RIG}(G_{0}^{\prime},\mathcal{R}^{\prime})$. Since $I_{\mathcal{H}^{\prime}}(M)\subseteq I_{\mathcal{R}^{\prime}}(M)\cup N_{\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})}(I_{\mathcal{R}^{\prime}}(M))$, it follows that $I_{\mathcal{H}^{\prime}}(M)$ has weak diameter at most $9240+2=9242$ in $\textbf{RIG}(G^{\prime},\mathcal{H}^{\prime})$. So, $(G^{\prime},\mathcal{M})$ is a $(9242,\infty)$-impression of $(G^{\prime},\mathcal{H}^{\prime})$.
If $H\in\mathcal{R}^{\prime}\subseteq\mathcal{H}^{\prime}$, then $I_{\mathcal{M}^{\prime}}(H)=I_{\mathcal{M}}(H)$ has weak diameter at most 80 in $\textbf{IM}(G_{0}^{\prime},\mathcal{M}^{\prime})$ and therefore also in $\textbf{IM}(G^{\prime},\mathcal{M})$. To show that $(G^{\prime},\mathcal{M}^{\prime})$ is a $(9242,80)$-impression of $(G^{\prime},\mathcal{H}^{\prime})$, it now just remains to show that if $H\in\mathcal{H}_{I}$, then $I_{\mathcal{M}}(H)$ has weak diameter at most $80$ in $\textbf{IM}(G^{\prime},\mathcal{M})$.
Fix some $H\in\mathcal{H}_{I}$. Consider some $u,v\in H$ and let $M_{u},M_{v}\in\mathcal{M}$ be such that $u\in M_{u}$ and $v\in M_{v}$. Since $\mathcal{B}$ is a 7-cage of $(G,\mathcal{H})$, there exists some sequence of vertex sets $A_{0},\ldots,A_{t}$ with $t\leqslant 7$ such that
  • each $A_{i}$ is either a single element of $\mathcal{B}$ or a single element of $\mathcal{F}$ or the union of a subcollection $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ such that some connected set $H^{*}\subseteq H$ of $G$ (or equivalently $G^{\prime}$) intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]=G^{\prime}[\bigcup_{B\in\mathcal{B}^{\prime}}B]$,
  • $u\in A_{0}$,
  • $v\in A_{t}$, and
  • for every $1\leqslant i<t$, we have that either $H$ intersects both $A_{i-1}$ and $A_{i}$, or neither $A_{i-1}$ or $A_{i}$ is an element of $\mathcal{F}$ and they touch.
If $A_{i}\in\mathcal{B}\subseteq\mathcal{M}$, then clearly $\{A_{i}\}$ has weak diameter 0 in $\textbf{IM}(G^{\prime},\mathcal{M})$. If $A_{i}$ is the union of a subcollection $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ such that some connected set $H^{*}\subseteq H$ of $G$ intersects each member of $\mathcal{B}^{\prime}$ and is contained in $G[\bigcup_{B\in\mathcal{B}^{\prime}}B]$, then it follows from $\mathcal{B}$ being a $(9240,9,4,7)$-encasing of $(G,\mathcal{H})$ that $A_{i}$ has weak diameter at most $9$ in $\textbf{IM}(G,\mathcal{B})$ and therefore in $\textbf{IM}(G^{\prime},\mathcal{M})$. If $A_{i}$ is a single element $F$ of $\mathcal{F}$, then $R_{F,H}$ contains $H\cap F$. As $R_{F,H}\in\mathcal{R}_{0}^{*}$, we have that $I_{\mathcal{M^{\prime}}}(R_{F,H})$ has weak diameter at most 9 in $\textbf{IM}(G_{0},\mathcal{M}^{\prime})$, and therefore $I_{\mathcal{M}}(R_{F,H})$ has weak diameter at most 9 in $\textbf{IM}(G^{\prime},\mathcal{M})$. As $t\leqslant 7$, it now follows that $d_{\textbf{IM}(G^{\prime},\mathcal{M})}(M_{u},M_{v})\leqslant 7+8\cdot 9=80$. Therefore, $I_{\mathcal{M}}(H)$ has weak diameter at most $80$ in $\textbf{IM}(G^{\prime},\mathcal{M})$.
Hence $(G^{\prime},\mathcal{M})$ is a $(9242,80)$-impression of $(G^{\prime},\mathcal{H}^{\prime})$, as desired. ∎
By Theorem˜5.1 and Lemma˜2.2, we obtain the following.
Theorem 5.2.
Let $G$ be a finite planar graph and $\mathcal{H}$ a spanning collection of connected vertex sets of $G$. Then there is a $(73936,80)$-impression $(G^{\prime},\mathcal{M})$ of $(G^{\prime},\mathcal{H})$, where $G^{\prime}$ is a fortification of $G$ and $\mathcal{H}$ 88-spans $G^{\prime}$.
Since for every finite string graph $S$ there is a finite planar graph $G$ and a spanning collection $\mathcal{H}$ of connected vertex sets of $G$ such that $\textbf{RIG}(G,\mathcal{H})=S$ [63], by Theorem˜5.2 and Proposition˜2.1, we finally get that string graphs are quasi-isometric to planar graphs.
Theorem 5.3.
Let $S$ be a finite string graph. Then there is a planar graph $G$ and a function $f:V(S)\to V(G)$ such that for all $u,v\in V(S)$,
$\frac{1}{73944}d_{S}(u,v)-\frac{73936}{73944}\leqslant d_{G}(f(u),f(v))\leqslant 160d_{S}(u,v),$
and for every $v\in V(G)$ there exists some $u\in V(S)$ such that $d_{G}(f(u),v)\leqslant 80$.
Theorem˜2.5 (and thus Theorem˜1.1) then finally follows from Theorem˜5.3 and Lemma˜2.3.
6 Riemannian surfaces
In this section, we prove Theorem˜1.7.
In addition to Theorem˜5.2, our main tool is the following recent theorem of Georgakopoulos and Vigolo [48]. (We delay giving certain definitions on Riemannian surfaces until after the proof of our last required graph theoretic lemma (which is Lemma˜6.4).)
Theorem 6.1 (Georgakopoulos, Vigolo [48]).
Let $\Sigma$ be a complete Riemannian surface. Then $\Sigma$ admits a piecewise smooth triangulation $H\subset\Sigma$ by a locally finite $(0,1]$-metric graph $H$ such that $H\hookrightarrow\Sigma$ is a quasi-isometry.
Roughly speaking, Theorem˜6.1 allows us to just deal with modifying some triangulation by a metric graph to obtain a quasi-isometric triangulation by a simplicial graph as in Theorem˜1.7.
We begin with considering metric graphs in the plane. A graph is an internally triangulated planar graph if it has an embedding in the plane in which every face except possibly the outerface is a triangle. Observe that in the case of internally triangulated planar graphs, Theorem˜5.2 simplifies to directly finding an impression of $(G,\mathcal{H})$ since in this case $G^{\prime}=G$. We require a version more directly applicable to metric graphs. We obtain this from Theorem˜5.2 with essentially the same trick as in Lemma˜2.4 and using Proposition˜2.1.
Lemma 6.2.
Let $H$ be a finite $(0,1]$-metric internally triangulated planar graph. Then there exists a partition of $V(H)$ into connected vertex sets $\mathcal{M}$ such that for every $u,v\in V(H)$, we have that
$\frac{1}{225000}d_{H}(u,v)-1\leqslant d_{\textbf{IM}(H,\mathcal{M})}(f(u),f(v))\leqslant 160d_{H}(u,v)+160,$
where $f(u)=M$ for every $u\in M\in\mathcal{M}$.
Доказательство.
Similarly to in the proof of Lemma˜2.4, for each $u\in V(H)$, we let $S_{u}$ be the set of vertices of $H$ at distance at most $\frac{3}{2}$ from $u$. Let $\mathcal{S}=\{S_{u}:u\in V(H)\}$. Clearly $\mathcal{S}$ is a spanning collection of connected vertex sets of $H$.
Consider some $u,v\in V(H)$ and let $P$ be a geodesic path from $u$ to $v$ in $H$. Choose vertices $x_{0},\ldots,x_{t}$ of $H$ along $P$ so that $x_{0}=u$, $x_{t}=v$ and for each $0\leqslant i<t$, the distance along $P$ between $x_{i}$ and $x_{i+1}$ is in $[1,2]$ (or if $d_{H}(u,v)<1$, we just take $x_{0}=u$, $x_{1}=v$). So, $\lceil\frac{1}{2}d_{H}(u,v)\rceil\leqslant t\leqslant\lfloor d_{H}(u,v)\rfloor$. For each $0\leqslant i<t$, we have that $S_{x_{i}}$ intersects $S_{x_{i+1}}$ since $H$ is a $(0,1]$-metric graph and $d_{H}(x_{i},x_{i+1})+1\leqslant 3=2\cdot\frac{3}{2}$. It therefore follows that for every pair $u,v\in V(H)$, we have that
$\frac{1}{3}d_{H}(u,v)\leqslant d_{\textbf{RIG}(H,\mathcal{S})}(S_{u},S_{v})\leqslant d_{H}(u,v)+1.$
By Theorem˜5.2, there is a $(73936,80)$-impression $(H,\mathcal{M})$ of $(H,\mathcal{S})$. By Proposition˜2.1, it then follows that for every $u,v\in V(H)$, we have that
$\frac{1}{73944}d_{\textbf{RIG}(H,\mathcal{S})}(S_{u},S_{v})-\frac{73936}{73944}\leqslant d_{\textbf{IM}(H,\mathcal{M})}(f(u),f(v))\leqslant 160d_{\textbf{RIG}(H,\mathcal{S})}(S_{u},S_{v}).$
Therefore, for every $u,v\in V(H)$, we have that
$\frac{1}{225000}d_{H}(u,v)-1\leqslant d_{\textbf{IM}(H,\mathcal{M})}(f(u),f(v))\leqslant 160d_{H}(u,v)+160,$
as desired. ∎
Next we extend Lemma˜6.2 to triangulations of higher genus surfaces. A graph triangulates a surface without boundary if it has an embedding in which every face is a non-degenerate triangle (meaning it is bounded by three distinct vertices and edges) and is homeomorphic to a disk. In extending Lemma˜6.2, we just get a some extra additive distortion which is linear in the Euler genus of the surface.
Lemma 6.3.
Let $H$ be a finite $(0,1]$-metric graph that triangulates a surface $\Sigma$ without boundary and with Euler genus $g$. Then there exists a partition of $V(H)$ into connected vertex sets $\mathcal{M}$ such that for every $u,v\in V(H)$, we have that
$\frac{1}{225000}d_{H}(u,v)-1\leqslant d_{\textbf{IM}(H,\mathcal{M})}(f(u),f(v))\leqslant 160d_{H}(u,v)+160+1280g,$
where $f(u)=M$ for every $u\in M\in\mathcal{M}$.
Доказательство.
Choose a vertex $r$ of $H$ and construct a spanning tree $T$ of $H$ with root $r$ by ordering the vertices of $H$ according to their distance from $r$ and adding them and some incident edge greedily to the tree so that for each $x\in V(H)$, we have that $d_{T}(r,x)=d_{H}(r,x)$.
In the dual of the embedding of $H^{\prime}$ in $\Sigma$ (defined as the graph $H^{*}$ with a vertex in each face, the same edge set as $H$ and with each edge $e$ in $H^{*}$ being incident to the two vertices corresponding to the two faces that $e$ is incident to in $H$, (possibly allowing loops or multiple edges)), choose some spanning tree $T^{*}$ with $E(T)\cap E(T^{*})=\emptyset$. Let $F$ be the edges of $H$ not in $T$ or $T^{*}$. By Euler’s formula, we have that $|F|=g$. Let $T^{\prime}$ be the minimal subtree of $T$ containing the ends of the edges of $F$. Note that $T^{\prime}$ is the union of at most $2g$ geodesics of $H$. Now observe that $G=H-V(T^{\prime})$ is a $(0,1[$-metric internally triangulated planar graph.
By Lemma˜6.2, there is a partition of $V(G)$ into connected vertex sets $\mathcal{X}$ such that for every $u,v\in V(G)$, we have that
$\frac{1}{225000}d_{G}(u,v)-1\leqslant d_{\textbf{IM}(G,\mathcal{X})}(f_{\mathcal{X}}(u),f_{\mathcal{X}}(v))\leqslant 160d_{G}(u,v)+160,$
where $f_{\mathcal{X}}(u)=X$ for all $u\in X\in\mathcal{X}$.
Let $\mathcal{Y}_{1}=\{N_{T^{\prime}}^{3}[r]\}$ and for each positive integer $i\geqslant 2$, let $\mathcal{Y}_{i}$ be the collection of maximal connected vertex sets in $T[N_{T^{\prime}}^{3i}[r]\backslash N_{T^{\prime}}^{3(i-1)}[r]]$. Let $\mathcal{Y}=\bigcup_{i=1}^{\infty}\mathcal{Y}_{i}$. Since $T$ is a $(0,1]$-metric tree, observe that for every $u,v\in V(T^{\prime})$, we have that
$\frac{1}{6}d_{T^{\prime}}(u,v)-1\leqslant d_{\textbf{IM}({T^{\prime}},\mathcal{Y})}(f_{\mathcal{Y}}(u),f_{\mathcal{Y}}(v))\leqslant d_{T^{\prime}}(u,v)+2,$
where $f_{\mathcal{Y}}(u)=Y$ for all $u\in Y\in\mathcal{Y}$.
Let $\mathcal{M}=\mathcal{X}\cup\mathcal{Y}$, and let $f(u)=M$ for every $u\in M\in\mathcal{M}$. Clearly $\mathcal{M}$ is a partition of $V(H)$ into connected vertex sets. Each $X\in\mathcal{X}$ has weak diameter at most 225000 in $G$ and therefore also in $H$. Similarly, each $Y\in\mathcal{Y}$ has weak diameter at most $6<225000$ in $T^{\prime}$ and therefore also in $H$. So, each $M\in\mathcal{M}$ has weak diameter at most 225000 in $H$, it follows that for every $u,v\in V(G)$, we have that $225000^{-1}d_{H}(u,v)-1\leqslant d_{\textbf{IM}(H,\mathcal{M})}(f(u),f(v))$.
Consider some $u,v\in V(H)$. Since $T^{\prime}$ is the union of at most $2g$ geodesic paths in $H$, we can find a path $P=P_{1}e_{1}P_{2}\cdots e_{n-1}P_{n}$ in $H$ between $u$ and $v$ such that
  • $P$ has length $d_{H}(u,v)$,
  • $n\leqslant 4g+1$,
  • for each $1\leqslant i<n$, $e_{i}$ is an edge of $H$ not contained in $E(T^{\prime})$ or $E(G)$, and
  • for each $1\leqslant i\leqslant n$, $P_{i}$ is a geodesic of either $T^{\prime}$ or $G$.
For each $1\leqslant i\leqslant n$, let $p_{2i-1}$ be the starting vertex of $P_{i}$, and let $p_{2i}$ be the end vertex of $P_{i}$ (these vertices need not be distinct). Clearly for each $1\leqslant i<n$, we have $d_{\textbf{IM}(H^{\prime},\mathcal{M})}(f(p_{2i}),f(p_{2i+1}))\leqslant 1$. So, for each $1\leqslant i<2n$, we have that $d_{\textbf{IM}(H^{\prime},\mathcal{M})}(f(p_{2i-1}),f(p_{2i}))\leqslant 160d_{H}(p_{2i-1},p_{2i})+160$, since each $P_{j}$ is a geodesic in both $H$ and either $T^{\prime}$ or $G$. Then,
$d_{\textbf{IM}(H,\mathcal{M})}(f(u),f(v))$ $\leqslant\sum_{i=1}^{2i-1}d_{\textbf{IM}(H,\mathcal{M})}(f(p_{i}),f(p_{i+1}))$
$\leqslant\sum_{i=1}^{2i-1}\left(160d_{H}(p_{i},p_{i+1})+160\right)$
$=160d_{H}(u,v)+160(2i-1)$
$\leqslant 160d_{H}(u,v)+160(8g+1)$
$=160d_{H}(u,v)+160+1280g,$
as desired. ∎
For our applications to Riemannian surfaces, we further need the following version of Lemma˜6.3 for locally finite graphs, which is a standard extension.
Lemma 6.4.
Let $H$ be a locally finite $(0,1]$-metric graph that triangulates a surface $\Sigma$ without boundary and with Euler genus $g$. Then there exists a partition of $V(H)$ into connected vertex sets $\mathcal{M}$ such that for every $u,v\in V(H)$, we have that
$\frac{1}{225000}d_{H}(u,v)-1\leqslant d_{\textbf{IM}(H,\mathcal{M})}(f(u),f(v))\leqslant 160d_{H}(u,v)+160+1280g,$
where $f(u)=M$ for every $u\in M\in\mathcal{M}$.
Доказательство.
Let $\mathcal{X}=\{X_{k}:k\geqslant 1\}$ be the collection of all finite subsets of $V(H)$. Fix a vertex $r$ of $H$. For each positive integer $i$, let $H_{i}$ be some finite $(0,1]$-metric graph that triangulates a surface $\Sigma^{\prime}$ of Euler genus at most $g$, contains the vertices $N_{H}^{i}[r]$ and with $N_{H_{i}}^{i}[r]=N_{H}^{i}[r]$ (note that $\Sigma^{\prime}$ may need to be chosen to be some compactification of $\Sigma$). For each $i$, let $\mathcal{M}_{i}$ be a partition of $V(H_{i})$ into connected vertex sets as given by Lemma˜6.3.
Let $J_{0}=\mathbb{N}$ and for each $k\geqslant 1$ in order, let $J_{k}\subseteq J_{k-1}$ be an infinite subset such that either $X_{k}\in\mathcal{M}_{i}$ for every $i\in J_{k}$, or $X_{k}\not\in\mathcal{M}_{i}$ for every $i\in J_{k}$. Let $\mathcal{M}$ be the subset of $\mathcal{X}$ consisting of all $X_{k}$ such that $X_{k}\in\mathcal{M}_{i}$ for every $i\in J_{k}$.
For each pair of positive integers $i,t$ with $i\geqslant t+22500$ and $t\geqslant 22500$, let $\mathcal{Y}_{i,t}$ be the subset of $\mathcal{M}_{i}$ consisting of all $M\in\mathcal{M}_{i}$ intersecting $N_{H}^{t}[r]$. Note that for every positive integer $t$, there are only finitely many possibilities for such $\mathcal{Y}_{i,t}$ and also $N_{H}^{t-22500}[r]\subseteq\bigcup_{Y\in\mathcal{Y}_{i,t}}Y$. So, it follows that $\mathcal{M}$ is a partition of $V(H)$ into connected sets. It is furthermore straightforward to check that $\mathcal{M}$ satisfies the conclusion of the lemma by using the fact that for every pair $i,t$ with $i\geqslant 10^{6}t$, we have for every $u,v\in N_{H}^{t}[r]$ that $d_{H}(u,v)=d_{H_{i}}(u,v)$ and $d_{\textbf{IM}(H,\mathcal{M}_{i})}(f_{i}(u),f_{i}(v))=d_{\textbf{IM}(H_{i},\mathcal{M}_{i})}(f_{i}(u),f_{i}(v))$, where $f_{i}(u)=M$ for every $u\in M\in\mathcal{M}_{i}$. ∎
With Lemma˜6.4 we are almost ready to prove Theorem˜1.7. First we give some required preliminaries on Riemannian surfaces.
A Riemannian surface is a surface $\Sigma$ together with a metric $d_{\Sigma}$ defined by a scalar product on the tangent space of every point. A Riemannian surface $\Sigma$ is complete if the metric space $(\Sigma,d_{\Sigma})$ is complete. Note that compact Riemannian surfaces are also complete. By the Hopf–Rinow [56] theorem, in a complete Riemannian surface $\Sigma$, subsets that are bounded and closed are compact. Restricting a complete Riemannian surface to a bounded compact subset homeomorphic to a disk gives a complete Riemannian surface homeomorphic to a disk. For more on Riemannian surfaces, see [90].
A cell decomposition of a complete Riemannian surface $\Sigma$ (without boundary) can be described as a locally finite embedded graph $G\subset\Sigma$ such that every face $F$ is homeomorphic to the unit open ball $B\subset\mathbb{R}^{2}$ via a homeomorphism $B\to\Sigma$ that extends to a continuous map of the unit disc $D$ mapping its boundary into $G$. This mapping defines a boundary path in $G$, and the boundary size of $F$ is the number of edges in its boundary path, counted with multiplicity. If every face has boundary size exactly 3 and furthermore the boundary paths all consist of 3 distinct vertices and edges, then such a cell decomposition is a triangulation. A cell decomposition is piecewise smooth if every edge in the embedding is piecewise smooth.
We need the following technical lemma of Georgakopoulos and Vigolo [48, Lemma 2.8] on modifying embedded graphs in Riemannian surfaces (actually this is a slightly less general version).
Lemma 6.5 (Georgakopoulos and Vigolo [48]).
Let $\Sigma$ be a complete Riemannian surface homeomorphic to a disk with boundary $\partial\Sigma$, and let $T\subset\Sigma$ be an embedded finite rooted tree consisting of piecewise smooth curves of finite length with root $o$ and leaves (distinct from $o$ if $o$ is a leaf) $\{p_{1},\ldots,p_{n}\}$ such that $T\cap\partial\Sigma=\{p_{1},\ldots,p_{n}\}$.
Then for every $\epsilon>0$ we may find piecewise smooth curves $\gamma_{i}$ joining $p_{i}$ to $o$ such that:
  • different $\gamma_{i}$ only meet at $o$;
  • $\bigcup_{i=1}^{n}\gamma_{i}$ is within Hausdorff distance $\epsilon$ from $T$;
  • $\left(\bigcup_{i=1}^{n}\gamma_{i}\right)\cap\partial\Sigma=\{p_{1},\ldots,p_{n}\}$.
Now, we prove the following slight weakening of Theorem˜1.7. With Lemma˜6.4 in hand, we roughly follow the idea of the proof of [48, Theorem 4.1], applying it along with Theorem˜6.1 and Lemma˜6.5.
Theorem 6.6.
Let $\Sigma$ be a complete Riemannian surface (without boundary) and with bounded Euler genus. Then there is a cell decomposition $G\subset\Sigma$ where every face has boundary of size at most three and $G^{(1)}\hookrightarrow\Sigma$ is a quasi-isometry.
Доказательство.
Let $H\subset\Sigma$ be a piecewise smooth triangulation where $H$ is a locally finite $(0,1]$-metric graph $H$ such that $H\hookrightarrow\Sigma$ is a quasi-isometry, as given by Theorem˜6.1. Since $H$ triangulates a surface of bounded genus, by Lemma˜6.4, there exists some $X\subseteq V(H)$ and a partition $\mathcal{M}=\{M_{x}:x\in X\}$ of $V(H)$ into connected sets with $x\in M_{x}$ for every $x\in X$ and such that the map $f:V(H)\to\mathcal{M}$ with $f(y)=x$ for every $y\in M_{x}$ is a quasi-isometry from $H$ to $\textbf{IM}(H,\mathcal{M})$. For each $x\in X$, let $T_{x}$ be a spanning tree of $H[M_{x}]$ with root $x$. For some sufficiently small $\epsilon_{x}>0$ we can choose some $N_{x}\subseteq\Sigma$ with $N_{x}=\{p:d_{\Sigma}(T_{x},p)\leqslant\epsilon_{x}\}$ such that $N_{x}\cap H$ is a tree $T_{x}^{*}$ (whose leaves were not vertices of $H$) and $N_{x}$ is homeomorphic to a disk. Restricting $\Sigma$ to $N_{x}$ yields a complete Riemannian surface homeomorphic to the disk. The leaves of $T_{x}^{*}$ are exactly the points of $T_{x}^{*}$ intersecting $\partial N_{x}$. Applying Lemma˜6.5 to every $T_{x}^{*}\subseteq N_{x}$ yields a cell decomposition $G\subset\Sigma$ where every face has boundary size at most three. For an illustration of this, see Fig.˜9. Since the graph $G$ (or essentially its simplicial 1-skeleton $G^{(1)}$) is obtained from $\textbf{IM}(H,\mathcal{M})$ by possibly adding loops and multiple edges, it can be seen from the quasi-isometry $f$ that $G^{(1)}\hookrightarrow\Sigma$ is also a quasi-isometry. ∎
Рис. 9: On the left is an illustration of the triangulation $H\subset\Sigma$. The induced subgraph $H[M_{x}]$ is the one consisting of the (black) rooted tree $T_{x}$ and the other (red) edges between vertices of $M_{x}$. The rest of $H$ is blue. For each such $x$, we apply Lemma˜6.5 to modify the cell decomposition as on the right. By doing this for every $x\in X$, we obtain the desired cell decomposition $G\subset\Sigma$.
Georgakopoulos and Vigolo [48, Lemma 2.7] showed that by taking some subgraph $G_{1}$ of $G$, such a cell decomposition can be refined so that every face has boundary of size exactly 3, while preserving the quasi-isometry $G_{1}^{(1)}\hookrightarrow\Sigma$. They also further observed [48, Remark 2.6] that by taking the barycentric subdivision of $G_{1}\subset\Sigma$, such a cell decomposition can be turned into a triangulation $G_{2}\subset\Sigma$, again while preserving the quasi-isometry $G_{2}^{(1)}\hookrightarrow\Sigma$. Thus, we obtain Theorem˜1.7 from Theorem˜6.6.
It can easily be checked that the proof of Theorem˜1.7 yields a $(M,A+O(g))$-quasi-isometry where $M$ and $A$ are absolute constants independent of the Euler genus. In fact, a close inspection shows that it yields a $(10^{6},10^{6}+10^{4}g)$-quasi-isometry.
Acknowledgements
This work was initiated at the Workshop on Graph Product Structure Theory (BIRS21w5235) at the Banff International Research Station, 21–26 November 2021. We are grateful to the organisers and participants for providing a stimulating research environment. We thank a number of people for helpful discussions and encouragement on the project including Sandra Albrechtsen, Marthe Bonamy, Linda Cook, Louis Esperet, Agelos Georgakopoulos, Robert Hickingbotham, Rose McCarty, Harry Petyt, Federico Vigolo, and Alexandra Wesolek.
Список литературы
  • [1] T. Abrishami, M. Briański, J. Davies, X. Du, J. Masaříková, P. Rzążewski, and B. Walczak (2025) Burling graphs in graphs with large chromatic number. arXiv preprint arXiv:2510.19650. Cited by: §1, §1.
  • [2] S. Albrechtsen and J. Davies (2025) Counterexample to the conjectured coarse grid theorem. arXiv preprint arXiv:2508.15342. Cited by: §1.
  • [3] S. Albrechtsen, M. Distel, and A. Georgakopoulos (2025) Excluding $K_{2,t}$} as a fat minor. arXiv preprint arXiv:2510.14644. Cited by: §1.
  • [4] S. Albrechtsen, T. Huynh, R. W. Jacobs, P. Knappe, and P. Wollan (2024) A menger-type theorem for two induced paths. SIAM Journal on Discrete Mathematics 38 (2), pp. 1438–1450. Cited by: §1.
  • [5] S. Albrechtsen, R. W. Jacobs, P. Knappe, and P. Wollan (2024) A characterisation of graphs quasi-isometric to $K_{4}$}-minor-free graphs. arXiv preprint arXiv:2408.15335. Cited by: §1.
  • [6] K. Appel, W. Haken, and J. Koch (1977) Every planar map is four colorable. Part II: Reducibility. Illinois Journal of Mathematics 21 (3), pp. 491–567. Cited by: §1.
  • [7] G. N. Arzhantseva and P. Cherix (2004) On the Cayley graph of a generic finitely presented group. Bulletin of the Belgian Mathematical Society-Simon Stevin 11 (4), pp. 589–601. Cited by: §1.
  • [8] B. Awerbuch and D. Peleg (1990) Sparse partitions. In Proceedings [1990] 31st Annual Symposium on Foundations of Computer Science, pp. 503–513. Cited by: §1.
  • [9] G. Bell and A. Dranishnikov (2008) Asymptotic dimension. Topology and its Applications 155 (12), pp. 1265–1296. Cited by: §1.
  • [10] S. Benzer (1959) On the topology of the genetic fine structure. Proceedings of the National Academy of Sciences 45 (11), pp. 1607–1620. Cited by: §1.
  • [11] A. Bernshteyn and J. Yu (2025) Large-scale geometry of borel graphs of polynomial growth. Advances in Mathematics 473, pp. 110290. Cited by: §1.
  • [12] M. Bonamy, N. Bousquet, L. Esperet, C. Groenland, C. Liu, F. Pirot, and A. Scott (2023) Asymptotic dimension of minor-closed families and assouad–nagata dimension of surfaces. Journal of the European Mathematical Society 26 (10), pp. 3739–3791. Cited by: §1, §1, §1, §1.
  • [13] M. Bonamy, C. Gavoille, T. Picavet, and A. Wesolek (2025) Distributed approximation algorithms for minimum dominating set in locally nice graphs. arXiv preprint arXiv:2507.04960. Cited by: §1.
  • [14] M. Bonamy, C. Gavoille, T. Picavet, and A. Wesolek (2025) Local constant approximation for dominating set on graphs excluding large minors. In Proceedings of the ACM Symposium on Principles of Distributed Computing, pp. 77–87. Cited by: §1.
  • [15] M. Bonk and B. Kleiner (2002) Quasisymmetric parametrizations of two-dimensional metric spheres. Inventiones mathematicae 150 (1), pp. 127–183. Cited by: §1.
  • [16] B. H. Bowditch (2004) Planar groups and the Seifert conjecture. Journal für die reine und angewandte Mathematik 576, pp. 11–62. Cited by: §1.
  • [17] B. H. Bowditch (2020) Bilipschitz triangulations of riemannian manifolds. Cited by: §1.
  • [18] D. Burago, Y. Burago, and S. Ivanov (2001) A course in metric geometry. Vol. 33, American Mathematical Soc.. Cited by: §1, §1.
  • [19] A. Casson and D. Jungreis (1994) Convergence groups and Seifert fibered 3-manifolds. Inventiones mathematicae 118 (1), pp. 441–456. Cited by: §1.
  • [20] H. Chang, J. Conroy, H. Le, L. Milenkovic, S. Solomon, and C. Than (2023) Covering planar metrics (and beyond): O (1) trees suffice. In 2023 IEEE 64th Annual Symposium on Foundations of Computer Science (FOCS), pp. 2231–2261. Cited by: Remark 1.10.
  • [21] H. Chang, J. Conroy, Z. Tan, and D. W. Zheng (2026) Cutting planarians: planar emulators for string graphs. Accepted to STOC 2026, 58th Annual ACM Symposium on Theory of Computing. Cited by: Remark 1.10, Remark 1.10.
  • [22] V. Chepoi, F. F. Dragan, I. Newman, Y. Rabinovich, and Y. Vaxes (2012) Constant approximation algorithms for embedding graph metrics into trees and outerplanar graphs. Discrete & Computational Geometry 47 (1), pp. 187–214. Cited by: §1, §3.
  • [23] P. Creutz and M. Romney (2022) Triangulating metric surfaces. Proceedings of the London Mathematical Society 125 (6), pp. 1426–1451. Cited by: §1, §1.
  • [24] J. Davies, M. Distel, and R. Hickingbotham Coarse structure for minor-free metrics. (In preparation). Cited by: Remark 1.10.
  • [25] J. Davies, A. Georgakopoulos, M. Hatzel, and R. McCarty (2025) Strongly sublinear separators and bounded asymptotic dimension for sphere intersection graphs. arXiv preprint arXiv:2504.00932. Cited by: §1, §1.
  • [26] J. Davies, R. Hickingbotham, F. Illingworth, and R. McCarty (2024) Fat minors cannot be thinned (by quasi-isometries). arXiv preprint arXiv:2405.09383. Cited by: §1.
  • [27] J. Davies (2026) Surfaces without quasi-isometric simplicial triangulations. arXiv preprint arXiv:2603.26652. Cited by: §1.
  • [28] A. Denise, M. Vasconcellos, and D. J. Welsh (1996) The random planar graph. Congressus numerantium, pp. 61–80. Cited by: §1.
  • [29] M. Distel (2023) Proper minor-closed classes of graphs have Assouad-Nagata dimension 2. arXiv preprint arXiv:2308.10377. Cited by: §1.
  • [30] C. Droms and D. Robinson (2006) Infinite-ended groups with planar Cayley graphs.. Journal of Group Theory 9 (4). Cited by: §1.
  • [31] C. Droms, B. Servatius, and H. Servatius (1998) Connectivity and planarity of Cayley graphs. Beiträge Algebra Geom 39 (2), pp. 269–282. Cited by: §1.
  • [32] M. J. Dunwoody (2007) Planar graphs and covers. arXiv preprint arXiv:0708.0920. Cited by: §1, §1.
  • [33] Z. Dvořák and S. Norin (2025) Asymptotic dimension of intersection graphs. European Journal of Combinatorics 123, pp. 103631. Cited by: §1.
  • [34] G. Ehrlich, S. Even, and R. E. Tarjan (1976) Intersection graphs of curves in the plane. Journal of Combinatorial Theory, Series B 21 (1), pp. 8–20. Cited by: §1.
  • [35] A. Erschler and I. Mitrofanov (2023) Assouad–nagata dimension and gap for ordered metric spaces. Commentarii Mathematici Helvetici 98 (2), pp. 217–260. Cited by: §1.
  • [36] L. Esperet, U. Giocanti, and C. Legrand-Duchesne (2024) The structure of quasi-transitive graphs avoiding a minor with applications to the domino problem. Journal of Combinatorial Theory, Series B 169, pp. 561–613. Cited by: §1.
  • [37] A. Filtser (2024) Scattering and sparse partitions, and their applications. ACM Transactions on Algorithms 20 (4), pp. 1–42. Cited by: §1.
  • [38] J. Fox and J. Pach (2008) Separator theorems and Turán-type results for planar intersection graphs. Advances in Mathematics 219 (3), pp. 1070–1080. Cited by: §1.
  • [39] J. Fox and J. Pach (2010) A separator theorem for string graphs and its applications. Combinatorics, Probability and Computing 19 (3), pp. 371–390. Cited by: §1.
  • [40] J. Fox and J. Pach (2012) String graphs and incomparability graphs. Advances in mathematics 230 (3), pp. 1381–1401. Cited by: §1.
  • [41] J. Fox and J. Pach (2014) Applications of a new separator theorem for string graphs. Combinatorics, Probability and Computing 23 (1), pp. 66–74. Cited by: §1.
  • [42] K. Fujiwara and P. Papasoglu (2021) Asymptotic dimension of planes and planar graphs. Transactions of the American Mathematical Society 374 (12), pp. 8887–8901. Cited by: §1.
  • [43] K. Fujiwara and P. Papasoglu (2023) A coarse-geometry characterization of cacti. arXiv preprint arXiv:2305.08512. Cited by: §1, §3.
  • [44] D. Gabai (1992) Convergence groups are Fuchsian groups. Annals of Mathematics 136 (3), pp. 447–510. Cited by: §1.
  • [45] A. Georgakopoulos and M. Hamann (2019) The planar Cayley graphs are effectively enumerable I: consistently planar graphs. Combinatorica 39 (5), pp. 993–1019. Cited by: §1.
  • [46] A. Georgakopoulos and M. Hamann (2019) The planar Cayley graphs are effectively enumerable II. arXiv preprint arXiv:1901.00347. Cited by: §1.
  • [47] A. Georgakopoulos and P. Papasoglu (2023) Graph minors and metric spaces. arXiv preprint arXiv:2305.07456. Cited by: §1, §1, §1.
  • [48] A. Georgakopoulos and F. Vigolo (2026) Tirangulating surfaces quasi-isometrically. arXiv preprint arXiv:2603.21189. Cited by: §1, §1, §1, Theorem 6.1, Lemma 6.5, §6, §6, §6, §6.
  • [49] A. Georgakopoulos (2017) The planar cubic Cayley graphs. Memoirs of the American Mathematical Society 250 (1190). Cited by: §1.
  • [50] A. Georgakopoulos (2020) On planar Cayley graphs and Kleinian groups. Transactions of the American Mathematical Society 373 (7), pp. 4649–4684. Cited by: §1, §1.
  • [51] M. Gromov (1993) Geometric group theory. Volume 2: Asymptotic invariants of infinite groups. Proceedings of the symposium held at the Sussex University, Brighton, July 14-19, 1991. Lond. Math. Soc. Lect. Note Ser., Vol. 182, Cambridge: Cambridge University Press (English). External Links: ISSN 0076-0552, ISBN 0-521-44680-5 Cited by: §1.
  • [52] A. Gupta, I. Newman, Y. Rabinovich, and A. Sinclair (2004) Cuts, trees and $\ell_{1}$-embeddings of graphs. Combinatorica 24 (2), pp. 233–269. Cited by: §1, §1.
  • [53] M. Hamann (2018) Accessibility in transitive graphs. Combinatorica 38 (4), pp. 847–859. Cited by: §1.
  • [54] M. Hamann (2018) Planar transitive graphs. The Electronic Journal of Combinatorics 25 (4), pp. 4–8. Cited by: §1.
  • [55] J. Hopcroft and R. Tarjan (1974) Efficient planarity testing. Journal of the ACM (JACM) 21 (4), pp. 549–568. Cited by: §1.
  • [56] H. Hopf and W. Rinow (1931) Ueber den begriff der vollständigen differentialgeometrischen fläche. Commentarii Mathematici Helvetici 3 (1), pp. 209–225. Cited by: §6.
  • [57] L. Jia, G. Lin, G. Noubir, R. Rajaraman, and R. Sundaram (2005) Universal approximations for TSP, Steiner tree, and set cover. In Proceedings of the thirty-seventh annual ACM symposium on Theory of computing, pp. 386–395. Cited by: §1.
  • [58] M. Jørgensen and U. Lang (2022) Geodesic spaces of low Nagata dimension. Annales Fennici Mathematici 47, pp. 83–88. Cited by: §1.
  • [59] P. Koebe (1936) Kontaktprobleme der konformen Abbildung. Berichte über die Verhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Physische Klasse 88, pp. 141–164. Cited by: §1.
  • [60] D. König (1927) Über eine schlussweise aus dem endlichen ins unendliche. Acta Sci. Math.(Szeged) 3 (2-3), pp. 121–130. Cited by: §2.
  • [61] J. Kratochvíl and J. Matoušek (1991) String graphs requiring exponential representations. Journal of Combinatorial Theory, Series B 53 (1), pp. 1–4. Cited by: §1.
  • [62] J. Kratochvíl (1991) String graphs. II. Recognizing string graphs is NP-hard. Journal of Combinatorial Theory, Series B 52 (1), pp. 67–78. Cited by: §1.
  • [63] J. R. Lee (2016) Separators in region intersection graphs. arXiv preprint arXiv:1608.01612. Cited by: §1, §2, §5.
  • [64] C. Liu (2025) Assouad–Nagata dimension of minor-closed metrics. Proceedings of the London Mathematical Society 130 (3), pp. e70032. Cited by: §1, §1.
  • [65] D. Lokshtanov, F. Panolan, S. Saurabh, J. Xue, and M. Zehavi (2024) A 1.9999-approximation algorithm for vertex cover on string graphs. In 40th International Symposium on Computational Geometry (SoCG 2024), Vol. 293, pp. 72. Cited by: §1.
  • [66] J. MacManus (2023) Accessibility, planar graphs, and quasi-isometries. arXiv preprint arXiv:2310.15242. Cited by: §1, §1, §1.
  • [67] J. MacManus (2024) A note on quasi-transitive graphs quasi-isometric to planar (Cayley) graphs. arXiv preprint arXiv:2407.13375. Cited by: §1.
  • [68] S. Maillot (2001) Quasi-isometries of groups, graphs and surfaces. Commentarii Mathematici Helvetici 76 (1), pp. 29–60. Cited by: §1, §1.
  • [69] J. F. Manning (2005) Geometry of pseudocharacters. Geometry & Topology 9 (2), pp. 1147–1185. Cited by: §1.
  • [70] H. Maschke (1896) The representation of finite groups, especially of the rotation groups of the regular bodies of three-and four-dimensional space, by Cayley’s color diagrams. American Journal of Mathematics 18 (2), pp. 156–194. Cited by: §1.
  • [71] J. Matoušek (2014) Near-optimal separators in string graphs. Combinatorics, Probability and Computing 23 (1), pp. 135–139. Cited by: §1.
  • [72] K. Menger (1927) Zur allgemeinen kurventheorie. Fundamenta mathematicae 10 (1), pp. 96–115. Cited by: §1.
  • [73] G. Mess (1988) The Seifert conjecture and groups which are coarse quasiisometric to planes. preprint. Cited by: §1.
  • [74] T. Nguyen, A. Scott, and P. Seymour (2025) A counterexample to the coarse menger conjecture. Journal of Combinatorial Theory, Series B 173, pp. 68–82. Cited by: §1.
  • [75] T. Nguyen, A. Scott, and P. Seymour (2025) Asymptotic structure. III. Excluding a fat tree. arXiv preprint arXiv:2509.09035. Cited by: §1.
  • [76] T. Nguyen, A. Scott, and P. Seymour (2025) Asymptotic structure. iv. a counterexample to the weak coarse menger conjecture. arXiv preprint arXiv:2508.14332. Cited by: §1.
  • [77] T. Nguyen, A. Scott, and P. Seymour (2025) Asymptotic structure. v. the coarse menger conjecture in bounded path-width. arXiv preprint arXiv:2509.08762. Cited by: §1.
  • [78] T. Nguyen, A. Scott, and P. Seymour (2025) Asymptotic structure. vi. distant paths across a disc. arXiv preprint arXiv:2509.07174. Cited by: §1.
  • [79] S. Norine, P. Seymour, R. Thomas, and P. Wollan (2006) Proper minor-closed families are small. Journal of Combinatorial Theory, Series B 96 (5), pp. 754–757. Cited by: §1.
  • [80] D. Ntalampekos and M. Romney (2023) Polyhedral approximation of metric surfaces and applications to uniformization. Duke Mathematical Journal 172 (9), pp. 1673–1734. Cited by: §1, §1.
  • [81] M. I. Ostrovskii and D. Rosenthal (2015) Metric dimensions of minor excluded graphs and minor exclusion in groups. International Journal of Algebra and Computation 25 (04), pp. 541–554. Cited by: §1.
  • [82] J. Pach and G. Tóth (2002) Recognizing string graphs is decidable. Discrete & Computational Geometry 28, pp. 593–606. Cited by: §1.
  • [83] J. Pach and G. Tóth (2006) How many ways can one draw a graph?. Combinatorica 26 (5), pp. 559–576. Cited by: §1.
  • [84] A. Pawlik, J. Kozik, T. Krawczyk, M. Lasoń, P. Micek, W. T. Trotter, and B. Walczak (2014) Triangle-free intersection graphs of line segments with large chromatic number. Journal of Combinatorial Theory, Series B 105, pp. 6–10. Cited by: §1.
  • [85] S. Rao (1999) Small distortion and volume preserving embeddings for planar and euclidean metrics. In Proceedings of the fifteenth annual symposium on Computational geometry, pp. 300–306. Cited by: §1.
  • [86] N. Robertson, D. Sanders, P. Seymour, and R. Thomas (1997) The four-colour theorem. Journal of combinatorial theory, Series B 70 (1), pp. 2–44. Cited by: §1.
  • [87] E. Saucan (2008) Intrinsic differential geometry and the existence of quasimeromorphic mappings. arXiv preprint arXiv:0901.0125. Cited by: §1, §1.
  • [88] M. Schaefer, E. Sedgwick, and D. Štefankovič (2002) Recognizing string graphs in NP. In Proceedings of the thiry-fourth annual ACM symposium on Theory of computing, pp. 1–6. Cited by: §1.
  • [89] F. W. Sinden (1966) Topology of thin film RC circuits. Bell System Technical Journal 45 (9), pp. 1639–1662. Cited by: §1.
  • [90] M. Spivak (1979) A comprehensive introduction to differential geometry, publish or perish. Inc., Berkeley 2. Cited by: §6.
  • [91] I. Tomon (2023) String graphs have the Erdős–Hajnal property. Journal of the European Mathematical Society 26 (1), pp. 275–287. Cited by: §1.
  • [92] T. W. Tucker (1983) Finite groups acting on surfaces and the genus of a group. Journal of Combinatorial Theory, Series B 34 (1), pp. 82–98. Cited by: §1.
  • [93] P. Tukia (1988) Homeomorphic conjugates of Fuchsian groups. Journal für die reine und angewandte Mathematik 391, pp. 1–54. Cited by: §1.
  • [94] K. Wagner (1967) Fastplättbare graphen. Journal of Combinatorial Theory 3 (4), pp. 326–365. Cited by: footnote 1.
  • [95] G. Yu (1998) The Novikov conjecture for groups with finite asymptotic dimension. Annals of Mathematics 147 (2), pp. 325–355. Cited by: §1.